The Objective Caml system release 1.07

Documentation and user’s manual

Xavier Leroy December 10, 1997

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1997 Institut National de Recherche en Informatique et Automatique 2

Contents

I An introduction to Objective Caml 7

1 The core language 9

2 Objects in Caml 23

3 The module system 39

II The Objective Caml language 49

4 The Objective Caml language 51 4.1 Lexical conventions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 4.2 Values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 4.3 Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 4.4 Type expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 4.5 Constants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 4.6 Patterns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 4.7 Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 4.8 Type and exception definitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 4.9 Classes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 4.10 Module types (module specifications) . . . . . . . . . . . . . . . . . . . . . . . . . . 76 4.11 Module expressions (module implementations) . . . . . . . . . . . . . . . . . . . . . 79 4.12 Compilation units . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82

5 Language extensions 83 5.1 Streams and stream parsers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 5.2 Range patterns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 5.3 Assertion checking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 5.4 Deferred computations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85

III The Objective Caml tools 87

6 Batch compilation (ocamlc) 89 6.1 Overview of the compiler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 6.2 Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 6.3 Modules and the file system . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 6.4 Common errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

7 The toplevel system (ocaml) 97 7.1 Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 7.2 Toplevel directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 7.3 The toplevel and the module system . . . . . . . . . . . . . . . . . . . . . . . . . . 101 7.4 Common errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 7.5 Building custom toplevel systems: ocamlmktop . . . . . . . . . . . . . . . . . . . . . 101 7.6 Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102

8 The runtime system (ocamlrun) 103 8.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 8.2 Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 8.3 Common errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104

9 Native-code compilation (ocamlopt) 107 9.1 Overview of the compiler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 9.2 Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 9.3 Common errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 9.4 Compatibility with the bytecode compiler . . . . . . . . . . . . . . . . . . . . . . . 110

10 Lexer and parser generators (ocamllex, ocamlyacc) 113 10.1 Overview of ocamllex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 10.2 Syntax of lexer definitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 10.3 Overview of ocamlyacc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 10.4 Syntax of grammar definitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 10.5 Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 10.6 A complete example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118

11 Dependency generator (ocamldep) 121 11.1 Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 11.2 A typical Makefile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121

12 The debugger (ocamldebug) 123 12.1 Compiling for debugging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 12.2 Invocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 12.3 Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 12.4 Executing a program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 12.5 Breakpoints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 12.6 The call stack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 12.7 Examining variable values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 12.8 Controlling the debugger . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 12.9 Miscellaneous commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 12.10 Running the debugger under Emacs . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 13 Profiling (ocamlprof) 137

13.1 Compiling for profiling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 13.2 Profiling an execution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 13.3 Printing profiling information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 13.4 Time profiling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138

14 Interfacing C with Objective Caml 139 14.1 Overview and compilation information . . . . . . . . . . . . . . . . . . . . . . . . . 139 14.2 The value type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 14.3 Representation of Caml data types . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 14.4 Operations on values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 14.5 Living in harmony with the garbage collector . . . . . . . . . . . . . . . . . . . . . 147 14.6 Callbacks from C to Caml . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 14.7 A complete example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 14.8 Advanced example with callbacks . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154

IV The Objective Caml library 157

15 The core library 159 15.1 Module Pervasives: the initially opened module . . . . . . . . . . . . . . . . . . . 159

16 The standard library 175 16.1 Module Arg: parsing of command line arguments . . . . . . . . . . . . . . . . . . . 176 16.2 Module Array: array operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 16.3 Module Callback: registering Caml values with the C runtime . . . . . . . . . . . . 180 16.4 Module Char: character operations . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 16.5 Module Digest: MD5 message digest . . . . . . . . . . . . . . . . . . . . . . . . . . 181 16.6 Module Filename: operations on file names . . . . . . . . . . . . . . . . . . . . . . 182 16.7 Module Format: pretty printing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 16.8 Module Gc: memory management control and statistics . . . . . . . . . . . . . . . . 189 16.9 Module Genlex: a generic lexical analyzer . . . . . . . . . . . . . . . . . . . . . . . 192 16.10 Module Hashtbl: hash tables and hash functions . . . . . . . . . . . . . . . . . . . 193 16.11 Module Lazy: deferred computations. . . . . . . . . . . . . . . . . . . . . . . . . . . 195 16.12 Module Lexing: the run-time library for lexers generated by camllex . . . . . . . . 195 16.13 Module List: list operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 16.14 Module Map: association tables over ordered types . . . . . . . . . . . . . . . . . . . 199 16.15 Module Marshal: marshaling of data structures . . . . . . . . . . . . . . . . . . . . 201 16.16 Module Oo: object-oriented extension . . . . . . . . . . . . . . . . . . . . . . . . . . 203 16.17 Module Parsing: the run-time library for parsers generated by camlyacc . . . . . . 203 16.18 Module Printexc: a catch-all exception handler . . . . . . . . . . . . . . . . . . . . 203 16.19 Module Printf: formatting printing functions . . . . . . . . . . . . . . . . . . . . . 204 16.20 Module Queue: first-in first-out queues . . . . . . . . . . . . . . . . . . . . . . . . . 205 16.21 Module Random: pseudo-random number generator . . . . . . . . . . . . . . . . . . 206 16.22 Module Set: sets over ordered types . . . . . . . . . . . . . . . . . . . . . . . . . . 206 16.23 Module Sort: sorting and merging lists . . . . . . . . . . . . . . . . . . . . . . . . . 208 16.24 Module Stack: last-in first-out stacks . . . . . . . . . . . . . . . . . . . . . . . . . . 208 16.25 Module Stream: streams and parsers . . . . . . . . . . . . . . . . . . . . . . . . . . 209 16.26 Module String: string operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 16.27 Module Sys: system interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 16.28 Module Weak: arrays of weak pointers . . . . . . . . . . . . . . . . . . . . . . . . . . 215

17 The unix library: Unix system calls 217 17.1 Module Unix: interface to the Unix system . . . . . . . . . . . . . . . . . . . . . . . 218

18 The num library: arbitrary-precision rational arithmetic 239 18.1 Module Num: operation on arbitrary-precision numbers . . . . . . . . . . . . . . . . 240 18.2 Module Arith_status: flags that control rational arithmetic . . . . . . . . . . . . . 243

19 The str library: regular expressions and string processing 245 19.1 Module Str: regular expressions and high-level string processing . . . . . . . . . . . 245

20 The threads library 249 20.1 Module Thread: lightweight threads . . . . . . . . . . . . . . . . . . . . . . . . . . . 250 20.2 Module Mutex: locks for mutual exclusion . . . . . . . . . . . . . . . . . . . . . . . 251 20.3 Module Condition: condition variables to synchronize between threads . . . . . . . 252 20.4 Module Event: first-class synchronous communication . . . . . . . . . . . . . . . . . 253 20.5 Module ThreadUnix: thread-compatible system calls . . . . . . . . . . . . . . . . . 254

21 The graphics library 257 21.1 Module Graphics: machine-independent graphics primitives . . . . . . . . . . . . . 258

22 The dbm library: access to NDBM databases 265 22.1 Module Dbm: interface to the NDBM databases . . . . . . . . . . . . . . . . . . . . 265

V Appendix 267

Index to the library 269

Index of keywords 279

Foreword

This manual documents the release 1.07 of the Objective Caml system. It is organized as follows.

• Part I, “An introduction to Objective Caml”, gives an overview of the language.

• Part II, “The Objective Caml language”, is the reference description of the language.

• Part III, “The Objective Caml tools, documents the compilers, toplevel system, and programming utilities.

• Part IV, “The Objective Caml library”, describes the modules provided in the standard library.

• Part V, “Appendix”, contains an index of all identifiers defined in the standard library, and an index of keywords.

Conventions

Objective Caml runs on several operating systems. The parts of this manual that are specific to one operating system are presented as shown below:

Unix:

This is material specific to Unix.

Windows:

This is material specific to MS Windows (NT and 95).

License

The Objective Caml system is copyright \begin{matrix}c \\ ◯ \end{matrix} 1996, 1997 Institut National de Recherche en Informatique et en Automatique (INRIA). INRIA holds all ownership rights to the Objective Caml system. See the file LICENSE in the distribution for licensing information.

The Objective Caml system can be freely redistributed. More precisely, INRIA grants any user of the Objective Caml system the right to reproduce it, provided that the copies are distributed under the conditions given in the LICENSE file. The present documentation is distributed under the same conditions.

Availability

The complete Objective Caml distribution resides on the machine ftp.inria.fr. The distribution files can be transferred by anonymous FTP:

               Host:  ftp.inria.fr (Internet address 192.93.2.54)

Login name: anonymous
Password:  your e-mail address

               Directory:  lang/caml-light

Files:  see the index in file README

More information on the Caml family of languages is also available on the World Wide Web,

http://pauillac.inria.fr/caml/.

Part I

Chapter 1

An introduction to Objective Caml The core language

This part of the manual is a tutorial introduction to the Objective Caml language. A good familiarity with programming in a conventional languages (say, Pascal or C) is assumed, but no prior exposure to functional languages is required. The present chapter introduces the core language. Chapter 2 deals with the object-oriented features, and chapter 3 with the module system.

Basics

For this overview of Caml, we use the interactive system, which is started by running ocaml from the Unix shell, or by launching the OCamlwin.exe application under Windows. This tutorial is presented as the transcript of a session with the interactive system: lines starting with # represent user input; the system responses are printed below, without a leading #.

Under the interactive system, the user types Caml phrases, terminated by ;;, in response to the # prompt, and the system compiles them on the fly, executes them, and prints the outcome of evaluation. Phrases are either simple expressions, or let definitions of identifiers (either values or functions).

# 1+2*3;;
- : int = 7
# let pi = 4.0 *. atan 1.0;;
val pi : float = 3.14159265359
# let square x = x *. x;;
val square : float -> float = <fun>
# square(sin pi) +. square(cos pi);;
- : float = 1

The Caml system computes both the value and the type for each phrase. Even function parameters need no explicit type declaration: the system infers their types from their usage in the function. Notice also that integers and floating-point numbers are distinct types, with distinct operators: + and * operate on integers, but +. and *. operate on floats.

# 1.0 * 2;;
Characters 0-3:
This expression has type float but is here used with type int

Recursive functions are defined with the let rec binding:

# let rec fib n =
#   if n < 2 then 1 else fib(n-1) + fib(n-2);;
val fib : int -> int = <fun>
# fib 10;;
- : int = 89

Data types

In addition to integers and floating-point numbers, Caml offers the usual basic data types: booleans, characters, and character strings.

# (1 < 2) = false;;
- : bool = false
# 'a';;
- : char = 'a'
# "Hello world";;
- : string = "Hello world"

Predefined data structures include tuples, arrays, and lists. General mechanisms for defining your own data structures are also provided. They will be covered in more details later; for now, we concentrate on lists. Lists are either given in extension as a bracketed list of semicolon-separated elements, or built from the empty list [] (pronounce “nil”) by adding elements in front using the :: (“cons”) operator.

# let l = ["is"; "a"; "tale"; "told"; "etc."];;
val l : string list = ["is"; "a"; "tale"; "told"; "etc."]
# "Life" :: l;;
- : string list = ["Life"; "is"; "a"; "tale"; "told"; "etc."]

As with all other Caml data structures, lists do not need to be explicitly allocated and deallocated from memory: all memory management is entirely automatic in Caml. Similarly, there is no explicit handling of pointers: the Caml compiler silently introduces pointers where necessary.

As with most Caml data structures, inspecting and destructuring lists is performed by patternmatching. List patterns have the exact same shape as list expressions, with identifier representing unspecified parts of the list. As an example, here is insertion sort on a list:

# let rec sort lst =
#   match lst with
#     [] -> []
#   | head :: tail -> insert head (sort tail)
# and insert elt lst =
#   match lst with
#     [] -> [elt]

#   | head :: tail -> if elt <= head then elt :: lst else head :: insert elt tail
# ;;
val sort : 'a list -> 'a list = <fun>
val insert : 'a -> 'a list -> 'a list = <fun>
# sort l;;
- : string list = ["a"; "etc."; "is"; "tale"; "told"]

The type inferred for sort, 'a list -> 'a list, means that sort can actually apply to lists of any type, and returns a list of the same type. The type 'a is a type variable, and stands for any given type. The reason why sort can apply to lists of any type is that the comparisons (=, <=, etc.) are polymorphic in Caml: they operate between any two values of the same type. This makes sort itself polymorphic over all list types.

# sort [6;2;5;3];;
- : int list = [2; 3; 5; 6]
# sort [3.14; 2.718];;
- : float list = [2.718; 3.14]

The sort function above does not modify its input list: it builds and returns a new list containing the same elements as the input list, in ascending order. There is actually no way in Caml to modify in-place a list once it is built: we say that lists are immutable data structures. Most Caml data structures are immutable, but a few (most notably arrays) are mutable, meaning that they can be modified in-place at any time.

Functions as values

Caml is a functional language: functions in the full mathematical sense are supported and can be passed around freely just as any other piece of data. For instance, here is a deriv function that takes any float function as argument and returns an approximation of its derivative function:

# let deriv f dx = function x -> (f(x +. dx) -. f(x)) /. dx;;
val deriv : (float -> float) -> float -> float -> float = <fun>
# let cos' = deriv sin 1e-6;;
val cos' : float -> float = <fun>
# cos' pi;;
- : float = -1.00000000014

Even function composition is definable:

# let compose f g = function x -> f(g(x));;
val compose : ('a -> 'b) -> ('c -> 'a) -> 'c -> 'b = <fun>
# let cos2 = compose square cos;;
val cos2 : float -> float = <fun>

Functions that take other functions as arguments are called “functionals”, or “higher-order functions”. Functionals are especially useful to provide iterators or similar generic operations over a data structure. For instance, the standard Caml library provides a List.map functional that applies a given function to each element of a list, and returns the list of the results:

# List.map (function n -> n * 2 + 1) [0;1;2;3;4];;
- : int list = [1; 3; 5; 7; 9]

This functional, along with a number of other list and array functionals, is predefined because it is often useful, but there is nothing magic with it: it can easily be defined as follows.

# let rec map f l =
#   match l with
#     [] -> []
#   | hd :: tl -> f hd :: map f tl;;
val map : ('a -> 'b) -> 'a list -> 'b list = <fun>

Records and variants

User-defined data structures include records and variants. Both are defined with the type declaration. Here, we declare a record type to represent rational numbers.

# type ratio = {num: int; denum: int};;
type ratio = { num: int; denum: int }
# let add_ratio r1 r2 =
#   {num = r1.num * r2.denum + r2.num * r1.denum;
#    denum = r1.denum * r2.denum};;
val add_ratio : ratio -> ratio -> ratio = <fun>
# add_ratio {num=1; denum=3} {num=2; denum=5};;
- : ratio = {num=11; denum=15}

The declaration of a variant type lists all possible shapes for values of that type. Each case is identified by a name, called a constructor, which serves both for constructing values of the variant type and inspecting them by pattern-matching. Constructor names are capitalized to distinguish them from variable names (which must start with a lowercase letter). For instance, here is a variant type for doing mixed arithmetic (integers and floats):

# type number = Int of int | Float of float | Error;;
type number = | Int of int | Float of float | Error

This declaration expresses that a value of type number is either an integer, a floating-point number, or the constant Error representing the result of an invalid operation (e.g. a division by zero).

Enumerated types are a special case of variant types, where all alternatives are constants:

# type sign = Positive | Negative;;
type sign = | Positive | Negative
# let sign_int n = if n >= 0 then Positive else Negative;;
val sign_int : int -> sign = <fun>

To define arithmetic operations for the number type, we use pattern-matching on the two numbers involved:

# let add_num n1 n2 =
#   match (n1, n2) with
#     (Int i1, Int i2) ->
#       (* Check for overflow of integer addition *)
#       if sign_int i1 = sign_int i2 && sign_int(i1 + i2) <> sign_int i1
#       then Float(float i1 +. float i2)
#       else Int(i1 + i2)
#   | (Int i1, Float f2) -> Float(float i1 +. f2)
#   | (Float f1, Int i2) -> Float(f1 +. float i2)
#   | (Float f1, Float f2) -> Float(f1 +. f2)
#   | (Error, _) -> Error
#   | (_, Error) -> Error;;
val add_num : number -> number -> number = <fun>
# add_num (Int 123) (Float 3.14159);;
- : number = Float 126.14159

The most common usage of variant types is to describe recursive data structures. Consider for example the type of binary trees:

# type 'a btree = Empty | Node of 'a * 'a btree * 'a btree;;
type 'a btree = | Empty | Node of 'a * 'a btree * 'a btree

This definition reads as follow: a binary tree containing values of type 'a (an arbitrary type) is either empty, or is a node containing one value of type 'a and two subtrees containing also values of type 'a, that is, two 'a btree.

Operations on binary trees are naturally expressed as recursive functions following the same structure as the type definition itself. For instance, here are functions performing lookup and insertion in ordered binary trees (elements increase from left to right):

# let rec member x btree =
#   match btree with
#     Empty -> false
#   | Node(y, left, right) ->
#       if x = y then true else
#       if x < y then member x left else member x right;;
val member : 'a -> 'a btree -> bool = <fun>
# let rec insert x btree =
#   match btree with
#     Empty -> Node(x, Empty, Empty)
#   | Node(y, left, right) ->
#       if x <= y then Node(y, insert x left, right)
#                 else Node(y, left, insert x right);;
val insert : 'a -> 'a btree -> 'a btree = <fun>

Imperative features

Though all examples so far were written in purely applicative style, Caml is also equipped with full imperative features. This includes the usual while and for loops, as well as mutable data structures such as arrays. Arrays are either given in extension between [| and |] brackets, or allocated and initialized with the Array.create function, then filled up later by assignments. For instance, the function below sums two vectors (represented as float arrays) componentwise.

# let add_vect v1 v2 =
#   let len = min (Array.length v1) (Array.length v2) in
#   let res = Array.create len 0.0 in
#   for i = 0 to len - 1 do
#     res.(i) <- v1.(i) +. v2.(i)
#   done;
#   res;;
val add_vect : float array -> float array -> float array = <fun>
# add_vect [| 1.0; 2.0 |] [| 3.0; 4.0 |];;
- : float array = [|4; 6|]

Record fields can also be modified by assignment, provided they are declared mutable in the definition of the record type:

# type mutable_point = { mutable x: float; mutable y: float };;
type mutable_point = { mutable x: float; mutable y: float }
# let translate p dx dy =
#   p.x <- p.x +. dx; p.y <- p.y +. dy;;
val translate : mutable_point -> float -> float -> unit = <fun>
# let mypoint = { x = 0.0; y = 0.0 };;
val mypoint : mutable_point = {x=0; y=0}
# translate mypoint 1.0 2.0;;
- : unit = ()
# mypoint;;
- : mutable_point = {x=1; y=2}

Caml has no built-in notion of variable – identifiers whose current value can be changed by assignment. (The let binding is not an assignment, it introduces a new identifier with a new scope.) However, the standard library provides references, which are mutable indirection cells (or one-element arrays), with operators ! to fetch the current contents of the reference and := to assign the contents. Variables can then be emulated by let-binding a reference. For instance, here is an in-place insertion sort over arrays:

# let insertion_sort a =
#   for i = 1 to Array.length a - 1 do
#     let val_i = a.(i) in
#     let j = ref i in

#     while !j > 0 && val_i < a.(!j - 1) do
#       a.(!j) <- a.(!j - 1);
#       j := !j - 1
#     done;
#     a.(!j) <- val_i
#   done;;
val insertion_sort : 'a array -> unit = <fun>

References are also useful to write functions that maintain a current state between two calls to the function. For instance, the following pseudo-random number generator keeps the last returned number in a reference:

# let current_rand = ref 0;;
val current_rand : int ref = {contents=0}
# let random () =
#   current_rand := !current_rand * 25713 + 1345;
#   !current_rand;;
val random : unit -> int = <fun>

Again, there is nothing magic with references: they are implemented as a one-field mutable record, as follows.

# type 'a ref = { mutable contents: 'a };;
type 'a ref = { mutable contents: 'a }
# let (!) r = r.contents;;
val ! : 'a ref -> 'a = <fun>
# let (:=) r newval = r.contents <- newval;;
val := : 'a ref -> 'a -> unit = <fun>

Exceptions

Caml provides exceptions for signalling and handling exceptional conditions. Exceptions can also be used as a general-purpose non-local control structure. Exceptions are declared with the exception construct, and signalled with the raise operator. For instance, the function below for taking the head of a list uses an exception to signal the case where an empty list is given.

# exception Empty_list;;
exception Empty_list
# let head l =
#   match l with
#     [] -> raise Empty_list
#   | hd :: tl -> hd;;
val head : 'a list -> 'a = <fun>
# head [1;2];;
- : int = 1

# head [];;
Uncaught exception: Empty_list

Exceptions are used throughout the standard library to signal cases where the library functions cannot complete normally. For instance, the List.assoc function, which returns the data associated with a given key in a list of (key, data) pairs, raises the predefined exception Not_found when the key does not appear in the list:

# List.assoc 1 [(0, "zero"); (1, "one")];;
- : string = "one"
# List.assoc 2 [(0, "zero"); (1, "one")];;
Uncaught exception: Not_found

Exceptions can be trapped with the try… with construct:

# let name_of_binary_digit digit =
#   try
#     List.assoc digit [0, "zero"; 1, "one"]
#   with Not_found ->
#     "not a binary digit";;
val name_of_binary_digit : int -> string = <fun>
# name_of_binary_digit 0;;
- : string = "zero"
# name_of_binary_digit (-1);;
- : string = "not a binary digit"

The with part is actually a regular pattern-matching on the exception value. Thus, several exceptions can be caught by one try… with construct. Also, finalization can be performed by trapping all exceptions, performing the finalization, then raising again the exception:

# let temporarily_set_reference ref newval funct =
#   let oldval = !ref in
#   try
#     ref := newval;
#     let res = funct () in
#     ref := oldval;
#     res
#   with x ->
#     ref := oldval;
#     raise x;;
val temporarily_set_reference : 'a ref -> 'a -> (unit -> 'b) -> 'b = <fun>

Symbolic processing of expressions

We finish this introduction with a more complete example representative of the use of Caml for symbolic processing: formal manipulations of arithmetic expressions containing variables. The following variant type describes the expressions we shall manipulate:

# type expression =
#     Const of float
#   | Var of string
#   | Sum of expression * expression    (* e1 + e2 *)
#   | Diff of expression * expression   (* e1 - e2 *)
#   | Prod of expression * expression   (* e1 * e2 *)
#   | Quot of expression * expression   (* e1 / e2 *)
# ;;
type expression =
  | Const of float
  | Var of string
  | Sum of expression * expression
  | Diff of expression * expression
  | Prod of expression * expression
  | Quot of expression * expression

We first define a function to evaluate an expression given an environment that maps variable names to their values. For simplicity, the environment is represented as an association list.

# exception Unbound_variable of string;;
exception Unbound_variable of string
# let rec eval env exp =
#   match exp with
#     Const c -> c
#   | Var v ->
#       (try List.assoc v env with Not_found -> raise(Unbound_variable v))
#   | Sum(f, g) -> eval env f +. eval env g
#   | Diff(f, g) -> eval env f -. eval env g
#   | Prod(f, g) -> eval env f *. eval env g
#   | Quot(f, g) -> eval env f /. eval env g;;
val eval : (string * float) list -> expression -> float = <fun>
# eval [("x", 1.0); ("y", 3.14)] (Prod(Sum(Var "x", Const 2.0), Var "y"));;
- : float = 9.42

Now for a real symbolic processing, we define the derivative of an expression with respect to a variable dv:

# let rec deriv exp dv =
#   match exp with
#     Const c -> Const 0.0
#   | Var v -> if v = dv then Const 1.0 else Const 0.0
#   | Sum(f, g) -> Sum(deriv f dv, deriv g dv)
#   | Diff(f, g) -> Diff(deriv f dv, deriv g dv)
#   | Prod(f, g) -> Sum(Prod(f, deriv g dv), Prod(deriv f dv, g))
#   | Quot(f, g) -> Quot(Diff(Prod(deriv f dv, g), Prod(f, deriv g dv)),
#                        Prod(g, g))

# ;;
val deriv : expression -> string -> expression = <fun>
# deriv (Quot(Const 1.0, Var "x")) "x";;
- : expression =
Quot
 (Diff (Prod (Const 0, Var "x"), Prod (Const 1, Const 1)),
  Prod (Var "x", Var "x"))

Pretty-printing and parsing

As shown in the examples above, the internal representation (also called abstract syntax) of expressions quickly becomes hard to read and write as the expressions get larger. We need a printer and a parser to go back and forth between the abstract syntax and the concrete syntax, which in the case of expressions is the familiar algebraic notation (e.g. 2*x+1).

For the printing function, we take into account the usual precedence rules (i.e. * binds tighter than +) to avoid printing unnecessary parentheses. To this end, we maintain the current operator precedence and print parentheses around an operator only if its precedence is less than the current precedence.

# let print_expr exp =
#   (* Local function definitions *)
#   let open_paren prec op_prec =
#     if prec > op_prec then print_string "(" in
#   let close_paren prec op_prec =
#     if prec > op_prec then print_string ")" in
#   let rec print prec exp =     (* prec is the current precedence *)
#     match exp with
#       Const c -> print_float c
#     | Var v -> print_string v
#     | Sum(f, g) ->
#         open_paren prec 0;
#         print 0 f; print_string " + "; print 0 g;
#         close_paren prec 0
#     | Diff(f, g) ->
#         open_paren prec 0;
#         print 0 f; print_string " - "; print 1 g;
#         close_paren prec 0
#     | Prod(f, g) ->
#         open_paren prec 2;
#         print 2 f; print_string " * "; print 2 g;
#         close_paren prec 2
#     | Quot(f, g) ->
#         open_paren prec 2;
#         print 2 f; print_string " / "; print 3 g;
#         close_paren prec 2

#   in print 0 exp;;
val print_expr : expression -> unit = <fun>
# let e = Sum(Prod(Const 2.0, Var "x"), Const 1.0);;
val e : expression = Sum (Prod (Const 2, Var "x"), Const 1)
# print_expr e; print_newline();;
2 * x + 1
- : unit = ()
# print_expr (deriv e "x"); print_newline();;
2 * 1 + 0 * x + 0
- : unit = ()

Parsing (transforming concrete syntax into abstract syntax) is usually more delicate. Caml offers several tools to help write parsers: on the one hand, Caml versions of the lexer generator Lex and the parser generator Yacc (see chapter 10), which handle LALR(1) languages using pushdown automata; on the other hand, a predefined type of streams (of characters or tokens) and pattern-matching over streams, which facilitate the writing of recursive-descent parsers for LL(1) languages. An example using ocamllex and ocamlyacc is given in chapter 10. Here, we will use stream parsers.

# open Genlex;;
# let lexer = make_lexer ["("; ")"; "+"; "-"; "*"; "/"];;
val lexer : char Stream.t -> Genlex.token Stream.t = <fun>

For the lexical analysis phase (transformation of the input text into a stream of tokens), we use a “generic” lexer provided in the standard library module Genlex. The make_lexer function takes a list of keywords and returns a lexing function that “tokenizes” an input stream of characters. Tokens are either identifiers, keywords, or literals (integer, floats, characters, strings). Whitespace and comments are skipped.

# let token_stream = lexer(Stream.of_string "1.0 +x");;
val token_stream : Genlex.token Stream.t = <abstr>
# Stream.next token_stream;;
- : Genlex.token = Float 1
# Stream.next token_stream;;
- : Genlex.token = Kwd "+"
# Stream.next token_stream;;
- : Genlex.token = Ident "x"

The parser itself operates by pattern-matching on the stream of tokens. As usual with recursive descent parsers, we use several intermediate parsing functions to reflect the precedence and associativity of operators. Pattern-matching over streams is more powerful than on regular data structures, as it allows recursive calls to parsing functions inside the patterns, for matching sub-components of the input stream. See chapter 5 for more details.

# let rec parse_expr = parser
#     [< e1 = parse_mult; e = parse_more_adds e1 >] -> e
# and parse_more_adds e1 = parser
#     [< 'Kwd "+"; e2 = parse_mult; e = parse_more_adds (Sum(e1, e2)) >] -> e
#   | [< 'Kwd "-"; e2 = parse_mult; e = parse_more_adds (Diff(e1, e2)) >] -> e
#   | [< >] -> e1
# and parse_mult = parser
#     [< e1 = parse_simple; e = parse_more_mults e1 >] -> e
# and parse_more_mults e1 = parser
#     [< 'Kwd "*"; e2 = parse_simple; e = parse_more_mults (Prod(e1, e2)) >] -> e
#   | [< 'Kwd "/"; e2 = parse_simple; e = parse_more_mults (Quot(e1, e2)) >] -> e
#   | [< >] -> e1
# and parse_simple = parser
#     [< 'Ident s >] -> Var s
#   | [< 'Int i >] -> Const(float i)
#   | [< 'Float f >] -> Const f
#   | [< 'Kwd "("; e = parse_expr; 'Kwd ")" >] -> e;;
val parse_expr : Genlex.token Stream.t -> expression = <fun>
val parse_more_adds : expression -> Genlex.token Stream.t -> expression =
  <fun>
val parse_mult : Genlex.token Stream.t -> expression = <fun>
val parse_more_mults : expression -> Genlex.token Stream.t -> expression =
  <fun>
val parse_simple : Genlex.token Stream.t -> expression = <fun>

Composing the lexer and parser, we finally obtain a function to read an expression from a character string:

# let read_expr s = parse_expr(lexer(Stream.of_string s));;
val read_expr : string -> expression = <fun>
# read_expr "2*(x+y)";;
- : expression = Prod (Const 2, Sum (Var "x", Var "y"))

Standalone Caml programs

All examples given so far were executed under the interactive system. Caml code can also be compiled separately and executed non-interactively using the batch compilers ocamlc or ocamlopt. The source code must be put in a file with extension .ml. It consists of a sequence of phrases, which will be evaluated at runtime in their order of appearance in the source file. Unlike in interactive mode, types and values are not printed automatically; the program must call printing functions explicitly to produce some output. Here is a sample standalone program to print Fibonacci numbers:

(* File fib.ml *)
let rec fib n =
  if n < 2 then 1 else fib(n-1) + fib(n-2);;

let main () =
  let arg = int_of_string Sys.argv.(1) in
  print_int(fib arg);
  print_newline();
  exit 0;;
main ();;

Sys.argv is an array of strings containing the command-line parameters. Sys.argv.(1) is thus the first command-line parameter. The program above is compiled and executed with the following shell commands:

$ ocamlc -o fib fib.ml
$ ./fib 10
89
$ ./fib 20
10946

22

Objects in Caml

(Chapter written by Jérôme Vouillon and Didier Rémy)

This chapter gives an overview of the object-oriented features of Objective Caml.

Classes and objects

The class point has one instance variable x and two methods get_x and move. The initial value of the instance variable is given here by the class parameter x_init. The variable x is declared mutable, so the method move can change its value.

# class point x_init =
#   val mutable x = x_init
#   method get_x = x
#   method move d = x <- x + d
# end;;
class point (int) =
  val mutable x : int
  method get_x : int
  method move : int -> unit
end

We now create a new point p, giving the initialization argument 7.

# let p = new point 7;;
val p : point = <obj>

Note that the type of p is point. This is an abbreviation automatically defined by the class definition above. It stands for the object type <get_x : int; move : int -> unit>, listing the methods of class point along with their types.

Let us apply some methods to p:

# p#get_x;;
- : int = 7
# p#move 3;;

- : unit = ()
# p#get_x;;
- : int = 10

The library function Oo.copy makes a shallow copy of an object. Its type is < .. > as 'a -> 'a (which is parsed as (< .. > as 'a) -> 'a). The keyword as in that type binds the type variable 'a to the object type < .. >. Therefore, Oo.copy takes an object with any methods (represented by the ellipsis), and returns an object of the same type. The type of Oo.copy is different from type < .. > -> < .. > as each ellipsis represents a different set of methods. Ellipsis actually behaves as a type variable.

# let q = Oo.copy p;;
val q : point = <obj>
# q#move 7; (p#get_x, q#get_x);;
- : int * int = 10, 17

Objects can be compared using the generic comparison functions (=, <, ...). Two objects are equal if and only if they are physically equal. In particular, an object and its copy are not equal.

# let q = Oo.copy p;;
val q : point = <obj>
# p = q, p = p;;
- : bool * bool = false, true

Inheritance

We now define a new class colored_point. This class inherits from class point. So, it has all the instance variable and all the methods of point, plus a new instance variable c and a new method

color.
# class colored_point x (c : string) =
#   inherit point x
#   val c = c
#   method color = c
# end;;
class colored_point (int) (string) =
  val c : string
  val mutable x : int
  method color : string
  method get_x : int
  method move : int -> unit
end
# let p' = new colored_point 5 "red";;
val p' : colored_point = <obj>
# p'#get_x, p'#color;;
- : int * string = 5, "red"

A point and a colored point have incompatible types: a point has no method color. However, the function get_x below is a generic function applying method get_x to any object p that has this method (and possibly some others, which are represented by an ellipsis in the type). Thus, it applies to both points and colored points.

# let get_succ_x p = p#get_x + 1;;
val get_succ_x : < get_x : int; .. > -> int = <fun>
# get_succ_x p + get_succ_x p';;
- : int = 17

Methods need not be declared previously, as shown by the example:

# let set_x p = p#set_x;;
val set_x : < set_x : 'a; .. > -> 'a = <fun>
# let incr p = set_x p (get_succ_x p);;
val incr : < get_x : int; set_x : int -> 'a; .. > -> 'a = <fun>

Parameterized classes

Reference cells can also be implemented as objects. The naive definition fails to typecheck:

# class ref x_init =
#   val mutable x = x_init
#   method get = x
#   method set y = x <- y
# end;;
Characters 5-85:
The type variable 'a is not bound in implicit type definition
  ref = < get : 'a; set : 'a -> unit >
It should be captured by a class type parameter

The reason is that at least one of the methods has a polymorphic type (here, the type of the value stored in the reference cell), thus the class should be parametric. A monomorphic instance of the class could be defined by:

# class ref (x_init:int) =
#   val mutable x = x_init
#   method get = x
#   method set y = x <- y
# end;;
class ref (int) =
  val mutable x : int
  method get : int
  method set : int -> unit
end

A class for polymorphic references must explicitly list the type parameters in its declaration. The type parameters must also be bound somewhere in the class body by a type constraint.

# class 'a ref x_init =
#   val mutable x = (x_init : 'a)
#   method get = x
#   method set y = x <- y
# end;;
class 'a ref ('a) =
  val mutable x : 'a
  method get : 'a
  method set : 'a -> unit
end
# let r = new ref 1 in r#set 2; (r#get);;
- : int = 2

The type parameter in the declaration may actually be constrained in the body of the class definition. In the class type, the actual value of the type parameter is displayed in the constraint clause.

# class 'a ref (x_init:'a) =
#   val mutable x = x_init + 1
#   method get = x
#   method set y = x <- y
# end;;
class 'a ref ('a) =
  constraint 'a = int
  val mutable x : int
  method get : int
  method set : int -> unit
end

Let us consider a more realistic example. We put an additional type constraint in method move, since no free variables must remain uncaptured by a type parameter.

# class 'a circle (c : 'a) =
#   val mutable center = c
#   method center = center
#   method set_center c = center <- c
#   method move = (center#move : int -> unit)
# end;;
class 'a circle ('a) =
  constraint 'a = < move : int -> unit; .. >
  val mutable center : 'a
  method center : 'a
  method move : int -> unit
  method set_center : 'a -> unit
end

An alternate definition of circle, using a constraint clause in the class definition, is shown below. The type #point used below in the constraint clause is an abbreviation produced by the definition of class point. This abbreviation unifies with the type of any object belonging to a subclass of class point. It actually expands to < get_x : int; move : int -> unit; .. >. This leads to the following alternate definition of circle, which has slightly stronger constraints on its argument, as we now expect center to have a method get_x.

# class 'a circle (c : 'a) =
#   constraint 'a = #point
#   val mutable center = c
#   method center = center
#   method set_center c = center <- c
#   method move = center#move
# end;;
class 'a circle ('a) =
  constraint 'a = #point
  val mutable center : 'a
  method center : 'a
  method move : int -> unit
  method set_center : 'a -> unit
end

The class colored_circle is a specialized version of class circle which requires the type of the center to unify with #colored_point, and adds a method color.

# class 'a colored_circle c =
#   constraint 'a = #colored_point
#   inherit ('a) circle c
#   method color = center#color
# end;;
class 'a colored_circle ('a) =
  constraint 'a = #colored_point
  val mutable center : 'a
  method center : 'a
  method color : string
  method move : int -> unit
  method set_center : 'a -> unit
end

Reference to self

A method can also send messages to the object that invoked the method. For that, self must be explicitly bound, here to the variable s.

# class printable_point y as s =
#   inherit point y
#   method print = print_int s#get_x
# end;;

class printable_point (int) =
  val mutable x : int
  method get_x : int
  method move : int -> unit
  method print : unit
end
# let p = new printable_point 7;;
val p : printable_point = <obj>
# p#print;;
7- : unit = ()

The variable s is bound at the invocation of a method. In particular, if the class printable_point is inherited, the variable s will correctly be bound to an object of the subclass.

Multiple inheritance

Multiple inheritance is allowed. Only the last definition of a method (or of an instance variable) is kept. Previous definitions of a method can be reused by binding the related ancestor. Below, super is bound to the ancestor printable_point. The name super is not actually a variable and can only be used to select a method as in super#print.

# class printable_colored_point y c as self =
#   inherit colored_point y c
#   inherit printable_point y as super
#   method print =
#     print_string "(";
#     super#print;
#     print_string ", ";
#     print_string (self#color);
#     print_string ")"
# end;;
class printable_colored_point (int) (string) =
  val c : string
  val mutable x : int
  method color : string
  method get_x : int
  method move : int -> unit
  method print : unit
end
# let p' = new printable_colored_point 7 "red";;
val p' : printable_colored_point = <obj>
# p'#print;;
(7, red)- : unit = ()

Non-mutable objects

It is possible to write a version of class point without assignments on the instance variables. The construct {< ... >} returns a copy of “self” (that is, the current object), possibly changing the value of some instance variables.

# class functional_point y =
#   val x = y
#   method get_x = x
#   method move d = {< x = x + d >}
# end;;
class functional_point (int) : 'a =
  val x : int
  method get_x : int
  method move : int -> 'a
end
# let p = new functional_point 7;;
val p : functional_point = <obj>
# p#get_x;;
- : int = 7
# (p#move 3)#get_x;;
- : int = 10
# p#get_x;;
- : int = 7

Note that the type abbreviation functional_point is recursive, which can be seen in the class type of functional_point: the type of self to 'a and 'a appears inside the type of the move method.

Virtual methods

The class comparable below is a template for classes with a binary method leq of type 'a -> bool where the type variable 'a is bound to the type of self. Since this class has a method declared but not defined, it must be flagged virtual and cannot be instantiated (that is, no object of this class can be created). It still defines abbreviations. In particular, #comparable expands to < leq : 'a -> bool; .. > as 'a. We see here that the binder as also allows to write recursive types.

# class virtual comparable () : 'a =
#   virtual leq : 'a -> bool
# end;;
class virtual comparable (unit) : 'a = virtual leq : 'a -> bool end

We then define a subclass of comparable that wraps integers as comparable objects. There is a type constraint on the class parameter x as the primitive <= is a polymorphic comparison function in Objective Caml. The inherit clause ensures that the type of objects of this class is an instance

of #comparable.

# class int_comparable (x : int) =
#   inherit comparable ()
#   val x = x
#   method x = x
#   method leq p = x <= p#x
# end;;
class int_comparable (int) : 'a =
  val x : int
  method leq : 'a -> bool
  method x : int
end

Objects of class int_comparable2 below can also modify the integer they hold. The status of instance variable x is changed. It is now mutable. Note that the type int_comparable2 is not a subtype of type int_comparable, as the self type appears in contravariant position in the type of method leq.

# class int_comparable2 x =
#   inherit int_comparable x
#   val mutable x
#   method set_x y = x <- y
# end;;
class int_comparable2 (int) : 'a =
  val mutable x : int
  method leq : 'a -> bool
  method set_x : int -> unit
  method x : int
end

The function min will return the minimum of any two objects whose type unifies with

#comparable. The type of min is not the same as #comparable  ->  #comparable  ->

#comparable, as the abbreviation #comparable hides a type variable (an ellipsis). Each occurrence of this abbreviation generates a new variable.

# let min (x : #comparable) y =
#   if x#leq y then x else y;;
val min : (#comparable as 'a) -> 'a -> 'a = <fun>

This function can be applied to objects of type int_comparable or int_comparable2.

# (min (new int_comparable  7) (new int_comparable 11))#x;;
- : int = 7
# (min (new int_comparable2 5) (new int_comparable2 3))#x;;
- : int = 3

Protected methods

Protected methods are methods that do not appear in object interfaces. They can only be invoked from other methods of the same object.

# class restricted_point x_init as self =
#   val mutable x = x_init
#   method get_x = x
#   method protected move d = x <- x + d
#   method bump = self#move 1
# end;;
class restricted_point (int) =
  val mutable x : int
  method bump : unit
  method get_x : int
  method protected move : int -> unit
end
# let p = new restricted_point 0;;
val p : restricted_point = <obj>
# p#move 10;;
Characters 0-1:
This expression has no method move
# p#bump;;
- : unit = ()

Protected methods are inherited. They can be hidden by signature matching, as described in the next section.

Class interfaces

Class interfaces are inferred from class definitions. They may also be defined directly in interfaces of modules. For instance, the following is the interface of a module defining class restricted_point.

# module type POINT = sig
#   class restricted_point (int) =
#     val mutable x : int
#     method get_x : int
#     method protected move : int -> unit
#     method bump : unit
#  end
# end;;
module type POINT =
  sig
    class restricted_point (int) =
      val mutable x : int
      method bump : unit

      method get_x : int
      method protected move : int -> unit
    end
  end
# module Point : POINT = struct
#   class restricted_point x = inherit restricted_point x end
# end;;
module Point : POINT

It is sometime necessary to restrict the interface of classes. Instance variables and protected methods can be hidden by signature matching, as shown below. However, public methods cannot be hidden.

# module type ABSPOINT = sig
#   class restricted_point (int) =
#     method get_x : int
#     method bump : unit
#  end
# end;;
module type ABSPOINT =
  sig
    class restricted_point (int) = method bump : unit method get_x : int end
  end
# module Abspoint : ABSPOINT = Point;;
module Abspoint : ABSPOINT

Using coercions

Subtyping is never implicit. There are, however, two ways to perform subtyping. The most general construction is fully explicit: both the domain and the codomain of the type coercion must be given.

We have seen that points and colored points have incompatible types. For instance, they cannot be mixed in the same list. However, a colored point can be coerced to a point, hiding its color method:

# let colored_point_to_point cp = (cp : colored_point :> point);;
val colored_point_to_point : colored_point -> point = <fun>
# let p = new point 3 and q = new colored_point 4 "blue";;
val p : point = <obj>
val q : colored_point = <obj>
# let l = [p; (colored_point_to_point q)];;
val l : point list = [<obj>; <obj>]

An object of type t can be seen as an object of type t' only if t is a subtype of t'. For instance, a point cannot be seen as a colored point.

# (p : point :> colored_point);;
Characters 0-28:
Type point = < get_x : int; move : int -> unit > is not a subtype of type
  colored_point = < get_x : int; move : int -> unit; color : string >

Indeed, backward coercions are unsafe, and should be combined with a type case, possibly raising a runtime error. However, there is not such operation available in the language.

Be aware that subtyping and inheritance are not related. Inheritance is a syntactic relation between classes while subtyping is a semantic relation between types. For instance, the class of colored points could have been defined directly, without inheriting from the class of points; the type of colored points would remain unchanged and thus still be a subtype of points. Conversely, the class int_comparable inherits from class comparable, but type int_comparable is not a subtype

of comparable.
# function x -> (x : int_comparable :> comparable);;
Characters 14-48:
Type int_comparable = < leq : int_comparable -> bool; x : int >
is not a subtype of type comparable = < leq : comparable -> bool >
Type int_comparable -> bool is not a subtype of type comparable -> bool
Type comparable = < leq : comparable -> bool > is not a subtype of type
  int_comparable = < leq : int_comparable -> bool; x : int >

Indeed, an object p of class int_comparable has a method leq that expects an argument of type int_comparable since it accesses its x method. Considering p of type comparable would allow to call method leq on p with an argument that does not have a method x, which would be an error.

The domain of a coercion can usually be omitted. For instance, one can define:

# let to_point cp = (cp :> point);;
val to_point : < get_x : int; move : int -> unit; .. > -> point = <fun>

In this case, the function colored_point_to_point is an instance of the function to_point. This is not always true, however. The fully explicit coercion is more precise and is sometimes unavoidable. Here is an example:

# class virtual c () = virtual m : c end;;
class virtual c (unit) = virtual m : c end
# class c' () as self =
#   inherit c ()
#   method m = (self :> c)
#   method m' = 1
# end;;
Characters 51-55:
This expression cannot be coerced to type c = < m : c >; it has type
  < m : c; m' : 'a; .. >
but is here used with type < m : 'b; m' : 'a; .. > as 'b
Type c = < m : c > is not compatible with type 'b

The type of the coercion to type c can be seen here:

# function x -> (x :> c);;
- : (< m : 'a; .. > as 'a) -> c = <fun>

As class c' inherits from class c, its method m must have type c. On the other hand, in expression (self :> c) the type of self and the domain of the coercion above must be unified. That is, the type of the method m in self (i.e. c) is also the type of self. So, the type of self is c. This is a contradiction, as the type of self has a method m', whereas type c does not.

The desired coercion of type <m : c;..> -> c can be obtained by using a fully explicit coercion:

# function x -> (x : #c :> c);;
- : #c -> c = <fun>

Thus one can define class c’ as follows:

# class c' () as self =
#   inherit c ()
#   method m = (self : #c :> c)
#   method m' = 1
# end;;
class c' (unit) = method m : c method m' : int end

An alternative is to define class c as follows (of course this definition is not equivalent to the previous one):

# class virtual c () : 'a = virtual m : 'a end;;
class virtual c (unit) : 'a = virtual m : 'a end

Then, a coercion operator is not even required.

# class c' () as self =
#   inherit c ()
#   method m = self
#   method m' = 1
# end;;
class c' (unit) : 'a = method m : 'a method m' : int end

Here, the simple coercion operator (e :> c) can be used to coerce an object expression e from type c' to type c. Semi implicit coercions are actually defined so as to work correctly with classes returning self.

# (new c' () :> c);;
- : c = <obj>

Another common problem may occur when one tries to define a coercion to a class c inside the definition of class c. The problem is due to the type abbreviation not being completely defined yet, and so its subtypes are not clearly known. Then, a coercion (_: #c :> c) is taken to be the identity function, as in

# function x -> (x :> 'a);;
- : 'a -> 'a = <fun>

As a consequence, if the coercion is applied to self, as in the following example, the type of self is unified with the closed type c. This constrains the class to be closed:

# class c () as self = method m = (self : #c :> c) end;;
Characters 32-48:
Type #c = < m : 'a; .. > is not a subtype of type c = < m : 'b; .. >

Although declaring the class as closed will typecheck, this is not usually what is desired.

# class closed c () as self = method m = (self : #c :> c) end;;
Characters 39-55:
Type < m : 'b; .. > as 'a is not a subtype of type c = 'a

This problem can sometimes be avoided by first defining the abbreviation, using a virtual class:

# class virtual c0 () = virtual m : c0 end;;
class virtual c0 (unit) = virtual m : c0 end
# class c () as self = method m = (self : #c0 :> c0) end;;
class c (unit) = method m : c0 end

The class c may be also declared to inherit from the virtual class c0, so as to simultaneously enforce all methods of c to have the same type as the methods of c0.

# class c () as self = inherit c0 () method m = (self : #c0 :> c0) end;;
class c (unit) = method m : c0 end

One could think of defining the type abbreviation directly:

# type c1 = <m : c1>;;
type c1 = < m : c1 >

However, the abbreviation #c1 cannot be defined this way (the abbreviation #c0 is defined from the class c0, not from the type c0), and should be expanded:

# class c () as self =  method m = (self : <m : c1; ..> as 'a :> c1) end;;
class c (unit) = method m : c1 end

Recursive classes

Recursive classes can be used to define objects whose types are mutually recursive.

# class window () =
#   val mutable top_widget = (None : widget option)
#   method top_widget = top_widget
# and widget (w : window) =
#   val window = w
#   method window = window
# end;;
class window (unit) =

  val mutable top_widget : widget option
  method top_widget : widget option
end
class widget (window) = val window : window method window : window end

Although their types are mutually recursive, the classes widget and window are themselves independent.

Simple modules as classes

There is sometime an alternative between using modules or classes. Indeed, there are situations when the two approaches are quite similar. For instance, a stack can be straightforwardly implemented as a class:

# exception Empty;;
exception Empty
# class 'a stack () =
#   val mutable l = ([] : 'a list)
#   method push x = l <- x::l
#   method pop = match l with [] -> raise Empty | a::l' -> l <- l'; a
#   method clear = l <- []
#   method length = List.length l
# end;;
class 'a stack (unit) =
  val mutable l : 'a list
  method clear : unit
  method length : int
  method pop : 'a
  method push : 'a -> unit
end

However, writing a method for iterating over a stack is more problematic. A method fold would have type ('b -> 'a -> 'b) -> 'b -> 'b. Here 'a is the parameter of the stack. The parameter 'b is not related to the class 'a stack but to the argument that will be passed to the method fold. The intuition is that method fold should be polymorphic, i.e. of type All ('a) ('b -> 'a -> 'b) -> 'b -> 'b, which is not currently possible. One possibility would be to make 'b an extra parameter of class stack

# class ('a, 'b) stack2 () =
#   inherit ('a) stack ()
#   method fold f (x : 'b) = List.fold_left f x l
# end;;
class ('a, 'b) stack2 (unit) =
  val mutable l : 'a list
  method clear : unit
  method fold : ('b -> 'a -> 'b) -> 'b -> 'b
  method length : int

  method pop : 'a
  method push : 'a -> unit
end

However, method fold of a given object can only be applied to functions that all have the same type:

# let s = new stack2 ();;
val s : ('_a, '_b) stack2 = <obj>
# s#fold (+) 0;;
- : int = 0
# s;;
- : (int, int) stack2 = <obj>

The best solution would be to make method fold polymorphic. However Ocaml does not currently allow methods to be polymorphic.

Thus, the current solution is to leave the function fold outside of the class.

# class 'a stack3 () =
#   inherit ('a) stack ()
#   method iter f = List.iter (f : 'a -> unit) l
# end;;
class 'a stack3 (unit) =
  val mutable l : 'a list
  method clear : unit
  method iter : ('a -> unit) -> unit
  method length : int
  method pop : 'a
  method push : 'a -> unit
end
# let stack_fold (s : 'a #stack3) f x =
#   let accu = ref x in
#   s#iter (fun e -> accu := f !accu e);
#   !accu;;
val stack_fold : 'a #stack3 -> ('b -> 'a -> 'b) -> 'b -> 'b = <fun>

Implementing sets leads to another difficulty. Indeed, the method union needs to be able to access the internal representation of another object of the same class. For that, a set class must have an additional method returning this representation. However, this representation should not be public. This result is obtained by making the type of the representation abstract via a module signature constraint. From outside, the additional method appears like a tag ensuring that an object belongs to class set.

# module type SET =
#   sig
#     type 'a tag

#     class 'a c (unit) : 'b =
#       method is_empty : bool
#       method mem : 'a -> bool
#       method add : 'a -> 'b
#       method union : 'b -> 'b
#       method iter : ('a -> unit) -> unit
#       method tag : 'a tag
#     end
#   end;;
module type SET =
  sig
    type 'a tag
    class 'a c (unit) : 'b =
      method add : 'a -> 'b
      method is_empty : bool
      method iter : ('a -> unit) -> unit
      method mem : 'a -> bool
      method tag : 'a tag
      method union : 'b -> 'b
    end
  end
# module Set : SET =
#   struct
#     let rec merge l1 l2 =
#       match l1 with
#         [] -> l2
#       | h1 :: t1 ->
#           match l2 with
#             [] -> l1
#           | h2 :: t2 ->
#               if h1 < h2 then h1 :: merge t1 l2
#               else if h1 > h2 then h2 :: merge l1 t2
#               else merge t1 l2
#     type 'a tag = 'a list
#     class 'a c () : 'b =
#       val repr = ([] : 'a list)
#       method is_empty = (repr = [])
#       method mem x = List.exists ((=) x) repr
#       method add x = {< repr = merge [x] repr >}
#       method union (s : 'b) = {< repr = merge repr s#tag >}
#       method iter (f : 'a -> unit) = List.iter f repr
#       method tag = repr
#     end
#   end;;
module Set : SET

The module system

This chapter introduces the module system of Objective Caml.

Structures

A primary motivation for modules is to package together related definitions (such as the definitions of a data type and associated operations over that type) and enforce a consistent naming scheme for these definitions. This avoids running out of names or accidentally confusing names. Such a package is called a structure and is introduced by the struct… end construct, which contains an arbitrary sequence of definitions. The structure is usually given a name with the module binding. Here is for instance a structure packaging together a type of priority queues and their operations:

# module PrioQueue =
#   struct
#     type priority = int
#     type 'a queue = Empty | Node of priority * 'a * 'a queue * 'a queue
#     let empty = Empty
#     let rec insert queue prio elt =
#       match queue with
#         Empty -> Node(prio, elt, Empty, Empty)
#       | Node(p, e, left, right) ->
#           if prio <= p
#           then Node(prio, elt, insert right p e, left)
#           else Node(p, e, insert right prio elt, left)
#     exception Queue_is_empty
#     let rec remove_top = function
#         Empty -> raise Queue_is_empty
#       | Node(prio, elt, left, Empty) -> left
#       | Node(prio, elt, Empty, right) -> right
#       | Node(prio, elt, (Node(lprio, lelt, _, _) as left),
#                         (Node(rprio, relt, _, _) as right)) ->
#           if lprio <= rprio
#           then Node(lprio, lelt, remove_top left, right)
#           else Node(rprio, relt, left, remove_top right)

#     let extract = function
#         Empty -> raise Queue_is_empty
#       | Node(prio, elt, _, _) as queue -> (prio, elt, remove_top queue)
#   end;;
module PrioQueue :
  sig
    type priority = int
    type 'a queue = | Empty | Node of priority * 'a * 'a queue * 'a queue
    val empty : 'a queue
    val insert : 'a queue -> priority -> 'a -> 'a queue
    exception Queue_is_empty
    val remove_top : 'a queue -> 'a queue
    val extract : 'a queue -> priority * 'a * 'a queue
  end

Outside the structure, its components can be referred to using the “dot notation”, that is, identifiers qualified by a structure name. For instance, PrioQueue.insert in a value context is the function insert defined inside the structure PrioQueue. Similarly, PrioQueue.queue in a type context is the type queue defined in PrioQueue.

# PrioQueue.insert PrioQueue.empty 1 "hello";;
- : string PrioQueue.queue =
PrioQueue.Node (1, "hello", PrioQueue.Empty, PrioQueue.Empty)

Signatures

Signatures are interfaces for structures. A signature specifies which components of a structure are accessible from the outside, and with which type. It can be used to hide some components of a structure (e.g. local function definitions) or export some components with a restricted type. For instance, the signature below specifies the three priority queue operations empty, insert and extract, but not the auxiliary function remove_top. Similarly, it makes the queue type abstract (by not providing its actual representation as a concrete type).

# module type PRIOQUEUE =
#   sig
#     type priority = int         (* still concrete *)
#     type 'a queue               (* now abstract *)
#     val empty : 'a queue
#     val insert : 'a queue -> int -> 'a -> 'a queue
#     val extract : 'a queue -> int * 'a * 'a queue
#     exception Queue_is_empty
#   end;;
module type PRIOQUEUE =
  sig
    type priority = int
    type 'a queue

    val empty : 'a queue
    val insert : 'a queue -> int -> 'a -> 'a queue
    val extract : 'a queue -> int * 'a * 'a queue
    exception Queue_is_empty
  end

Restricting the PrioQueue structure by this signature results in another view of the PrioQueue structure where the remove_top function is not accessible and the actual representation of priority queues is hidden:

# module AbstractPrioQueue = (PrioQueue : PRIOQUEUE);;
module AbstractPrioQueue : PRIOQUEUE
# AbstractPrioQueue.remove_top;;
Characters 0-28:
Unbound value AbstractPrioQueue.remove_top
# AbstractPrioQueue.insert AbstractPrioQueue.empty 1 "hello";;
- : string AbstractPrioQueue.queue = <abstr>

The restriction can also be performed during the definition of the structure, as in

module PrioQueue = (struct ... end : PRIOQUEUE);;

An alternate syntax is provided for the above:

module PrioQueue : PRIOQUEUE = struct ... end;;

Functors

Functors are “functions” from structures to structures. They are used to express parameterized structures: a structure A parameterized by a structure B is simply a functor F with a formal parameter B (along with the expected signature for B) which returns the actual structure A itself. The functor F can then be applied to one or several implementations B_{1} … B_{n} of B, yielding the corresponding structures A_{1} … A_{n}.

For instance, here is a structure implementing sets as sorted lists, parameterized by a structure providing the type of the set elements and an ordering function over this type (used to keep the sets sorted):

# type comparison = Less | Equal | Greater;;
type comparison = | Less | Equal | Greater
# module type ORDERED_TYPE =
#   sig
#     type t
#     val cmp: t -> t -> comparison
#   end;;
module type ORDERED_TYPE = sig type t val cmp : t -> t -> comparison end
# module Set =

#   functor (Elt: ORDERED_TYPE) ->
#     struct
#       type element = Elt.t
#       type set = element list
#       let empty = []
#       let rec add x s =
#         match s with
#           [] -> [x]
#         | hd::tl ->
#            match Elt.cmp x hd with
#              Equal   -> s         (* x is already in s *)
#            | Less    -> x :: s    (* x is smaller than all elements of s *)
#            | Greater -> hd :: add x tl
#       let rec member x s =
#         match s with
#           [] -> false
#         | hd::tl ->
#             match Elt.cmp x hd with
#               Equal   -> true     (* x belongs to s *)
#             | Less    -> false    (* x is smaller than all elements of s *)
#             | Greater -> member x tl
#     end;;
module Set :
  functor(Elt : ORDERED_TYPE) ->
    sig
      type element = Elt.t
      type set = element list
      val empty : 'a list
      val add : Elt.t -> Elt.t list -> Elt.t list
      val member : Elt.t -> Elt.t list -> bool
    end

By applying the Set functor to a structure implementing an ordered type, we obtain set operations for this type:

# module OrderedString =
#   struct
#     type t = string
#     let cmp x y = if x = y then Equal else if x < y then Less else Greater
#   end;;
module OrderedString :
  sig type t = string val cmp : 'a -> 'a -> comparison end
# module StringSet = Set(OrderedString);;
module StringSet :
  sig
    type element = OrderedString.t

    type set = element list
    val empty : 'a list
    val add : OrderedString.t -> OrderedString.t list -> OrderedString.t list
    val member : OrderedString.t -> OrderedString.t list -> bool
  end
# StringSet.member "bar" (StringSet.add "foo" StringSet.empty);;
- : bool = false

Functors and type abstraction

As in the PrioQueue example, it would be good style to hide the actual implementation of the type set, so that users of the structure will not rely on sets being lists, and we can switch later to another, more efficient representation of sets without breaking their code. This can be achieved by restricting Set by a suitable functor signature:

# module type SETFUNCTOR =
#   functor (Elt: ORDERED_TYPE) ->
#     sig
#       type element = Elt.t      (* concrete *)
#       type set                  (* abstract *)
#       val empty : set
#       val add : element -> set -> set
#       val member : element -> set -> bool
#     end;;
module type SETFUNCTOR =
  functor(Elt : ORDERED_TYPE) ->
    sig
      type element = Elt.t
      type set
      val empty : set
      val add : element -> set -> set
      val member : element -> set -> bool
    end
# module AbstractSet = (Set : SETFUNCTOR);;
module AbstractSet : SETFUNCTOR
# module AbstractStringSet = AbstractSet(OrderedString);;
module AbstractStringSet :
  sig
    type element = OrderedString.t
    type set = AbstractSet(OrderedString).set
    val empty : set
    val add : element -> set -> set
    val member : element -> set -> bool
  end

# AbstractStringSet.add "gee" AbstractStringSet.empty;;
- : AbstractStringSet.set = <abstr>

In an attempt to write the type constraint above more elegantly, one may wish to name the signature of the structure returned by the functor, then use that signature in the constraint:

# module type SET =
#   sig
#     type element
#     type set
#     val empty : set
#     val add : element -> set -> set
#     val member : element -> set -> bool
#   end;;
module type SET =
  sig
    type element
    type set
    val empty : set
    val add : element -> set -> set
    val member : element -> set -> bool
  end
# module WrongSet = (Set : functor(Elt: ORDERED_TYPE) -> SET);;
module WrongSet : functor(Elt : ORDERED_TYPE) -> SET
# module WrongStringSet = WrongSet(OrderedString);;
module WrongStringSet :
  sig
    type element = WrongSet(OrderedString).element
    type set = WrongSet(OrderedString).set
    val empty : set
    val add : element -> set -> set
    val member : element -> set -> bool
  end
# WrongStringSet.add "gee" WrongStringSet.empty;;
Characters 19-24:
This expression has type string but is here used with type
  WrongStringSet.element = WrongSet(OrderedString).element

The problem here is that SET specifies the type element abstractly, so that the type equality between element in the result of the functor and t in its argument is forgotten. Consequently,

WrongStringSet.element is not the same type as string, and the operations of WrongStringSet

cannot be applied to strings. As demonstrated above, it is important that the type element in the signature SET be declared equal to Elt.t; unfortunately, this is impossible above since SET is defined in a context where Elt does not exist. To overcome this difficulty, Objective Caml provides a with type construct over signatures that allows to enrich a signature with extra type equalities:

# module AbstractSet =
#   (Set : functor(Elt: ORDERED_TYPE) -> (SET with type element = Elt.t));;
module AbstractSet :
  functor(Elt : ORDERED_TYPE) ->
    sig
      type element = Elt.t
      type set
      val empty : set
      val add : element -> set -> set
      val member : element -> set -> bool
    end

As in the case of simple structures, an alternate syntax is provided for defining functors and restricting their result:

module AbstractSet(Elt: ORDERED_TYPE) : (SET with type element = Elt.t) =
  struct ... end;;

Abstracting a type component in a functor result is a powerful technique that provides a high degree of type safety, as we now illustrate. Consider an ordering over character strings that is different from the standard ordering implemented in the OrderedString structure. For instance, we compare strings without distinguishing upper and lower case.

# module NoCaseString =
#   struct
#     type t = string
#     let cmp s1 s2 =
#       OrderedString.cmp (String.lowercase s1) (String.lowercase s2)
#   end;;
module NoCaseString :
  sig type t = string val cmp : string -> string -> comparison end
# module NoCaseStringSet = AbstractSet(NoCaseString);;
module NoCaseStringSet :
  sig
    type element = NoCaseString.t
    type set = AbstractSet(NoCaseString).set
    val empty : set
    val add : element -> set -> set
    val member : element -> set -> bool
  end
# NoCaseStringSet.add "FOO" AbstractStringSet.empty;;
Characters 26-49:
This expression has type
  AbstractStringSet.set = AbstractSet(OrderedString).set
but is here used with type
  NoCaseStringSet.set = AbstractSet(NoCaseString).set

Notice that the two types AbstractStringSet.set and NoCaseStringSet.set are not compatible, and values of these two types do not match. This is the correct behavior: even though both set types contain elements of the same type (strings), both are built upon different orderings of that type, and different invariants need to be maintained by the operations (being strictly increasing for the standard ordering and for the case-insensitive ordering). Applying operations from AbstractStringSet to values of type NoCaseStringSet.set could give incorrect results, or build lists that violate the invariants of NoCaseStringSet.

Modules and separate compilation

All examples of modules so far have been given in the context of the interactive system. However, modules are most useful for large, batch-compiled programs. For these programs, it is a practical necessity to split the source into several files, called compilation units, that can be compiled separately, thus minimizing recompilation after changes.

In Objective Caml, compilation units are special cases of structures and signatures, and the relationship between the units can be explained easily in terms of the module system. A compilation unit a comprises two files:

• the implementation file a.ml, which contains a sequence of definitions, analogous to the inside of a struct… end construct;

• the interface file a.mli, which contains a sequence of specifications, analogous to the inside of a sig… end construct.

Both files define a structure named A (same name as the base name a of the two files, with the first letter capitalized), as if the following definition was entered at top-level:

module A: sig (* contents of file a.mli *) end
        = struct (* contents of file a.ml *) end;;

The files defining the compilation units can be compiled separately using the ocaml -c command (the -c option means “compile only, do not try to link”); this produces compiled interface files (with extension .cmi) and compiled object code files (with extension .cmo). When all units have been compiled, their .cmo files are linked together using the ocaml command. For instance, the following commands compile and link a program composed of two compilation units aux and main:

$ ocamlc -c aux.mli                     # produces aux.cmi
$ ocamlc -c aux.ml                      # produces aux.cmo
$ ocamlc -c main.mli                    # produces main.cmi
$ ocamlc -c main.ml                     # produces main.cmo
$ ocamlc -o theprogram aux.cmo main.cmo

The program behaves exactly as if the following phrases were entered at top-level:

module Aux: sig (* contents of aux.mli *) end
          = struct (* contents of aux.ml *) end;;
module Main: sig (* contents of main.mli *) end
           = struct (* contents of main.ml *) end;;

In particular, Main can refer to Aux: the definitions and declarations contained in main.ml and main.mli can refer to definition in aux.ml, using the Aux.ident notation, provided these definitions are exported in aux.mli.

The order in which the .cmo files are given to ocaml during the linking phase determines the order in which the module definitions occur. Hence, in the example above, Aux appears first and Main can refer to it, but Aux cannot refer to Main.

Notice that only top-level structures can be mapped to separately-compiled files, but not functors nor module types. However, all module-class objects can appear as components of a structure, so the solution is to put the functor or module type inside a structure, which can then be mapped to a file.

48

Part II

The Objective Caml language The Objective Caml language

Foreword

This document is intended as a reference manual for the Objective Caml language. It lists the language constructs, and gives their precise syntax and informal semantics. It is by no means a tutorial introduction to the language: there is not a single example. A good working knowledge of Caml is assumed.

No attempt has been made at mathematical rigor: words are employed with their intuitive meaning, without further definition. As a consequence, the typing rules have been left out, by lack of the mathematical framework required to express them, while they are definitely part of a full formal definition of the language.

Notations

The syntax of the language is given in BNF-like notation. Terminal symbols are set in typewriter font (like this). Non-terminal symbols are set in italic font (like that). Square brackets […] denote optional components. Curly brackets {…} denotes zero, one or several repetitions of the enclosed components. Curly bracket with a trailing plus sign {…}^{+} denote one or several repetitions of the enclosed components. Parentheses (…) denote grouping.

4.1 Lexical conventions

Blanks

The following characters are considered as blanks: space, newline, horizontal tabulation, carriage return, line feed and form feed. Blanks are ignored, but they separate adjacent identifiers, literals and keywords that would otherwise be confused as one single identifier, literal or keyword.

Comments

Comments are introduced by the two characters (*, with no intervening blanks, and terminated by the characters *), with no intervening blanks. Comments are treated as blank characters. Comments do not occur inside string or character literals. Nested comments are handled correctly.

Identifiers

ident ::= (letter | _) {letter | 0 … 9 | _| '}
letter ::= A … Z | a … z

Identifiers are sequences of letters, digits, _(the underscore character), and ' (the single quote), starting with a letter or an underscore. Letters contain at least the 52 lowercase and uppercase letters from the ASCII set. The current implementation also recognizes as letters all accented characters from the ISO 8859-1 (“ISO Latin 1”) set, and also allows an underscore _as the first character of an identifier. All characters in an identifier are meaningful. The current implementation places no limits on the number of characters of an identifier.

Integer literals

integer-literal ::= [-] {0 … 9}^{+}
| [-] (0x | 0X) {0 … 9 | A … F | a … f}^{+}
| [-] (0o | 0O) {0 … 7}^{+}
| [-] (0b | 0B) {0 … 1}^{+}

An integer literal is a sequence of one or more digits, optionally preceded by a minus sign. By default, integer literals are in decimal (radix 10). The following prefixes select a different radix:

Prefix Radix
0x, 0X hexadecimal (radix 16)
0o, 0O octal (radix 8)
0b, 0B binary (radix 2)

(The initial 0 is the digit zero; the O for octal is the letter O.) The interpretation of integer literals that fall outside the range of representable integer values is undefined.

Floating-point literals

float-literal ::= [-] {0 … 9}^{+} [. {0 … 9}] [(e | E) [+ | -] {0 … 9}^{+}]

Floating-point decimals consist in an integer part, a decimal part and an exponent part. The integer part is a sequence of one or more digits, optionally preceded by a minus sign. The decimal part is a decimal point followed by zero, one or more digits. The exponent part is the character e or E followed by an optional + or - sign, followed by one or more digits. The decimal part or the exponent part can be omitted, but not both to avoid ambiguity with integer literals. The interpretation of floating-point literals that fall outside the range of representable floating-point values is undefined.

Character literals

char-literal ::= ' regular-char '

                                  | '\ (\ | ' | n | t | b | r) '

| '\ (0 … 9) (0 … 9) (0 … 9) '

Character literals are delimited by ' (single quote) characters. The two single quotes enclose either one character different from ' and \, or one of the escape sequences below:

Sequence Character denoted
\\  backslash (\)
\'  single quote (')
\n  newline (LF)
\r  return (CR)
\t  horizontal tabulation (TAB)
\b  backspace (BS)
\ddd  the character with ASCII code ddd in decimal

String literals

string-literal ::= " {string-character} "
string-character ::= regular-char
| \ (\ | " | n | t | b | r)
| \ (0 … 9) (0 … 9) (0 … 9)

String literals are delimited by " (double quote) characters. The two double quotes enclose a sequence of either characters different from " and \, or escape sequences from the table below:

Sequence Character denoted
\\  backslash (\)
\"  double quote (")
\n  newline (LF)
\r  return (CR)
\t  horizontal tabulation (TAB)
\b  backspace (BS)
\ddd  the character with ASCII code ddd in decimal

The current implementation places no restrictions on the length of string literals.

Prefix and infix symbols

infix-symbol ::= (= | < | > | @ | ^| | | & | + | - | * | / | $ | %) {operator-char}
prefix-symbol ::= (! | ? | ~) {operator-char}
operator-char ::= ! | $ | % | & | * | + | - | . | / | : | < | = | > | ? | @ | ^| | | ~

Sequences of “operator characters”, such as <=> or !!, are read as a single token from the infix-symbol or prefix-symbol class. These symbols are parsed as prefix and infix operators inside expressions, but otherwise behave much as identifiers.

Keywords

The identifiers below are reserved as keywords, and cannot be employed otherwise:

      and         as          assert      asr         begin       class
      closed      constraint  do          done        downto      else
      end         exception   external    false       for         fun
      function    functor     if          in          include     inherit
      land        lazy        let         lor         lsl         lsr
      lxor        match       method      mod         module      mutable
      new         of          open        or          parser      private
      rec         sig         struct      then        to          true
      try         type        val         virtual     when        while
      with

The following character sequences are also keywords:

    #     &     '     (     )     *     ,     ->    ?
    .     ..    .(    .[    :     ::    :=    ;     ;;
    <-    =     [     [|    [<    {<    ]     |]    >]
    >}    _{     |     }

Ambiguities

Lexical ambiguities are resolved according to the “longest match” rule: when a character sequence can be decomposed into two tokens in several different ways, the decomposition retained is the one with the longest first token.

4.2 Values

This section describes the kinds of values that are manipulated by Caml Light programs.

4.2.1 Base values

Integer numbers

Integer values are integer numbers from −2^{30} to 2^{30} − 1, that is −1073741824 to 1073741823. The implementation may support a wider range of integer values: on 64-bit platforms, the current implementation supports integers ranging from −2^{62} to 2^{62} − 1.

Floating-point numbers

Floating-point values are numbers in floating-point representation. The current implementation uses double-precision floating-point numbers conforming to the IEEE 754 standard, with 53 bits of mantissa and an exponent ranging from −1022 to 1023.

Characters

Character values are represented as 8-bit integers between 0 and 255. Character codes between 0 and 127 are interpreted following the ASCII standard. The current implementation interprets character codes between 128 and 255 following the ISO 8859-1 standard.

Character strings

String values are finite sequences of characters. The current implementation supports strings containing up to 2^{24} − 6 characters (16777210 characters).

4.2.2 Tuples

Tuples of values are written (v_{1}, …, v_{n}), standing for the n-tuple of values v_{1} to v_{n}. The current implementation supports tuple of up to 2^{22} − 1 elements (4194303 elements).

4.2.3 Records

Record values are labeled tuples of values. The record value written {label_{1} = v_{1}; …; label_{n} = v_{n}} associates the value v_{i} to the record label label_{i}, for i = 1 … n. The current implementation supports records with up to 2^{22} − 1 fields (4194303 fields).

4.2.4 Arrays

Arrays are finite, variable-sized sequences of values of the same type. The current implementation supports arrays containing to 2^{22} − 1 elements (4194303 elements).

4.2.5 Variant values

Variant values are either a constant constructor, or a pair of a non-constant constructor and a value. The former case is written cconstr; the latter case is written ncconstr(v), where v is said to be the argument of the non-constant constructor ncconstr.

The following constants are treated like built-in constant constructors:

Constant Constructor
false  the boolean false
true  the boolean true
()  the “unit” value
[]  the empty list

The current implementation limits the number of distinct constructors in a given variant type to at most 249.

4.2.6 Functions

Functional values are mappings from values to values.

4.2.7 Objects

Objects are composed of a hidden internal state which is a record of instance variables, and a set of methods for accessing and modifying these variables. The structure of an object is described by the toplevel class that created it.

4.3 Names

Identifiers are used to give names to several classes of language objects and refer to these objects by name later:

• value names (syntactic class value-name),

• value constructors (constant – class cconstr-name – or non-constant – class ncconstr-name),

• type constructors (typeconstr-name),

• record labels (label-name),

• class names (class-name),

• method names (method-name),

• instance variable names (inst-var-name),

• module names (module-name),

• module type names (modtype-name). These nine name spaces are distinguished both by the context and by the capitalization of the identifier: whether the first letter of the identifier is in lowercase (written lowercase-ident below) or in uppercase (written capitalized-ident). Underscore is considered a lowercase letter for this purpose.

Naming objects

value-name ::= lowercase-ident
| ( operator-name )
operator-name ::= prefix-symbol | infix-symbol | * | = | or | & | :=
cconstr-name ::= capitalized-ident

     | false
     | true
     | []
     | ()

ncconstr-name ::= capitalized-ident

     | ::

typeconstr-name ::= lowercase-ident
label-name ::= lowercase-ident
module-name ::= capitalized-ident
modtype-name ::= ident
class-name ::= lowercase-ident
inst-var-name ::= lowercase-ident
method-name ::= lowercase-ident

As shown above, prefix and infix symbols as well as some keywords can be used as value names, provided they are written between parentheses. Keywords such as ’::’ and ’false’ are also constructor names. The capitalization rules are summarized in the table below.

Name space  Case of first letter
Values  lowercase
Constructors  uppercase
Type constructors lowercase
Record labels  lowercase
Classes  lowercase
Methods  lowercase
Modules  uppercase
Module types  any

Referring to named objects

value-path ::= value-name
| module-path . lowercase-ident
cconstr ::= cconstr-name
| module-path . capitalized-ident
ncconstr ::= ncconstr-name
| module-path . capitalized-ident
typeconstr ::= typeconstr-name
| extended-module-path . lowercase-ident
label ::= label-name
| module-path . lowercase-ident
module-path ::= module-name
| module-path . capitalized-ident
extended-module-path ::= module-name
| extended-module-path . capitalized-ident
| extended-module-path ( extended-module-path )
modtype-path ::= modtype-name
| extended-module-path . ident
class-path ::= class-name
| module-path . lowercase-ident

A named object can be referred to either by its name (following the usual static scoping rules for names) or by an access path prefix . name, where prefix designates a module and name is the name of an object defined in that module. The first component of the path, prefix, is either a simple module name or an access path name_{1} . name_{2} …, in case the defining module is itself nested inside other modules. For referring to type constructors or module types, the prefix can also contain simple functor applications (as in the syntactic class extended-module-path above), in case the defining module is the result of a functor application.

Instance variable names and method names need not be qualified: the former are local to a class while the latter are global labels.

4.4 Type expressions

typexpr ::= ' ident
| ( typexpr )
| typexpr -> typexpr
| typexpr {* typexpr}^{+}
| typeconstr
| typexpr typeconstr
| ( typexpr {, typexpr} ) typeconstr
| typexpr as' ident

                              | < [..] >

| < method-type {; method-type} [;..] >
| # class-path
| typexpr # class-path
| ( typexpr {, typexpr}) # class-path
method-type ::= method-name : typexpr

The table below shows the relative precedences and associativity of operators and non-closed type constructions. The constructions with higher precedences come first.

Operator  Associativity
Type constructor application –

                      *  –

->  right

                      as  –

Type expressions denote types in definitions of data types as well as in type constraints over patterns and expressions.

Type variables

The type expression ' ident stands for the type variable named ident. In data type definitions, type variables are names for the data type parameters. In type constraints, they represent unspecified types that can be instantiated by any type to satisfy the type constraint.

Parenthesized types

The type expression ( typexpr ) denotes the same type as typexpr.

Function types

The type expression typexpr_{1} -> typexpr_{2} denotes the type of functions mapping arguments of type typexpr_{1} to results of type typexpr_{2}.

Tuple types

The type expression typexpr_{1} * … * typexpr_{n} denotes the type of tuples whose elements belong to types typexpr_{1}, … typexpr_{n} respectively.

Constructed types

Type constructors with no parameter, as in typeconstr, are type expressions.

The type expression typexpr typeconstr, where typeconstr is a type constructor with one parameter, denotes the application of the unary type constructor typeconstr to the type typexpr.

The type expression (typexpr_{1}, …, typexpr_{n}) typeconstr, where typeconstr is a type constructor with n parameters, denotes the application of the n-ary type constructor typeconstr to the types typexpr_{1} through typexpr_{n}.

Recursive types

The type expression typexpr as' ident denotes the same type as typexpr, and also binds the type variable ident to type typexpr both in typexpr and in the remaining part of the type. If the type variable ident actually occurs in typexpr, a recursive type is created. Recursive types are only allowed when any recursion crosses an object type.

Object types

An object type < method-type {; method-type} > is a record of method types.

The type < method-type {; method-type} ; .. > is the type of an object with methods and their associated types are described by method-type_{1}, …, method-type_{n}, and possibly some other methods represented by the ellipsis. This ellipsis actually is a special kind of type variable, named row variable.

#-types

The type # class-path is a special kind of abbreviation. This abbreviation unifies with the type of any object belonging to a subclass of class class-path. It is handled in a special way as it usually hides a type variable (an ellipsis, representing the methods that may be added in a subclass). In particular, it vanishes when the ellipsis gets instantiated. Each type expression # class-path defines a new type variable, so type # class-path -> # class-path is usually not the same as type # class-path as' ident ->' ident.

4.5 Constants

constant ::= integer-literal
| float-literal
| char-literal
| string-literal
| cconstr

The syntactic class of constants comprises literals from the four base types (integers, floatingpoint numbers, characters, character strings), and constant constructors.

4.6 Patterns

pattern ::= value-name

                             | _

| constant
| pattern as value-name
| ( pattern )
| ( pattern : typexpr )
| pattern | pattern
| ncconstr pattern
| pattern {, pattern}
| { label = pattern {; label = pattern} }
| [ pattern {; pattern} ]
| pattern :: pattern

The table below shows the relative precedences and associativity of operators and non-closed pattern constructions. The constructions with higher precedences come first.

Operator  Associativity
Constructor application –
::  right

                         ,  –

|  left

                         as  –

Patterns are templates that allow selecting data structures of a given shape, and binding identifiers to components of the data structure. This selection operation is called pattern matching; its outcome is either “this value does not match this pattern”, or “this value matches this pattern, resulting in the following bindings of names to values”.

Variable patterns

A pattern that consists in a value name matches any value, binding the name to the value. The pattern _also matches any value, but does not bind any name.

Patterns are linear: a variable cannot appear several times in a given pattern. In particular, there is no way to test for equality between two parts of a data structure using only a pattern (but when guards can be used for this purpose).

Constant patterns

A pattern consisting in a constant matches the values that are equal to this constant.

Alias patterns

The pattern pattern_{1} as value-name matches the same values as pattern_{1}. If the matching against pattern_{1} is successful, the name name is bound to the matched value, in addition to the bindings performed by the matching against pattern_{1}.

Parenthesized patterns

The pattern ( pattern_{1} ) matches the same values as pattern_{1}. A type constraint can appear in a parenthesized pattern, as in ( pattern_{1} : typexpr ). This constraint forces the type of pattern_{1} to be compatible with type.

“Or” patterns

The pattern pattern_{1} | pattern_{2} represents the logical “or” of the two patterns pattern_{1} and pattern_{2}. A value matches pattern_{1} | pattern_{2} either if it matches pattern_{1} or if it matches pattern_{2}. The two sub-patterns pattern_{1} and pattern_{2} must contain no identifiers. Hence no bindings are returned by matching against an “or” pattern.

Variant patterns

The pattern ncconstr pattern_{1} matches all variants whose constructor is equal to ncconstr, and whose argument matches pattern_{1}.

The pattern pattern_{1} :: pattern_{2} matches non-empty lists whose heads match pattern_{1}, and whose tails match pattern_{2}. This pattern behaves like (::)( pattern_{1} , pattern_{2} ).

The pattern [ pattern_{1} ; …; pattern_{n} ] matches lists of length n whose elements match pattern_{1} … pattern_{n}, respectively. This pattern behaves like pattern_{1} :: …:: pattern_{n} ::[].

Tuple patterns

The pattern pattern_{1} , …, pattern_{n} matches n-tuples whose components match the patterns pattern_{1} through pattern_{n}. That is, the pattern matches the tuple values (v_{1}, …, v_{n}) such that pattern_{i} matches v_{i} for i = 1, …, n.

Record patterns

The pattern { label_{1} = pattern_{1} ; …; label_{n} = pattern_{n} } matches records that define at least the labels label_{1} through label_{n}, and such that the value associated to label_{i} match the pattern pattern_{i}, for i = 1, …, n. The record value can define more labels than label_{1} … label_{n}; the values associated to these extra labels are not taken into account for matching.

Chapter 4. The Objective Caml language 63

4.7 Expressions

expr ::= value-path
| constant
| ( expr )

| begin expr end

| ( expr : typexpr )
| expr , expr {, expr}
| ncconstr expr
| expr :: expr
| [ expr {; expr} ]
| [| expr {; expr} |]
| { label = expr {; label = expr} }
| expr expr
| prefix-symbol expr
| expr infix-op expr
| expr . label
| expr . label <- expr
| expr .( expr )
| expr .( expr )<- expr
| expr .[ expr ]
| expr .[ expr ]<- expr
| if expr then expr [else expr]

| while expr do expr done

| for ident = expr (to | downto) expr do expr done
| expr ; expr
| match expr with pattern-matching
| function pattern-matching
| fun multiple-matching
| try expr with pattern-matching
| let [rec] let-binding {and let-binding} in expr
| new class-path
| expr # method-name
| ( expr :> typexpr )
| ( expr : typexpr :> typexpr )
| {< inst-var-name = expr {; inst-var-name = expr} >}
pattern-matching ::= pattern [when expr] -> expr {| pattern [when expr] -> expr}
multiple-matching ::= {pattern}^{+} [when expr] -> expr
let-binding ::= pattern = expr
| value-name {pattern}^{+} [: typexpr] = expr
infix-op ::= infix-symbol

| * | = | or | &

The table below shows the relative precedences and associativity of operators and non-closed constructions. The constructions with higher precedence come first. For infix and prefix symbols, we write “*… ” to mean “any symbol starting with *”.

Construction or operator  Associativity
prefix-symbol  –

.   .(    .[  –

function application  left
constructor application  –
-   -. (prefix)  –
**…  right
*… /… %… mod  left
+… -…  left
::  right
@ ^right
comparisons (=  ==  < etc.), all other infix symbols left

not  –

&   &&  left

or  ||  left
,  –

<-   :=  right

if  –

;  right

let  match  fun  function  try  –

4.7.1 Basic expressions

Constants

Expressions consisting in a constant evaluate to this constant.

Value paths

Expressions consisting in an access path evaluate to the value bound to this path in the current evaluation environment. The path can be either a value name or an access path to a value component of a module.

Parenthesized expressions

The expressions ( expr ) and begin expr end have the same value as expr. Both constructs are semantically equivalent, but it is good style to use begin … end inside control structures:

if . . . then begin . . . ; . . . end else begin . . . ; . . . end

and ( …) for the other grouping situations.

Parenthesized expressions can contain a type constraint, as in ( expr : type ). This constraint forces the type of expr to be compatible with type.

Parenthesized expressions can also contain coercions ( expr [: type] :> type ) (see subsection 4.7.5 below).

Function application

Function application is denoted by juxtaposition of expressions. The expression expr_{1} expr_{2} … expr_{n} evaluates the expressions expr_{1} to expr_{n}. The expression expr_{1} must evaluate to a functional value, which is then applied to the values of expr_{2}, …, expr_{n}. The order in which the expressions expr_{1}, …, expr_{n} are evaluated is not specified.

Function definition

Two syntactic forms are provided to define functions. The first form is introduced by the keyword

function:

function pattern_{1} -> expr_{1}
| …
| pattern_{n} -> expr_{n}

This expression evaluates to a functional value with one argument. When this function is applied to a value v, this value is matched against each pattern pattern_{1} to pattern_{n}. If one of these matchings succeeds, that is, if the value v matches the pattern pattern_{i} for some i, then the expression expr_{i} associated to the selected pattern is evaluated, and its value becomes the value of the function application. The evaluation of expr_{i} takes place in an environment enriched by the bindings performed during the matching.

If several patterns match the argument v, the one that occurs first in the function definition is selected. If none of the patterns matches the argument, the exception Match_failure is raised.

The other form of function definition is introduced by the keyword fun:

fun pattern_{1} … pattern_{n} -> expr

This expression is equivalent to:

function pattern_{1} -> … function pattern_{n} -> expr

That is, the fun expression above evaluates to a curried function with n arguments: after applying this function n times to the values v_{1} … v_{m}, the values will be matched in parallel against the patterns pattern_{1} … pattern_{n}. If the matching succeeds, the function returns the value of expr in an environment enriched by the bindings performed during the matchings. If the matching fails, the exception Match_failure is raised.

Guards in pattern-matchings

Cases of a pattern matching (in the function, fun, match and try constructs) can include guard expressions, which are arbitrary boolean expressions that must evaluate to true for the match case to be selected. Guards occur just before the -> token and are introduced by the when keyword:

function pattern_{1} [when cond_{1}] -> expr_{1}
| …
| pattern_{n} [when cond_{n}] -> expr_{n}

Matching proceeds as described before, except that if the value matches some pattern pattern_{i} which has a guard cond_{i}, then the expression cond_{i} is evaluated (in an environment enriched by the bindings performed during matching). If cond_{i} evaluates to true, then expr_{i} is evaluated and its value returned as the result of the matching, as usual. But if cond_{i} evaluates to false, the matching is resumed against the patterns following pattern_{i}.

Local definitions

The let and let rec constructs bind value names locally. The construct

let pattern_{1} = expr_{1} and … and pattern_{n} = expr_{n} in expr

evaluates expr_{1} … expr_{n} in some unspecified order, then matches their values against the patterns pattern_{1} … pattern_{n}. If the matchings succeed, expr is evaluated in the environment enriched by the bindings performed during matching, and the value of expr is returned as the value of the whole let expression. If one of the matchings fails, the exception Match_failure is raised.

An alternate syntax is provided to bind variables to functional values: instead of writing

let ident =fun pattern_{1} … pattern_{m} -> expr

in a let expression, one may instead write

let ident pattern_{1} … pattern_{m} = expr

Recursive definitions of names are introduced by let rec:

letrec pattern_{1} = expr_{1} and … and pattern_{n} = expr_{n} in expr

The only difference with the let construct described above is that the bindings of names to values performed by the pattern-matching are considered already performed when the expressions expr_{1} to expr_{n} are evaluated. That is, the expressions expr_{1} to expr_{n} can reference identifiers that are bound by one of the patterns pattern_{1}, …, pattern_{n}, and expect them to have the same value as in expr, the body of the let rec construct.

The recursive definition is guaranteed to behave as described above if the expressions expr_{1} to expr_{n} are function definitions (fun … or function …), and the patterns pattern_{1} … pattern_{n} are just value names, as in:

letrec name_{1} =fun … and … and name_{n} =fun … in expr

This defines name_{1} … name_{n} as mutually recursive functions local to expr.

The behavior of other forms of let rec definitions is implementation-dependent. The current implementation also supports a certain class of recursive definitions of non-functional values, such as

letrec name_{1} =1:: name_{2} and name_{2} =2:: name_{1} in expr

which binds name_{1} to the cyclic list 1::2::1::2::…, and name_{2} to the cyclic list 2::1::2::1::… Informally, the class of accepted definitions consists of those definitions where the defined names occur only inside function bodies or as argument to a data constructor.

4.7.2 Control structures

Sequence

The expression expr_{1} ; expr_{2} evaluates expr_{1} first, then expr_{2}, and returns the value of expr_{2}.

Conditional

The expression if expr_{1} then expr_{2} else expr_{3} evaluates to the value of expr_{2} if expr_{1} evaluates to the boolean true, and to the value of expr_{3} if expr_{1} evaluates to the boolean false.

The else expr_{3} part can be omitted, in which case it defaults to else().

Case expression

The expression

                             match expr

with pattern_{1} -> expr_{1}
| …
| pattern_{n} -> expr_{n}

matches the value of expr against the patterns pattern_{1} to pattern_{n}. If the matching against pattern_{i} succeeds, the associated expression expr_{i} is evaluated, and its value becomes the value of the whole match expression. The evaluation of expr_{i} takes place in an environment enriched by the bindings performed during matching. If several patterns match the value of expr, the one that occurs first in the match expression is selected. If none of the patterns match the value of expr, the exception Match_failure is raised.

Boolean operators

The expression expr_{1} & expr_{2} evaluates to true if both expr_{1} and expr_{2} evaluate to true; otherwise, it evaluates to false. The first component, expr_{1}, is evaluated first. The second component, expr_{2}, is not evaluated if the first component evaluates to false. Hence, the expression expr_{1} & expr_{2} behaves exactly as

                           if expr_{1} then expr_{2} elsefalse.

The expression expr_{1} or expr_{2} evaluates to true if one of expr_{1} and expr_{2} evaluates to true; otherwise, it evaluates to false. The first component, expr_{1}, is evaluated first. The second component, expr_{2}, is not evaluated if the first component evaluates to true. Hence, the expression expr_{1} or expr_{2} behaves exactly as

                            if expr_{1} thentrueelse expr_{2}.

Loops

The expression while expr_{1} do expr_{2} done repeatedly evaluates expr_{2} while expr_{1} evaluates to true. The loop condition expr_{1} is evaluated and tested at the beginning of each iteration. The whole while … done expression evaluates to the unit value ().

The expression for name = expr_{1} to expr_{2} do expr_{3} done first evaluates the expressions expr_{1} and expr_{2} (the boundaries) into integer values n and p. Then, the loop body expr_{3} is repeatedly evaluated in an environment where name is successively bound to the values n, n + 1, …, p − 1, p. The loop body is never evaluated if n > p.

The expression for name = expr_{1} downto expr_{2} do expr_{3} done evaluates similarly, except that name is successively bound to the values n, n − 1, …, p + 1, p. The loop body is never evaluated if n < p.

In both cases, the whole for expression evaluates to the unit value ().

Exception handling

The expression

try expr
with pattern_{1} -> expr_{1}
| …
| pattern_{n} -> expr_{n}

evaluates the expression expr and returns its value if the evaluation of expr does not raise any exception. If the evaluation of expr raises an exception, the exception value is matched against the patterns pattern_{1} to pattern_{n}. If the matching against pattern_{i} succeeds, the associated expression expr_{i} is evaluated, and its value becomes the value of the whole try expression. The evaluation of expr_{i} takes place in an environment enriched by the bindings performed during matching. If several patterns match the value of expr, the one that occurs first in the try expression is selected. If none of the patterns matches the value of expr, the exception value is raised again, thereby transparently “passing through” the try construct.

4.7.3 Operations on data structures

Products

The expression expr_{1} , …, expr_{n} evaluates to the n-tuple of the values of expressions expr_{1} to expr_{n}. The evaluation order for the subexpressions is not specified.

Variants

The expression ncconstr expr evaluates to the variant value whose constructor is ncconstr, and whose argument is the value of expr.

For lists, some syntactic sugar is provided. The expression expr_{1} :: expr_{2} stands for the constructor (::) applied to the argument ( expr_{1} , expr_{2} ), and therefore evaluates to the list whose head is the value of expr_{1} and whose tail is the value of expr_{2}. The expression [ expr_{1} ; …; expr_{n} ] is equivalent to expr_{1} :: …:: expr_{n} ::[], and therefore evaluates to the list whose elements are the values of expr_{1} to expr_{n}.

Records

The expression { label_{1} = expr_{1} ; …; label_{n} = expr_{n} } evaluates to the record value { label_{1} = v_{1} ; …; label_{n} = v_{n} }, where v_{i} is the value of expr_{i} for i = 1, …, n. The labels label_{1} to label_{n} must all belong to the same record types; all labels belonging to this record type must appear exactly once in the record expression, though they can appear in any order. The order in which expr_{1} to expr_{n} are evaluated is not specified.

The expression expr_{1} . label evaluates expr_{1} to a record value, and returns the value associated to label in this record value.

The expression expr_{1} . label <- expr_{2} evaluates expr_{1} to a record value, which is then modified in-place by replacing the value associated to label in this record by the value of expr_{2}. This operation is permitted only if label has been declared mutable in the definition of the record type. The whole expression expr_{1} . label <- expr_{2} evaluates to the unit value ().

Arrays

The expression [| expr_{1} ; …; expr_{n} |] evaluates to a n-element array, whose elements are initialized with the values of expr_{1} to expr_{n} respectively. The order in which these expressions are evaluated is unspecified.

The expression expr_{1} .( expr_{2} ) returns the value of element number expr_{2} in the array denoted by expr_{1}. The first element has number 0; the last element has number n − 1, where n is the size of the array. The exception Invalid_argument is raised if the access is out of bounds.

The expression expr_{1} .( expr_{2} )<- expr_{3} modifies in-place the array denoted by expr_{1}, replacing element number expr_{2} by the value of expr_{3}. The exception Invalid_argument is raised if the access is out of bounds. The value of the whole expression is ().

Strings

The expression expr_{1} .[ expr_{2} ] returns the value of character number expr_{2} in the string denoted by expr_{1}. The first character has number 0; the last character has number n − 1, where n is the length of the string. The exception Invalid_argument is raised if the access is out of bounds.

The expression expr_{1} .[ expr_{2} ] <- expr_{3} modifies in-place the string denoted by expr_{1}, replacing character number expr_{2} by the value of expr_{3}. The exception Invalid_argument is raised if the access is out of bounds. The value of the whole expression is ().

4.7.4 Operators

Symbols from the class infix-symbols, as well as the keywords *, =, or and &, can appear in infix position (between two expressions). Symbols from the class prefix-symbols can appear in prefix position (in front of an expression).

Infix and prefix symbols do not have a fixed meaning: they are simply interpreted as applications of functions bound to the names corresponding to the symbols. The expression prefix-symbol expr is interpreted as the application ( prefix-symbol ) expr. Similarly, the expression expr_{1} infix-symbol expr_{2} is interpreted as the application ( infix-symbol ) expr_{1} expr_{2}.

The table below lists the symbols defined in the initial environment and their initial meaning. (See the description of the standard library module Pervasive in chapter 15 for more details). Their meaning may be changed at any time using let( infix-op ) name_{1} name_{2} = … Operator Initial meaning + Integer addition.

- (infix) Integer subtraction.

- (prefix) Integer negation.

* Integer multiplication. / Integer division. Raise Division_by_zero if second argument is zero.

The result is unspecified if either argument is negative.

mod Integer modulus. Raise Division_by_zero if second argument is zero.

The result is unspecified if either argument is negative.

land Bitwise logical “and” on integers. lor Bitwise logical “or on integers. lxor Bitwise logical “exclusive or” on integers. lsl Bitwise logical shift left on integers. lsr Bitwise logical shift right on integers. asr Bitwise arithmetic shift right on integers. +. Floating-point addition. -. (infix) Floating-point subtraction. -. (prefix) Floating-point negation. *. Floating-point multiplication. /. Floating-point division. ** Floating-point exponentiation. @ List concatenation. ^String concatenation. ! Dereferencing (return the current contents of a reference). := Reference assignment (update the reference given as first argument with

the value of the second argument).

= Structural equality test. <> Structural inequality test. == Physical equality test. != Physical inequality test. < Test “less than”. <= Test “less than or equal”. > Test “greater than”. >= Test “greater than or equal”

4.7.5 Objects

Object creation

The expression new class-path denotes a function that takes some initialization arguments and returns a new object of class class-path.

Message sending

The expression expr # method-name invokes the method method-name of the object denoted by expr.

Coercion

The type of an object can be coerced (weakened) to a supertype. The expression ( expr :> typexpr ) coerces the expression expr to type typexpr. The expression ( expr : typexpr_{1} :> typexpr_{2} ) coerces the expression expr from type typexpr_{1} to type typexpr_{2}. The former operator will sometimes fail to coerce an expression expr from a type t_{1} to a type t_{2} even if type t_{1} is a subtype of type t_{2}. In this case, the latter operator should be used.

In a class definition, coercion to the type this class defines is the identity, as this type abbreviation is not yet completely defined.

Object duplication

An object can be duplicated using the library function Oo.copy (see section 16.16). Inside a method, the expression {< inst-var-name = expr {; inst-var-name = expr} >} returns a copy of self with the given instance variables replaced by the values of the associated expressions; other instance variables have the same value in the returned object as in self.

4.8 Type and exception definitions

4.8.1 Type definitions

Type definitions bind type constructors to data types: either variant types, record types, type abbreviations, or abstract data types. They also bind the value constructors and record labels associated with the definition.

type-definition ::= type typedef {and typedef }
typedef ::= [type-params] typeconstr-name [type-equation] [type-representation] {constraint}
type-equation ::= = typexpr
type-representation ::= = constr-decl {| constr-decl}
| ={ label-decl {; label-decl} }
type-params ::= ' ident
| (' ident {,' ident} )
constr-decl ::= cconstr-name
| ncconstr-name of typexpr
label-decl ::= label-name : typexpr
| mutable label-name : typexpr
constraint ::= constraint' ident = typexpr

Type definitions are introduced by the type keyword, and consist in one or several simple definitions, possibly mutually recursive, separated by the and keyword. Each simple definition defines one type constructor.

A simple definition consists in a lowercase identifier, possibly preceded by one or several type parameters, and followed by an optional type equation, then an optional type representation, and then a constraint clause. The identifier is the name of the type constructor being defined.

The optional type parameters are either one type variable ' ident, for type constructors with one parameter, or a list of type variables (' ident_{1}, …, ' ident_{n}), for type constructors with several parameters. These type parameters can appear in the type expressions of the right-hand side of the definition.

The optional type equation = typexpr makes the defined type equivalent to the type expression typexpr on the right of the = sign: one can be substituted for the other during typing. If no type equation is given, a new type is generated: the defined type is incompatible with any other type.

The optional type representation describes the data structure representing the defined type, by giving the list of associated constructors (if it is a variant type) or associated labels (if it is a record type). If no type representation is given, nothing is assumed on the structure of the type besides what is stated in the optional type equation.

The type representation = constr-decl {| constr-decl} describes a variant type. The constructor declarations constr-decl_{1}, …, constr-decl_{n} describe the constructors associated to this variant type. The constructor declaration ncconstr-name of typexpr declares the name ncconstr-name as a nonconstant constructor, whose argument has type typexpr. The constructor declaration cconstr-name declares the name cconstr-name as a constant constructor. Constructor names must be capitalized.

The type representation = { label-decl {; label-decl} } describes a record type. The label declarations label-decl_{1}, …, label-decl_{n} describe the labels associated to this record type. The label declaration label-name : typexpr declares label-name as a label whose argument has type typexpr. The label declaration mutable label-name : typexpr behaves similarly; in addition, it allows physical modification over the argument to this label.

The two components of a type definition, the optional equation and the optional representation, can be combined independently, giving rise to four typical situations:

Abstract type: no equation, no representation.

When appearing in a module signature, this definition specifies nothing on the type constructor, besides its number of parameters: its representation is hidden and it is assumed incompatible with any other type.

Type abbreviation: an equation, no representation.

This defines the type constructor as an abbreviation for the type expression on the right of the = sign.

New variant type or record type: no equation, a representation.

This generates a new type constructor and defines associated constructors or labels, through which values of that type can be directly built or inspected.

Re-exported variant type or record type: an equation, a representation.

In this case, the type constructor is defined as an abbreviation for the type expression given in the equation, but in addition the constructors or labels given in the representation remain attached to the defined type constructor. The type expression in the equation part must agree with the representation: it must be of the same kind (record or variant) and have exactly the same constructors or labels, in the same order, with the same arguments.

The construct constraint ' ident = typexpr allows to specify type parameters. Any actual type argument corresponding to the type parameter ident have to be an instance of typexpr (more precisely, ident and typexpr are unified). Type variables of typexpr can appear in the type equation and the type declaration.

4.8.2 Exception definitions

exception-definition ::= exception constr-decl

Exception definitions add new constructors to the built-in variant type exn of exception values. The constructors are declared as for a definition of a variant type.

4.9 Classes

4.9.1 Class definitions

class-definition ::= class-header = {constraint} {class-fields}
class-header ::= class-tags parameterized-class-name class-params class-binders
class-tags ::= [virtual] [closed]
parameterized-class-name ::= class-name
| ' ident class-name
| (' ident {,' ident} ) class-name
class-params ::= {pattern}^{+}
class-binders ::= [as value-name] [:' ident]
class-fields ::= inherit ancestor
| val value

                          | virtual [protected] method-type
                          | method [protected] method

ancestor ::= [( typexpr {, typexpr} )] class-path {expr}^{+} [as value-name]
value ::= [private] [mutable] inst-var-name [= expr]
method ::= method-name {pattern} = expr
method-type ::= method-name : typexpr

Class definitions

A class definition class class-definition {and class-definition} is recursive. Each class-definition defines a class-name that can be used in the whole expression except for inheritance. It can also be used for inheritance, but only in the definitions that follow its own.

Type parameters

A class class-name automatically defines two abbreviations : class-name and # class-name. The first one is the type of objects of this class, while the second is more general as it unifies with the type of any object belonging to a subclass (see section 4.4).

The class type parameters correspond to the ones of these two abbreviations. They must be bound to actual types in the class definition using type constraints. So that the abbreviations are well-formed, type variables of the inferred type of the class must either be type parameters or be bound in the constraint clause.

Class parameters

The parameters class-params are the ones of the object creation function new class-path, as well as the ones of the inheritance construct.

Self and self type binders

The binders class-binders, i.e. as value-name and :' ident, allow to bind self (the current object) and its type, respectively. The variable value-name can then be used as any variable in method body.

Constraints on type parameters

The construct constraint ' ident = typexpr allows to specify type parameters. The value of the type parameter ident will be an instance of typexpr (more precisely, ident and typexpr are unified).

Inheritance

The inheritance construct inherit [( typexpr {, typexpr} )] class-path {expr}^{+} allows to reuse methods from other classes. It adds the instance variables and methods from class class-path into the current class, possibly overriding previously defined ones of the same name. Parent instance variables are initialized with parent class parameters bound to the arguments {expr}^{+}. Parent type parameters are unified with type arguments typexpr_{1}, …, typexpr_{n}.

An ancestor can be bound by prepending the construct as value-name to the inheritance construct above. value-name is not a true variable and can only be used to select a method, i.e. in an expression value-name # method-name. This gives access to the method method-name as it was defined in the parent class even if it is redefined in the current class.

Instance variable definition

Instance variables can be defined or their status can be changed using the construct val value.

The definition val [private] [mutable] inst-var-name = expr adds an instance variable inst-var-name whose initial value is the value of expression expr. If the variable was previously defined, its previous initial value is overridden. By default, this variable is visible in subclasses of current class. The flag private makes this variable only visible in the current class. The flag mutable allows physical modification of this variable by methods.

The construct val [private] [mutable] inst-var-name enables to change the variable status (makes it private and/or mutable), while keeping the same initial value.

Method definition

Method definition is written method method. The definition of a method overrides any previous definition of this method. The method will be public (that is, not protected) if any of the definition states so.

A protected method, method␣protected method or virtualΩprotected method, is a method that can only be invoked on self (from other methods of the current class as well as of subclasses of the current class). This invocation is performed using the expression value-name # method-name, where value-name is directly bound to self at the beginning of the class definition. Protected methods do not appear in object types.

Method bodies do not have access to class parameters, but to instance variables. Some special expressions are available in method bodies for manipulating instance variables and duplicating self:

expr ::= …
| inst-var-name
| inst-var-name <- expr
| {< [inst-var-name = expr {; inst-var-name = expr}] >}

The expression inst-var-name evaluates to the value of the corresponding instance variable in the current object, while the expression inst-var-name <- expr modifies in-place the current object by replacing the value associated to inst-var-name by the value of expr. Of course, this instance variable must have been declared mutable.

The expression {< [inst-var-name = expr {; inst-var-name = expr}] >} evaluates to a copy of the current object in which the values of instance variables inst-var-name_{1}, …, inst-var-name_{n} have been replaced by the values of the corresponding expressions expr_{1}, …, expr_{n}.

Virtual class

Methods can be declared, without being defined, with the construct virtual method-type. Methods that are declared in this way or applied to self but not actually defined are said to be virtual. A class must be flagged virtual if one of its methods is virtual. Objects cannot be created from a virtual class.

Closed class

A class can be flagged as closed. A closed class is a class to which subclasses cannot add methods. If the compiler complains that a class must be closed, this usually means that the type of self has been unintentionally unified with a closed object type.

4.9.2 Class types

Class types are specifications for class definitions. The syntax of class types is closely modeled on class definitions.

The type of a concrete method can be omitted. That way, the status of a method can be changed from virtual to concrete without having the method type be given.

Some type information can be hidden in a class type: instance variables and protected concrete methods can be omitted; previously non-mutable instance variables can be flagged mutable.

class-type ::= class-type-header = {constraint} {class-type-fields}
class-type-header ::= class-tags parameterized-class-name class-type-params [:' ident]
class-type-params ::= {( typexpr )}^{+}
class-type-fields ::= inherit ancestor-type
| val value-type

     | virtual [protected] method-type

| method [protected] method-name [: typexpr]
ancestor-type ::= [( typexpr {, typexpr} )] class-path
value-type ::= [private] [mutable] inst-var-name [: typexpr]
method-type ::= method-name : typexpr

4.10 Module types (module specifications)

Module types are the module-level equivalent of type expressions: they specify the general shape and type properties of modules.

module-type ::= modtype-path
| sig {specification [;;]} end
| functor( module-name : module-type )-> module-type
| module-type with constraint {and constraint}
| ( module-type )
specification ::= val value-name : typexpr
| external value-name : typexpr = external-declaration
| type-definition
| exception-definition
| class class-type {and class-type} end
| module module-name : module-type
| module module-name {( module-name : module-type )} : module-type
| moduletype modtype-name
| moduletype modtype-name = module-type
| open module-path
constraint ::= type [type-parameters] typeconstr = typexp
| module module-path = extended-module-path

4.10.1 Simple module types

The expression modtype-path is equivalent to the module type bound to the name modtype-path. The expression ( module-type ) denotes the same type as module-type.

4.10.2 Signatures

Signatures are type specifications for structures. Signatures sig … end are collections of type specifications for value names, type names, exceptions, module names and module type names. A structure will match a signature if the structure provides definitions (implementations) for all the names specified in the signature (and possibly more), and these definitions meet the type requirements given in the signature.

For compatibility with Caml Light, an optional ;; is allowed after each specification in a signature. The ;; has no semantic meaning.

Value specifications

A specification of a value component in a signature is written val value-name : typexpr, where value-name is the name of the value and typexpr its expected type.

The form external value-name : typexpr = external-declaration is similar, except that it requires in addition the name to be implemented as the external function specified in external-declaration (see chapter 14).

Type specifications

A specification of one or several type components in a signature is written type typedef {and typedef } and consists of a sequence of mutually recursive definitions of type names.

Each type definition in the signature specifies an optional type equation = typexp and an optional type representation = constr-decl … or = { label-decl … }. The implementation of the type name in a matching structure must be compatible with the type expression specified in the equation (if given), and have the specified representation (if given). Conversely, users of that signature will be able to rely on the type equation or type representation, if given. More precisely, we have the following four situations:

Abstract type: no equation, no representation.

Names that are defined as abstract types in a signature can be implemented in a matching structure by any kind of type definition (provided it has the same number of type parameters). The exact implementation of the type will be hidden to the users of the structure. In particular, if the type is implemented as a variant type or record type, the associated constructors and labels will not be accessible to the users; if the type is implemented as an abbreviation, the type equality between the type name and the right-hand side of the abbreviation will be hidden from the users of the structure. Users of the structure consider that type as incompatible with any other type: a fresh type has been generated.

Type abbreviation: an equation = typexp, no representation.

The type name must be implemented by a type compatible with typexp. All users of the structure know that the type name is compatible with typexp.

New variant type or record type: no equation, a representation.

The type name must be implemented by a variant type or record type with exactly the constructors or labels specified. All users of the structure have access to the constructors or labels, and can use them to create or inspect values of that type. However, users of the structure consider that type as incompatible with any other type: a fresh type has been generated.

Re-exported variant type or record type: an equation, a representation.

This case combines the previous two: the representation of the type is made visible to all users, and no fresh type is generated.

Exception specification

The specification exception constr-decl in a signature requires the matching structure to provide an exception with the name and arguments specified in the definition, and makes the exception available to all users of the structure.

Class specifications

A class specification is composed of one or several mutually recursive class types, as described in section 4.9.2.

Module specifications

A specification of a module component in a signature is written module module-name : module-type, where module-name is the name of the module component and module-type its expected type. Modules can be nested arbitrarily; in particular, functors can appear as components of structures and functor types as components of signatures.

For specifying a module component that is a functor, one may write

module module-name ( name_{1} : module-type_{1} ) … ( name_{n} : module-type_{n} ): module-type

instead of

module module-name :functor( name_{1} : module-type_{1} )-> … -> module-type

Module type specifications

A module type component of a signature can be specified either as a manifest module type or as an abstract module type.

An abstract module type specification module type modtype-name allows the name modtype-name to be implemented by any module type in a matching signature, but hides the implementation of the module type to all users of the signature.

A manifest module type specification module type modtype-name = module-type requires the name modtype-name to be implemented by the module type module-type in a matching signature, but makes the equality between modtype-name and module-type apparent to all users of the signature.

Opening a module path

The expression open module-path in a signature does not specify any components. It simply affects the parsing of the following items of the signature, allowing components of the module denoted by module-path to be referred to by their simple names name instead of path accesses module-path . name. The scope of the open stops at the end of the signature expression.

4.10.3 Functor types

The module type expression functor ( module-name : module-type_{1} ) -> module-type_{2} is the type of functors (functions from modules to modules) that take as argument a module of type module-type_{1} and return as result a module of type module-type_{2}. The module type module-type_{2} can use the name module-name to refer to type components of the actual argument of the functor. No restrictions are placed on the type of the functor argument; in particular, a functor may take another functor as argument (“higher-order” functor).

4.10.4 The with operator

Assuming module-type denotes a signature, the expression module-type with constraint {and constraint} denotes the same signature where type equations have been added to some of the type specifications, as described by the constraints following the with keyword. The constraint type [type-parameters] typeconstr = typexp adds the type equation = typexp to the specification of the type component named typeconstr of the constrained signature. The constraint module module-path = extended-module-path adds type equations to all type components of the sub-structure denoted by module-path, making them equivalent to the corresponding type components of the structure denoted by extended-module-path.

For instance, if the module type name S is bound to the signature

        sig type t module M: (sig type u end) end

then S with type t=int denotes the signature

        sig type t=int module M: (sig type u end) end

and S with module M = N denotes the signature

        sig type t module M: (sig type u=N.u end) end

A functor taking two arguments of type S that share their t component is written

        functor (A: S) (B: S with type t = A.t) ...

4.11 Module expressions (module implementations)

Module expressions are the module-level equivalent of value expressions: they evaluate to modules, thus providing implementations for the specifications expressed in module types.

module-expr ::= module-path
| struct {definition [;;]} end
| functor( module-name : module-type )-> module-expr
| module-expr ( module-expr )
| ( module-expr )
| ( module-expr : module-type )
definition ::= let [rec] let-binding {and let-binding}
| external value-name : typexpr = external-declaration
| type-definition
| exception-definition
| class class-definition {and class-definition} end
| module module-name {( module-name : module-type )} [: module-type] = module-expr
| moduletype modtype-name = module-type
| open module-path

4.11.1 Simple module expressions

The expression module-path evaluates to the module bound to the name module-path. The expression ( module-expr ) evaluates to the same module as module-expr. The expression ( module-expr : module-type ) checks that the type of module-expr is a

subtype of module-type, that is, that all components specified in module-type are implemented in module-expr, and their implementation meets the requirements given in module-type. In other terms, it checks that the implementation module-expr meets the type specification module-type. The whole expression evaluates to the same module as module-expr, except that all components not specified in module-type are hidden and can no longer be accessed.

4.11.2 Structures

Structures struct … end are collections of definitions for value names, type names, exceptions, module names and module type names. The definitions are evaluated in the order in which they appear in the structure. The scope of the bindings performed by the definitions extend to the end of the structure. As a consequence, a definition may refer to names bound by earlier definitions in the same structure.

For compatibility with toplevel phrases (chapter 7) and with Caml Light, an optional ;; is allowed after each definition in a structure. The ;; has no semantic meaning. Also for compatibility, ;; expr is allowed as a component of a structure, meaning let _= expr, i.e. evaluate expr for its side-effects.

Value definitions

A value definition let [rec] let-binding {and let-binding} bind value names in the same way as a let … in … expression (see section 4.7.1). The value names appearing in the left-hand sides of the bindings are bound to the corresponding values in the right-hand sides.

A value definition external value-name : typexpr = external-declaration implements value-name as the external function specified in external-declaration (see chapter 14).

Type definitions

A definition of one or several type components is written type typedef {and typedef } and consists of a sequence of mutually recursive definitions of type names.

Exception definitions

Exceptions are defined with the syntax exception constr-decl.

Class definitions

The class keyword introduces one or several mutually recursive class definitions, as described in section 4.9.1.

Module definitions

The basic form for defining a module component is module module-name = module-expr, which evaluates module-expr and binds the result to the name module-name.

One can write

module module-name : module-type = module-expr

instead of

module module-name =( module-expr : module-type ).

Another derived form is

module module-name ( name_{1} : module-type_{1} ) … ( name_{n} : module-type_{n} )= module-expr

which is equivalent to

module module-name =functor( name_{1} : module-type_{1} )-> … -> module-expr

Module type definitions

A definition for a module type is written moduletype modtype-name = module-type. It binds the name modtype-name to the module type denoted by the expression module-type.

Opening a module path

The expression open module-path in a structure does not define any components nor perform any bindings. It simply affects the parsing of the following items of the structure, allowing components of the module denoted by module-path to be referred to by their simple names name instead of path accesses module-path . name. The scope of the open stops at the end of the structure expression.

4.11.3 Functors

Functor definition

The expression functor ( module-name : module-type ) -> module-expr evaluates to a functor that takes as argument modules of the type module-type_{1}, binds module-name to these modules, evaluates module-expr in the extended environment, and returns the resulting modules as results. No restrictions are placed on the type of the functor argument; in particular, a functor may take another functor as argument (“higher-order” functor).

Functor application

The expression module-expr_{1} ( module-expr_{2} ) evaluates module-expr_{1} to a functor and module-expr_{2} to a module, and applies the former to the latter. The type of module-expr_{2} must match the type expected for the arguments of the functor module-expr_{1}.

4.12 Compilation units

unit-interface ::= {specification [;;]}
unit-implementation ::= {definition [;;]}

Compilation units bridge the module system and the separate compilation system. A compilation unit is composed of two parts: an interface and an implementation. The interface contains a sequence of specifications, just as the inside of a sig … end signature expression. The implementation contains a sequence of definitions, just as the inside of a struct … end module expression. A compilation unit also has a name unit-name, derived from the names of the files containing the interface and the implementation (see chapter 6 for more details). A compilation unit behaves roughly as the module definition

module unit-name :sig unit-interface end=struct unit-implementation end

A compilation unit can refer to other compilation units by their names, as if they were regular modules. For instance, if U is a compilation unit that defines a type t, other compilation units can refer to that type under the name U.t; they can also refer to U as a whole structure. Except for names of other compilation units, a unit interface or unit implementation must not have any other free variables. In other terms, the type-checking and compilation of an interface or implementation proceeds in the initial environment

name_{1} :sig interface_{1} end … name_{n} :sig interface_{n} end

where name_{1} … name_{N} are the names of the other compilation units available in the search path (see chapter 6 for more details) and interface_{1} … interface_{n} are their respective interfaces.

Language extensions

This chapter describes the language features that are implemented in Objective Caml, but not described in the Objective Caml reference manual. In contrast with the fairly stable kernel language that is described in the reference manual, the extensions presented here are still experimental, and may be removed or changed in the future.

5.1 Streams and stream parsers

Objective Caml comprises a library type for streams (possibly infinite sequences of elements, that are evaluated on demand), and associated stream expressions, to build streams, and stream patterns, to destructure streams. Streams and stream patterns provide a natural approach to the writing of recursive-descent parsers.

Streams are presented by the following extensions to the syntactic classes of expressions:

expr ::= …

                    | [<>]

| [< stream-component {; stream-component} >]
| parser [pattern] stream-matching
| match expr withparser [pattern] stream-matching
stream-component ::= ' expr
| expr
stream-matching ::= stream-pattern [pattern] -> expr {| stream-pattern [pattern] -> expr}
stream-pattern ::= [<>]
| [< stream-pat-comp {; stream-pat-comp [? expr]} >]
stream-pat-comp ::= ' pattern [when expr]
| pattern = expr
| ident

Stream expressions are bracketed by [< and >]. They represent the concatenation of their components. The component ' expr represents the one-element stream whose element is the value of expr. The component expr represents a sub-stream. For instance, if both s and t are streams of integers, then [<'1; s; t; '2>] is a stream of integers containing the element 1, then the elements of s, then those of t, and finally 2. The empty stream is denoted by [< >].

Unlike any other kind of expressions in the language, stream expressions are submitted to lazy evaluation: the components are not evaluated when the stream is built, but only when they are accessed during stream matching. The components are evaluated once, the first time they are accessed; the following accesses reuse the value computed the first time.

Stream patterns, also bracketed by [< and >], describe initial segments of streams. In particular, the stream pattern [< >] matches all streams. Stream pattern components are matched against the corresponding elements of a stream. The component ' pattern matches the corresponding stream element against the pattern; if followed by when, the match is accepted only if the result of the guard expression is true. The component pattern = expr applies the function denoted by expr to the current stream, then matches the result of the function against pattern. Finally, the component ident simply binds the identifier to the stream being matched.

Stream matching proceeds destructively: once a component has been matched, it is discarded from the stream (by in-place modification).

Stream matching proceeds in two steps: first, a pattern is selected by matching the stream against the first components of the stream patterns; then, the following components of the selected pattern are checked against the stream. If the following components do not match, the exception Stream.Parse_error is raised. There is no backtracking here: stream matching commits to the pattern selected according to the first element. If none of the first components of the stream patterns

match, the exception Stream.Parse_failure is raised. The Stream.Parse_failure exception

causes the next alternative to be tried, if it occurs during the matching of the first element of a stream, before matching has committed to one pattern.

The streams hold the count of their elements discarded. The optional pattern before the first stream pattern is bound to the stream count before the matching. The one after each stream pattern (optional, too) is bound to the stream count after the matching.

The exception Parse_error has a string parameter coming from the optional ? expr after the stream pattern components (its default is the empty string). This expression is evaluated only in case of error.

See Functional programming using Caml Light for a more gentle introductions to streams, and for some examples of their use in writing parsers. A more formal presentation of streams, and a discussion of alternate semantics, can be found in Parsers in ML by Michel Mauny and Daniel de Rauglaudre, in the proceedings of the 1992 ACM conference on Lisp and Functional Programming.

5.2 Range patterns

In patterns, Objective Caml recognizes the form ' c ' .. ' d ' (two character literals separated by ..) as shorthand for the pattern

                       ' c '|' c_{1} '|' c_{2} '| . . . |' c_{n} '|' d '

where c_{1}, c_{2}, …, c_{n} are the characters that occur between c and d in the ASCII character set. For instance, the pattern '0'..'9' matches all characters that are digits.

5.3 Assertion checking

Objective Caml supports the assert construct to check debugging assertions. The expression assert expr evaluates the expression expr and returns () if expr evaluates to true. Otherwise, the exception Assert_failure is raised with the location of expr in the source code as argument. As a special case, assert false is reduced to raise (Assert_failure ...), which is polymorphic. Assertion checking can be turned off with the -noassert compiler option.

5.4 Deferred computations

The expression lazy expr returns a value v of type Lazy.t that encapsulates the computation of expr. The argument expr is not evaluated at this point in the program. Instead, its evaluation will be performed the first time Lazy.force is applied to the value v, returning the actual value of expr. Subsequent applications of Lazy.force to v do not evaluate expr again.

The expression lazy expr is equivalent to ref (Lazy.Delayed (fun () -> expr)). For more information, see the description of module Lazy in the standard library (section 16.11).

86

Part III

The Objective Caml tools Batch compilation (ocamlc)

This chapter describes the Objective Caml batch compiler ocamlc, which compiles Caml source files to bytecode object files and link these object files to produce standalone bytecode executable files. These executable files are then run by the bytecode interpreter ocamlrun.

6.1 Overview of the compiler

The ocamlc command has a command-line interface similar to the one of most C compilers. It accepts several types of arguments:

• Arguments ending in .mli are taken to be source files for compilation unit interfaces. Interfaces specify the names exported by compilation units: they declare value names with their types, define public data types, declare abstract data types, and so on. From the file x.mli, the ocamlc compiler produces a compiled interface in the file x.cmi.

• Arguments ending in .ml are taken to be source files for compilation unit implementations. Implementations provide definitions for the names exported by the unit, and also contain expressions to be evaluated for their side-effects. From the file x.ml, the ocamlc compiler produces compiled object bytecode in the file x.cmo.

If the interface file x.mli exists, the implementation x.ml is checked against the corresponding compiled interface x.cmi, which is assumed to exist. If no interface x.mli is provided, the compilation of x.ml produces a compiled interface file x.cmi in addition to the compiled object code file x.cmo. The file x.cmi produced corresponds to an interface that exports everything that is defined in the implementation x.ml.

• Arguments ending in .cmo are taken to be compiled object bytecode. These files are linked together, along with the object files obtained by compiling .ml arguments (if any), and the Caml Light standard library, to produce a standalone executable program. The order in which .cmo and .ml arguments are presented on the command line is relevant: compilation units are initialized in that order at run-time, and it is a link-time error to use a component of a unit before having initialized it. Hence, a given x.cmo file must come before all .cmo files that refer to the unit x.

• Arguments ending in .cma are taken to be libraries of object bytecode. A library of object bytecode packs in a single file a set of object bytecode files (.cmo files). Libraries are built with ocamlc -a (see the description of the -a option below). The object files contained in the library are linked as regular .cmo files (see above), in the order specified when the .cma file was built. The only difference is that if an object file contained in a library is not referenced anywhere in the program, then it is not linked in.

• Arguments ending in .c are passed to the C compiler, which generates a .o object file. This object file is linked with the program if the -custom flag is set (see the description of -custom below).

• Arguments ending in .o or .a are assumed to be C object files and libraries. They are passed to the C linker when linking in -custom mode (see the description of -custom below).

The output of the linking phase is a file containing compiled bytecode that can be executed by the Objective Caml bytecode interpreter: the command named ocamlrun. If caml.out is the name of the file produced by the linking phase, the command

   ocamlrun caml.out arg_{1} arg_{2} . . . arg_{n}

executes the compiled code contained in caml.out, passing it as arguments the character strings arg_{1} to arg_{n}. (See chapter 8 for more details.)

On most Unix systems, the file produced by the linking phase can be run directly, as in:

./caml.out arg_{1} arg_{2} … arg_{n}

The produced file has the executable bit set, and it manages to launch the bytecode interpreter by itself.

6.2 Options

The following command-line options are recognized by ocamlc.

-a Build a library (.cma file) with the object files (.cmo files) given on the command line, instead of linking them into an executable file. The name of the library can be set with the -o option. The default name is library.cma.

-c Compile only. Suppress the linking phase of the compilation. Source code files are turned into compiled files, but no executable file is produced. This option is useful to compile modules separately.

-cclib -llibname

Pass the -llibname option to the C linker when linking in “custom runtime” mode (see the -custom option). This causes the given C library to be linked with the program.

-ccopt option

Pass the given option to the C compiler and linker, when linking in “custom runtime” mode (see the -custom option). For instance, -ccopt -Ldir causes the C linker to search for C libraries in directory dir.

-custom

Link in “custom runtime” mode. In the default linking mode, the linker produces bytecode that is intended to be executed with the shared runtime system, ocamlrun. In the custom runtime mode, the linker produces an output file that contains both the runtime system and the bytecode for the program. The resulting file is larger, but it can be executed directly, even if the ocamlrun command is not installed. Moreover, the “custom runtime” mode enables linking Caml code with user-defined C functions, as described in chapter 14.

-g Add debugging information while compiling and linking. This option is required in order to be able to debug the program with ocamldebug (see chapter 12).

-i Cause the compiler to print all defined names (with their inferred types or their definitions) when compiling an implementation (.ml file). This can be useful to check the types inferred by the compiler. Also, since the output follows the syntax of interfaces, it can help in writing an explicit interface (.mli file) for a file: just redirect the standard output of the compiler to a .mli file, and edit that file to remove all declarations of unexported names.

-I directory Add the given directory to the list of directories searched for compiled interface files (.cmi) and compiled object code files (.cmo). By default, the current directory is searched first, then the standard library directory. Directories added with -I are searched after the current directory, in the order in which they were given on the command line, but before the standard library directory.

-impl filename Compile the file filename as an implementation file, even if its extension is not .ml.

-intf filename Compile the file filename as an interface file, even if its extension is not .mli.

-linkall

Force all modules contained in libraries to be linked in. If this flag is not given, unreferenced modules are not linked in. When building a library (-a flag), setting the -linkall flag forces all subsequent links of programs involving that library to link all the modules contained in the library.

-noassert

Turn assertion checking off: assertions are not compiled. This flag has no effect when linking already compiled files.

-o exec-file Specify the name of the output file produced by the linker. The default output name is a.out, in keeping with the Unix tradition. If the -a option is given, specify the name of the library produced. If the -output-obj option is given, specify the name of the output file produced.

-output-obj

Cause the linker to produce a C object file instead of a bytecode executable file. This is useful to wrap Caml code as a C library, callable from any C program. See chapter 14, section 14.6.5. The name of the output object file is camlprog.o by default; it can be set with the -o option.

-pp command Cause the compiler to call the given command as a preprocessor for each source file. The output of command is redirected to an intermediate file, which is compiled. If there are no compilation errors, the intermediate file is deleted afterwards. The name of this file is built from the basename of the source file with the extension .ppi for an interface (.mli) file and .ppo for an implementation (.ml) file.

-thread

Compile or link multithreaded programs, in combination with the threads library described in chapter 20. What this option actually does is select a special, thread-safe version of the standard library.

-unsafe

Turn bound checking off on array and string accesses (the v.(i) and s.[i] constructs). Programs compiled with -unsafe are therefore slightly faster, but unsafe: anything can happen if the program accesses an array or string outside of its bounds.

-v Print the version number of the compiler.

6.3 Modules and the file system

This short section is intended to clarify the relationship between the names of the modules corresponding to compilation units and the names of the files that contain their compiled interface and compiled implementation.

The compiler always derives the module name by taking the capitalized base name of the source file (.ml or .mli file). That is, it strips the leading directory name, if any, as well as the .ml or .mli suffix; then, it set the first letter to uppercase, in order to comply with the requirement that module names must be capitalized. For instance, compiling the file mylib/misc.ml provides an implementation for the module named Misc. Other compilation units may refer to components defined in mylib/misc.ml under the names Misc.name; they can also do open Misc, then use unqualified names name.

The .cmi and .cmo files produced by the compiler have the same base name as the source file. Hence, the compiled files always have their base name equal (modulo capitalization of the first letter) to the name of the module they describe (for .cmi files) or implement (for .cmo files).

When the compiler encounters a reference to a free module identifier Mod, it looks in the search path for a file mod.cmi (note lowercasing of first letter) and loads the compiled interface contained in that file. As a consequence, renaming .cmi files is not advised: the name of a .cmi file must always correspond to the name of the compilation unit it implements. It is admissible to move them to another directory, if their base name is preserved, and the correct -I options are given to the compiler. The compiler will flag an error if it loads a .cmi file that has been renamed.

Compiled bytecode files (.cmo files), on the other hand, can be freely renamed once created. That’s because the linker never attempts to find by itself the .cmo file that implements a module with a given name: it relies instead on the user providing the list of .cmo files by hand.

6.4 Common errors

This section describes and explains the most frequently encountered error messages.

Cannot find file filename The named file could not be found in the current directory, nor in the directories of the search path. The filename is either a compiled interface file (.cmi file), or a compiled bytecode file (.cmo file). If filename has the format mod.cmi, this means you are trying to compile a file that references identifiers from module mod, but you have not yet compiled an interface for module mod. Fix: compile mod.mli or mod.ml first, to create the compiled interface

mod.cmi.

If filename has the format mod.cmo, this means you are trying to link a bytecode object file that does not exist yet. Fix: compile mod.ml first.

If your program spans several directories, this error can also appear because you haven’t specified the directories to look into. Fix: add the correct -I options to the command line.

Corrupted compiled interface filename The compiler produces this error when it tries to read a compiled interface file (.cmi file) that has the wrong structure. This means something went wrong when this .cmi file was written: the disk was full, the compiler was interrupted in the middle of the file creation, and so on. This error can also appear if a .cmi file is modified after its creation by the compiler. Fix: remove the corrupted .cmi file, and rebuild it.

This expression has type t_{1}, but is used with type t_{2} This is by far the most common type error in programs. Type t_{1} is the type inferred for the expression (the part of the program that is displayed in the error message), by looking at the expression itself. Type t_{2} is the type expected by the context of the expression; it is deduced by looking at how the value of this expression is used in the rest of the program. If the two types t_{1} and t_{2} are not compatible, then the error above is produced.

In some cases, it is hard to understand why the two types t_{1} and t_{2} are incompatible. For instance, the compiler can report that “expression of type foo cannot be used with type foo”, and it really seems that the two types foo are compatible. This is not always true. Two type constructors can have the same name, but actually represent different types. This can happen if a type constructor is redefined. Example:

        type foo = A | B
        let f = function A -> 0 | B -> 1
        type foo = C | D
        f C

This result in the error message “expression C of type foo cannot be used with type foo”.

The type of this expression, t, contains type variables that cannot be generalized Type variables ('a, 'b, …) in a type t can be in either of two states: generalized (which means that the type t is valid for all possible instantiations of the variables) and not generalized (which means that the type t is valid only for one instantiation of the variables). In a let binding let name = expr, the type-checker normally generalizes as many type variables as possible in the type of expr. However, this leads to unsoundness (a well-typed program can crash) in conjunction with polymorphic mutable data structures. To avoid this, generalization is performed at let bindings only if the bound expression expr belongs to the class of “syntactic values”, which includes constants, identifiers, functions, tuples of syntactic values, etc. In all other cases (for instance, expr is a function application), a polymorphic mutable could have been created and generalization is therefore turned off.

Non-generalized type variables in a type cause no difficulties inside a given structure or compilation unit (the contents of a .ml file, or an interactive session), but they cannot be allowed inside signatures nor in compiled interfaces (.cmi file), because they could be used inconsistently later. Therefore, the compiler flags an error when a structure or compilation unit defines a value name whose type contains non-generalized type variables. There are two ways to fix this error:

• Add a type constraint or a .mli file to give a monomorphic type (without type variables) to name. For instance, instead of writing

        let sort_int_list = Sort.list (<)
        (* inferred type 'a list -> 'a list, with 'a not generalized *)

write

        let sort_int_list = (Sort.list (<) : int list -> int list);;

• If you really need name to have a polymorphic type, turn its defining expression into a function by adding an extra parameter. For instance, instead of writing

        let map_length = List.map Array.length
        (* inferred type 'a array list -> int list, with 'a not generalized *)

write

        let map_length lv = List.map Array.length lv

Reference to undefined global mod This error appears when trying to link an incomplete or incorrectly ordered set of files. Either you have forgotten to provide an implementation for the compilation unit named mod on the command line (typically, the file named mod.cmo, or a library containing that file). Fix: add the missing .ml or .cmo file to the command line. Or, you have provided an implementation for the module named mod, but it comes too late on the command line: the implementation of mod must come before all bytecode object files that reference mod. Fix: change the order of .ml and .cmo files on the command line.

Of course, you will always encounter this error if you have mutually recursive functions across modules. That is, function Mod1.f calls function Mod2.g, and function Mod2.g calls function Mod1.f. In this case, no matter what permutations you perform on the command line, the program will be rejected at link-time. Fixes:

• Put f and g in the same module.

• Parameterize one function by the other. That is, instead of having

mod1.ml:    let f x = ... Mod2.g ...
mod2.ml:    let g y = ... Mod1.f ...

define

mod1.ml:    let f g x = ... g ...
mod2.ml:    let rec g y = ... Mod1.f g ...
and link mod1.cmo before mod2.cmo.

• Use a reference to hold one of the two functions, as in :

mod1.ml:    let forward_g =
                ref((fun x -> failwith "forward_g") : <type>)
            let f x = ... !forward_g ...
mod2.ml:    let g y = ... Mod1.f ...
            let _= Mod1.forward_g := g

This will not work if g is a polymorphic function, however.

The external function f is not available This error appears when trying to link code that calls external functions written in C in “default runtime” mode. As explained in chapter 14, such code must be linked in “custom runtime” mode. Fix: add the -custom option, as well as the C libraries and C object files that implement the required external functions.

96

The toplevel system (ocaml)

This chapter describes the toplevel system for Objective Caml, that permits interactive use of the Objective Caml system through a read-eval-print loop. In this mode, the system repeatedly reads Caml phrases from the input, then typechecks, compile and evaluate them, then prints the inferred type and result value, if any. The system prints a # (sharp) prompt before reading each phrase.

Input to the toplevel can span several lines. It is terminated by ;; (a double-semicolon). The toplevel input consists in one or several toplevel phrases, with the following syntax:

toplevel-input ::= {toplevel-phrase} ;;
toplevel-phrase ::= definition
| expr
| # ident directive-argument
definition ::= let [rec] let-binding {and let-binding}
| external value-name : typexpr = external-declaration
| type-definition
| exception-definition
| module module-name [: module-type] = module-expr
| moduletype modtype-name = module-type
| open module-path
directive-argument ::= nothing
| string-literal
| integer-literal
| value-path

A phrase can consist of a definition, similar to those found in implementations of compilation units or in struct … end module expressions. The definition can bind value names, type names, an exception, a module name, or a module type name. The toplevel system performs the bindings, then prints the types and values (if any) for the names thus defined.

A phrase may also consist in a open directive (see section 4.11), or a value expression (section 4.7). Expressions are simply evaluated, without performing any bindings, and the value of the expression is printed.

Finally, a phrase can also consist in a toplevel directive, starting with # (the sharp sign). These directives control the behavior of the toplevel; they are listed below in section 7.2.

Unix:

The toplevel system is started by the command ocaml, as follows:

        ocaml options  # interactive mode
        ocaml options scriptfile  # script mode

If no filename is given on the command line, the toplevel system enters interactive mode: phrases are read on standard input, results are printed on standard output, errors on standard error. End-of-file on standard input terminates ocaml (see also the #quit directive in section 7.2).

On start-up (before the first phrase is read), if the file .ocamlinit exists in the current directory, its contents are read as a sequence of Objective Caml phrases and executed as per the #use directive described in section 7.2. The evaluation outcode for each phrase are not displayed.

The toplevel system does not perform line editing, but it can easily be used in conjunction with an external line editor such as fep; just run fep -emacs ocaml or fep -vi ocaml. Another option is to use ocaml under Gnu Emacs, which gives the full editing power of Emacs (see the subdirectory emacs of the Objective Caml distribution).

At any point, the parsing, compilation or evaluation of the current phrase can be interrupted by pressing ctrl-C (or, more precisely, by sending the sigintr signal to the ocaml process). The toplevel then immediately returns to the # prompt.

If a filename is given on the command-line to ocaml, the toplevel system enters script mode: the contents of the file are read as a sequence of Objective Caml phrases and executed, as per the #use directive (section 7.2). The outcome of the evaluation is not printed. On reaching the end of file, the ocaml command exits immediately. No commands are read from standard input.

In script mode, the first line of the script is ignored if it starts with #!. Thus, it is theoretically possible to make the script itself executable and put as first line #!/usr/local/bin/ocaml, thus calling the toplevel system automatically when the script is run. However, ocaml itself is a #! script on most installations of Objective Caml, and Unix kernels usually do not handle nested #! scripts.

Windows:

In addition to the text-only command ocaml.exe, which works exactly as under Unix (see above), a graphical user interface for the toplevel is available under the name ocamlwin.exe. It should be launched from the Windows file manager or program manager.

The “Terminal” windows is split in two panes. Phrases are entered and edited in the bottom pane. The top pane displays a copy of the input phrases as they are processed by the Caml Light toplevel, interspersed with the toplevel responses. The “Return” key sends the contents of the bottom pane to the Caml Light toplevel. The “Enter” key inserts a newline without sending the contents of the Input window. (This can be configured with the “Preferences” menu item.)

The contents of the input window can be edited at all times, with the standard Windows interface. An history of previously entered phrases is maintained and displayed in a separate window.

To quit the Camlwin application, either select “Quit” from the “File” menu, or use the quit function described below.

At any point, the parsing, compilation or evaluation of the current phrase can be interrupted by selecting the “Interrupt Caml Light” menu item. This goes back to the # prompt.

7.1 Options

The following command-line options are recognized by the ocaml command.

-I directory Add the given directory to the list of directories searched for source and compiled files. By default, the current directory is searched first, then the standard library directory. Directories added with -I are searched after the current directory, in the order in which they were given on the command line, but before the standard library directory.

Directories can also be added to the search path once the toplevel is running with the #directory directive (section 7.2).

-unsafe

See the corresponding option for ocamlc, chapter 6. Turn bound checking off on array and string accesses (the v.(i) and s.[i] constructs). Programs compiled with -unsafe are therefore slightly faster, but unsafe: anything can happen if the program accesses an array or string outside of its bounds.

Unix:

The following environment variables are also consulted:

LC_CTYPE

If set to iso_8859_1, accented characters (from the ISO Latin-1 character set) in string and character literals are printed as is; otherwise, they are printed as decimal escape sequences (\ddd).

TERM

When printing error messages, the toplevel system attempts to underline visually the location of the error. It consults the TERM variable to determines the type of output terminal and look up its capabilities in the terminal database.

7.2 Toplevel directives

The following directives control the toplevel behavior, load files in memory, and trace program execution.

#quit;;

Exit the toplevel loop and terminate the ocaml command.

#directory "dir-name";;

Add the given directory to the list of directories searched for source and compiled files.

#cd "dir-name";; Change the current working directory.

#load "file-name";;

Load in memory a bytecode object file (.cmo file) produced by the batch compiler ocamlc.

#use "file-name";;

Read, compile and execute source phrases from the given file. This is textual inclusion: phrases are processed just as if they were typed on standard input. The reading of the file stops at the first error encountered.

#install_printer printer-name;;

This directive registers the function named printer-name (a value path) as a printer for objects whose types match the argument type of the function. That is, the toplevel loop will call printer-name when it has such an object to print. The printing function printer-name must use the Format library module to produce its output, otherwise its output will not be correctly located in the values printed by the toplevel loop.

#remove_printer printer-name;;

Remove the named function from the table of toplevel printers.

#trace function-name;; After executing this directive, all calls to the function named function-name will be “traced”. That is, the argument and the result are displayed for each call, as well as the exceptions escaping out of the function, raised either by the function itself or by another function it calls. If the function is curried, each argument is printed as it is passed to the function.

#untrace function-name;; Stop tracing the given function.

#untrace_all;;

Stop tracing all functions traced so far.

#print_depth n;;

Limit the printing of values to a maximal depth of n. The parts of values whose depth exceeds n are printed as ... (ellipsis).

#print_length n;;

Limit the number of value nodes printed to at most n. Remaining parts of values are printed as ... (ellipsis).

7.3 The toplevel and the module system

Toplevel phrases can refer to identifiers defined in compilation units with the same mechanisms as for separately compiled units: either by using qualified names (Modulename.localname), or by using the open construct and unqualified names (see section 4.3).

However, before referencing another compilation unit, an implementation of that unit must be present in memory. At start-up, the toplevel system contains implementations for all the modules in the the standard library. Implementations for user modules can be entered with the #load directive described above. Referencing a unit for which no implementation has been provided results in the error “Reference to undefined global ‘… ’ ”.

Note that entering open mod merely accesses the compiled interface (.cmi file) for mod, but does not load the implementation of mod, and does not cause any error if no implementation of mod has been loaded. The error “reference to undefined global mod” will occur only when executing a value or module definition that refers to mod.

7.4 Common errors

This section describes and explains the most frequently encountered error messages.

Cannot find file filename The named file could not be found in the current directory, nor in the directories of the search path.

If filename has the format mod.cmi, this means you have referenced the compilation unit mod, but its compiled interface could not be found. Fix: compile mod.mli or mod.ml first, to create the compiled interface mod.cmi.

If filename has the format mod.cmo, this means you are trying to load with #load a bytecode object file that does not exist yet. Fix: compile mod.ml first.

If your program spans several directories, this error can also appear because you haven’t specified the directories to look into. Fix: use the #directory directive to add the correct directories to the search path.

This expression has type t_{1}, but is used with type t_{2} See section 6.4.

Reference to undefined global mod You have neglected to load in memory an implementation for a module with #load. See section 7.3 above.

7.5 Building custom toplevel systems: ocamlmktop

The ocamlmktop command builds Objective Caml toplevels that contain user code preloaded at start-up.

The ocamlmktop command takes as argument a set of .cmo and .cma files, and links them with the object files that implement the Objective Caml toplevel. The typical use is:

        ocamlmktop -o mytoplevel foo.cmo bar.cmo gee.cmo

This creates the bytecode file mytoplevel, containing the Objective Caml toplevel system, plus the code from the three .cmo files. This toplevel is directly executable and is started by:

        ./mytoplevel

This enters a regular toplevel loop, except that the code from foo.cmo, bar.cmo and gee.cmo is already loaded in memory, just as if you had typed:

        #load "foo.cmo";;
        #load "bar.cmo";;
        #load "gee.cmo";;

on entrance to the toplevel. The modules Foo, Bar and Gee are not opened, though; you still have to do

        open Foo;;

yourself, if this is what you wish.

7.6 Options

The following command-line options are recognized by ocamlmktop.

-cclib libname Pass the -llibname option to the C linker when linking in “custom runtime” mode. See the corresponding option for ocamlc, in chapter 6.

-ccopt option

Pass the given option to the C compiler and linker, when linking in “custom runtime” mode. See the corresponding option for ocamlc, in chapter 6.

-custom

Link in “custom runtime” mode. See the corresponding option for ocamlc, in chapter 6.

-I directory Add the given directory to the list of directories searched for compiled object code files (.cmo and .cma).

-o exec-file Specify the name of the toplevel file produced by the linker. The default is a.out.

The runtime system (ocamlrun)

The ocamlrun command executes bytecode files produced by the linking phase of the ocamlc command.

8.1 Overview

The ocamlrun command comprises three main parts: the bytecode interpreter, that actually executes bytecode files; the memory allocator and garbage collector; and a set of C functions that implement primitive operations such as input/output.

The usage for ocamlrun is:

ocamlrun options bytecode-executable arg_{1} ... arg_{n}

The first non-option argument is taken to be the name of the file containing the executable bytecode. (That file is searched in the executable path as well as in the current directory.) The remaining arguments are passed to the Caml Light program, in the string array Sys.argv. Element 0 of this array is the name of the bytecode executable file; elements 1 to n are the remaining arguments arg_{1} to arg_{n}.

As mentioned in chapter 6, in most cases, the bytecode executable files produced by the ocamlc command are self-executable, and manage to launch the ocamlrun command on themselves automatically. That is, assuming caml.out is a bytecode executable file,

        caml.out arg_{1} ... arg_{n}

works exactly as

        ocamlrun caml.out arg_{1} ... arg_{n}

Notice that it is not possible to pass options to ocamlrun when invoking caml.out directly.

8.2 Options

The following command-line option is recognized by ocamlrun.

-v When set, the memory manager prints verbose messages on standard error to signal garbage collections and heap extensions.

The following environment variable are also consulted:

CAMLRUNPARAM

Set the garbage collection parameters. This variable must be a sequence of parameter specifications. A parameter specification is an option letter followed by an = sign, a decimal number, and an optional multiplier. There are seven options, the first six correspond to the fields of the control record documented in section 16.8:

s (minor_heap_size) Size of the minor heap. i (major_heap_increment) Minimum size increment for the major heap. o (space_overhead) The major GC speed setting. O (max_overhead) The heap compaction trigger setting. v (verbose) Whether to print GC messages or not. 0 is false; 1 is true; other values may

give unexpected results. l (stack_limit) The limit (in words) of the stack size. h The initial size of the major heap (in words).

The multiplier is k, M, or G, for multiplication by 2^{10}, 2^{20}, and 2^{30} respectively. For example, on a 32-bit machine, under bash the command

        export CAMLRUNPARAM='s=256k,v=1'

tells a subsequent ocamlrun to set its initial minor heap size to 1 megabyte and to print its GC messages.

PATH

List of directories searched to find the bytecode executable file.

8.3 Common errors

This section describes and explains the most frequently encountered error messages.

filename: no such file or directory

If filename is the name of a self-executable bytecode file, this means that either that file does not exist, or that it failed to run the ocamlrun bytecode interpreter on itself. The second possibility indicates that Objective Caml has not been properly installed on your system.

Cannot exec camlrun

(When launching a self-executable bytecode file.) The ocamlrun could not be found in the executable path. Check that Objective Caml has been properly installed on your system.

Cannot find the bytecode file

The file that ocamlrun is trying to execute (e.g. the file given as first non-option argument to ocamlrun) either does not exist, or is not a valid executable bytecode file.

Truncated bytecode file

The file that ocamlrun is trying to execute is not a valid executable bytecode file. Probably it has been truncated or mangled since created. Erase and rebuild it.

Uncaught exception

The program being executed contains a “stray” exception. That is, it raises an exception at some point, and this exception is never caught. This causes immediate termination of the program. The name of the exception is printed, but not its arguments.

Out of memory

The program being executed requires more memory than available. Either the program builds excessively large data structures; or the program contains too many nested function calls, and the stack overflows. In some cases, your program is perfectly correct, it just requires more memory than your machine provides. In other cases, the “out of memory” message reveals an error in your program: non-terminating recursive function, allocation of an excessively large array or string, attempts to build an infinite list or other data structure, …

To help you diagnose this error, run your program with the -v option to ocamlrun. If it displays lots of “Growing stack… ” messages, this is probably a looping recursive function. If it displays lots of “Growing heap… ” messages, with the heap size growing slowly, this is probably an attempt to construct a data structure with too many (infinitely many?) cells. If it displays few “Growing heap… ” messages, but with a huge increment in the heap size, this is probably an attempt to build an excessively large array or string.

106

Native-code compilation (ocamlopt)

This chapter describes the Objective Caml high-performance native-code compiler ocamlopt, which compiles Caml source files to native code object files and link these object files to produce standalone executables.

The native-code compiler is only available on certain platforms. It produces code that runs faster than the bytecode produced by ocamlc, at the cost of increased compilation time and executable code size. Compatibility with the bytecode compiler is extremely high: the same source code should run identically when compiled with ocamlc and ocamlopt.

It is not possible to mix native-code object files produced by ocamlc with bytecode object files produced by ocamlopt: a program must be compiled entirely with ocamlopt or entirely with ocamlc. Native-code object files produced by ocamlopt cannot be loaded in the toplevel system

ocaml.

9.1 Overview of the compiler

The ocamlopt command has a command-line interface very close to that of ocamlc. It accepts the same types of arguments:

• Arguments ending in .mli are taken to be source files for compilation unit interfaces. Interfaces specify the names exported by compilation units: they declare value names with their types, define public data types, declare abstract data types, and so on. From the file x.mli, the ocamlopt compiler produces a compiled interface in the file x.cmi. The interface produced is identical to that produced by the bytecode compiler ocamlc.

• Arguments ending in .ml are taken to be source files for compilation unit implementations. Implementations provide definitions for the names exported by the unit, and also contain expressions to be evaluated for their side-effects. From the file x.ml, the ocamlopt compiler produces two files: x.o, containing native object code, and x.cmx, containing extra information for linking and optimization of the clients of the unit. The compiled implementation should always be referred to under the name x.cmx (when given a .o file, ocamlopt assumes that it contains code compiled from C, not from Caml).

The implementation is checked against the interface file x.mli (if it exists) as described in the manual for ocamlc (chapter 6).

• Arguments ending in .cmx are taken to be compiled object code. These files are linked together, along with the object files obtained by compiling .ml arguments (if any), and the Caml Light standard library, to produce a native-code executable program. The order in which .cmx and .ml arguments are presented on the command line is relevant: compilation units are initialized in that order at run-time, and it is a link-time error to use a component of a unit before having initialized it. Hence, a given x.cmx file must come before all .cmx files that refer to the unit x.

• Arguments ending in .cmxa are taken to be libraries of object code. Such a library packs in two files (lib.cmxa and lib.a) a set of object files (.cmx/.o files). Libraries are build with ocamlopt -a (see the description of the -a option below). The object files contained in the library are linked as regular .cmx files (see above), in the order specified when the library was built. The only difference is that if an object file contained in a library is not referenced anywhere in the program, then it is not linked in.

• Arguments ending in .c are passed to the C compiler, which generates a .o object file. This object file is linked with the program.

• Arguments ending in .o or .a are assumed to be C object files and libraries. They are linked with the program.

The output of the linking phase is a regular Unix executable file. It does not need ocamlrun to run.

9.2 Options

The following command-line options are recognized by ocamlopt.

-a Build a library (.cmxa/.a file) with the object files (.cmx/.o files) given on the command line, instead of linking them into an executable file. The name of the library can be set with the -o option. The default name is library.cmxa.

-c Compile only. Suppress the linking phase of the compilation. Source code files are turned into compiled files, but no executable file is produced. This option is useful to compile modules separately.

-cclib -llibname

Pass the -llibname option to the linker. This causes the given C library to be linked with the program.

-ccopt option

Pass the given option to the C compiler and linker. For instance, -ccopt -Ldir causes the C linker to search for C libraries in directory dir.

-compact

Optimize the produced code for space rather than for time. This results in slightly smaller but slightly slower programs. The default is to optimize for speed.

-i Cause the compiler to print all defined names (with their inferred types or their definitions) when compiling an implementation (.ml file). This can be useful to check the types inferred by the compiler. Also, since the output follows the syntax of interfaces, it can help in writing an explicit interface (.mli file) for a file: just redirect the standard output of the compiler to a .mli file, and edit that file to remove all declarations of unexported names.

-I directory Add the given directory to the list of directories searched for compiled interface files (.cmi) and compiled object code files (.cmx). By default, the current directory is searched first, then the standard library directory. Directories added with -I are searched after the current directory, in the order in which they were given on the command line, but before the standard library directory.

-inline n

Set aggressiveness of inlining to n, where n is a positive integer. Specifying -inline 0 prevents all functions from being inlined, except those whose body is smaller than the call site. Thus, inlining causes no expansion in code size. The default aggressiveness, -inline 1, allows slightly larger functions to be inlined, resulting in a slight expansion in code size. Higher values for the -inline option cause larger and larger functions to become candidate for inlining, but can result in a serious increase in code size.

-linkall

Forces all modules contained in libraries to be linked in. If this flag is not given, unreferenced modules are not linked in. When building a library (-a flag), setting the -linkall flag forces all subsequent links of programs involving that library to link all the modules contained in the library.

-o exec-file Specify the name of the output file produced by the linker. The default output name is a.out, in keeping with the Unix tradition. If the -a option is given, specify the name of the library produced. If the -output-obj option is given, specify the name of the output file produced.

-output-obj

Cause the linker to produce a C object file instead of an executable file. This is useful to wrap Caml code as a C library, callable from any C program. See chapter 14, section 14.6.5. The name of the output object file is camlprog.o by default; it can be set with the -o option.

-pp command Cause the compiler to call the given command as a preprocessor for each source file. The output of command is redirected to an intermediate file, which is compiled. If there are no compilation errors, the intermediate file is deleted afterwards. The name of this file is built from the basename of the source file with the extension .ppi for an interface (.mli) file and .ppo for an implementation (.ml) file.

-S Keep the assembly code produced during the compilation. The assembly code for the source file x.ml is saved in the file x.s.

-thread

Compile or link multithreaded programs, in combination with the threads library described in chapter 20. What this option actually does is select a special, thread-safe version of the standard library.

-unsafe

Turn bound checking off on array and string accesses (the v.(i) and s.[i] constructs). Programs compiled with -unsafe are therefore faster, but unsafe: anything can happen if the program accesses an array or string outside of its bounds.

-v Print the version number of the compiler.

9.3 Common errors

The error messages are almost identical to those of ocamlc. See section 6.4.

9.4 Compatibility with the bytecode compiler

This section lists the known incompatibilities between the bytecode compiler and the native-code compiler. Except on those points, the two compilers should generate code that behave identically.

• The following operations abort the program (either by printing an error message or just via an hardware trap or fatal Unix signal) instead of raising an exception:

– out-of bounds accesses to arrays and strings;

– integer division by zero, modulus by zero;

– stack overflow;

– on the Alpha processor only, floating-point operations involving infinite or denormalized numbers (all other processors supported by ocamlopt treat these numbers correctly, as per the IEEE 754 standard).

In particular, notice that stack overflow caused by excessively deep recursion is reported by most Unix kernels as a “segmentation violation” signal.

• The following library functions print a fatal error message and abort the program instead of raising the Invalid_argument exception:

– structural comparisons (=), (<>), etc., when encountering a functional value;

– Array.create and String.create when the requested size is negative or exceeds the memory manager limits;

– output_value when encountering a functional value or pointers outside the Caml heap (such as input-output buffers).

• Signals are detected only when the program performs an allocation in the heap. That is, if a signal is delivered while in a piece of code that does not allocate, its handler will not be called until the next heap allocation.

The best way to avoid running into those incompatibilities is to never trap the

Invalid_argument, Division_by_zero, and Stack_overflow exceptions, thus also treat-

ing them as fatal errors with the bytecode compiler as well as with the native-code compiler. Test the divisor or array/string index before performing the operation instead of trapping the exception afterwards.

112

Lexer and parser generators (ocamllex, ocamlyacc)

This chapter describes two program generators: ocamllex, that produces a lexical analyzer from a set of regular expressions with associated semantic actions, and ocamlyacc, that produces a parser from a grammar with associated semantic actions.

These program generators are very close to the well-known lex and yacc commands that can be found in most C programming environments. This chapter assumes a working knowledge of lex and yacc: while it describes the input syntax for ocamllex and ocamlyacc and the main differences with lex and yacc, it does not explain the basics of writing a lexer or parser description in lex and yacc. Readers unfamiliar with lex and yacc are referred to “Compilers: principles, techniques, and tools” by Aho, Sethi and Ullman (Addison-Wesley, 1986), or “Lex & Yacc”, by Levine, Mason and Brown (O’Reilly, 1992).

10.1 Overview of ocamllex

The ocamllex command produces a lexical analyzer from a set of regular expressions with attached semantic actions, in the style of lex. Assuming the input file is lexer.mll, executing

        ocamllex lexer.mll

produces Caml code for a lexical analyzer in file lexer.ml. This file defines one lexing function per entry point in the lexer definition. These functions have the same names as the entry points. Lexing functions take as argument a lexer buffer, and return the semantic attribute of the corresponding entry point.

Lexer buffers are an abstract data type implemented in the standard library module Lexing.

The functions Lexing.from_channel, Lexing.from_string and Lexing.from_function create

lexer buffers that read from an input channel, a character string, or any reading function, respectively. (See the description of module Lexing in chapter 15.)

When used in conjunction with a parser generated by ocamlyacc, the semantic actions compute a value belonging to the type token defined by the generated parsing module. (See the description

of ocamlyacc below.)

10.2 Syntax of lexer definitions

The format of lexer definitions is as follows:

{ header } rule entrypoint =

parse regexp { action } | … | regexp { action }

and entrypoint =

  parse . . .
and . . .

{ trailer }

Comments are delimited by (* and *), as in Caml.

10.2.1 Header and trailer

The header and trailer sections are arbitrary Caml text enclosed in curly braces. Either or both can be omitted. If present, the header text is copied as is at the beginning of the output file and the trailer text at the end. Typically, the header section contains the open directives required by the actions, and possibly some auxiliary functions used in the actions.

10.2.2 Entry points

The names of the entry points must be valid value identifiers.

10.2.3 Regular expressions

The regular expressions are in the style of lex, with a more Caml-like syntax.

' char '

A character constant, with the same syntax as Objective Caml character constants. Match the denoted character.

_(Underscore.) Match any character.

eof Match the end of the lexer input.

Note: On some systems, with interactive input, and end-of-file may be followed by more characters. However, ocamllex will not correctly handle regular expressions that contain eof followed by something else.

" string "

A string constant, with the same syntax as Objective Caml string constants. Match the corresponding sequence of characters.

[ character-set ]

Match any single character belonging to the given character set. Valid character sets are: single character constants ' c '; ranges of characters ' c_{1} ' - ' c_{2} ' (all characters between c_{1} and c_{2}, inclusive); and the union of two or more character sets, denoted by concatenation.

[^character-set ]

Match any single character not belonging to the given character set.

regexp *

(Repetition.) Match the concatenation of zero or more strings that match regexp.

regexp +

(Strict repetition.) Match the concatenation of one or more strings that match regexp.

regexp ?

(Option.) Match either the empty string, or a string matching regexp.

regexp_{1} | regexp_{2}

(Alternative.) Match any string that matches either regexp_{1} or regexp_{2}

regexp_{1} regexp_{2}

(Concatenation.) Match the concatenation of two strings, the first matching regexp_{1}, the second matching regexp_{2}.

( regexp )

Match the same strings as regexp.

Concerning the precedences of operators, * and + have highest precedence, followed by ?, then concatenation, then | (alternation).

10.2.4 Actions

The actions are arbitrary Caml expressions. They are evaluated in a context where the identifier lexbuf is bound to the current lexer buffer. Some typical uses for lexbuf, in conjunction with the operations on lexer buffers provided by the Lexing standard library module, are listed below.

Lexing.lexeme lexbuf

Return the matched string.

Lexing.lexeme_char lexbuf n

Return the n^{th} character in the matched string. The first character corresponds to n = 0.

Lexing.lexeme_start lexbuf

Return the absolute position in the input text of the beginning of the matched string. The first character read from the input text has position 0.

Lexing.lexeme_end lexbuf

Return the absolute position in the input text of the end of the matched string. The first character read from the input text has position 0.

entrypoint lexbuf (Where entrypoint is the name of another entry point in the same lexer definition.) Recursively call the lexer on the given entry point. Useful for lexing nested comments, for example.

10.3 Overview of ocamlyacc

The ocamlyacc command produces a parser from a context-free grammar specification with attached semantic actions, in the style of yacc. Assuming the input file is grammar.mly, executing

ocamlyacc options grammar.mly

produces Caml code for a parser in the file grammar.ml, and its interface in file grammar.mli.

The generated module defines one parsing function per entry point in the grammar. These functions have the same names as the entry points. Parsing functions take as arguments a lexical analyzer (a function from lexer buffers to tokens) and a lexer buffer, and return the semantic attribute of the corresponding entry point. Lexical analyzer functions are usually generated from a lexer specification by the ocamllex program. Lexer buffers are an abstract data type implemented in the standard library module Lexing. Tokens are values from the concrete type token, defined in the interface file grammar.mli produced by ocamlyacc.

10.4 Syntax of grammar definitions

Grammar definitions have the following format:

%{

header

%}

declarations

%%

rules

%%

trailer

Comments are enclosed between /* and */ (as in C) in the “declarations” and “rules” sections, and between (* and *) (as in Caml) in the “header” and “trailer” sections.

10.4.1 Header and trailer

The header and the trailer sections are Caml code that is copied as is into file grammar.ml. Both sections are optional. The header goes at the beginning of the output file; it usually contains open directives and auxiliary functions required by the semantic actions of the rules. The trailer goes at the end of the output file.

10.4.2 Declarations

Declarations are given one per line. They all start with a % sign.

%token symbol … symbol

Declare the given symbols as tokens (terminal symbols). These symbols are added as constant constructors for the token concrete type.

%token< type > symbol … symbol

Declare the given symbols as tokens with an attached attribute of the given type. These symbols are added as constructors with arguments of the given type for the token concrete type. The type part is an arbitrary Caml type expression, except that all type constructor names must be fully qualified (e.g. Modname.typename) for all types except standard builtin types, even if the proper open directives (e.g. open Modname) were given in the header section. That’s because the header is copied only to the .ml output file, but not to the .mli output file, while the type part of a %token declaration is copied to both.

%start symbol … symbol

Declare the given symbols as entry points for the grammar. For each entry point, a parsing function with the same name is defined in the output module. Non-terminals that are not declared as entry points have no such parsing function. Start symbols must be given a type with the %type directive below.

%type< type > symbol … symbol

Specify the type of the semantic attributes for the given symbols. This is mandatory for start symbols only. Other nonterminal symbols need not be given types by hand: these types will be inferred when running the output files through the Objective Caml compiler (unless the -s option is in effect). The type part is an arbitrary Caml type expression, except that all type constructor names must be fully qualified, as explained above for %token.

%left symbol … symbol

%right symbol … symbol

%nonassoc symbol … symbol

Associate precedences and associativities to the given symbols. All symbols on the same line are given the same precedence. They have higher precedence than symbols declared before in a %left, %right or %nonassoc line. They have lower precedence than symbols declared after in a %left, %right or %nonassoc line. The symbols are declared to associate to the left (%left), to the right (%right), or to be non-associative (%nonassoc). The symbols are usually tokens. They can also be dummy nonterminals, for use with the %prec directive inside the rules.

10.4.3 Rules

The syntax for rules is as usual:

nonterminal :

symbol … symbol { semantic-action } | … | symbol … symbol { semantic-action }

;

Rules can also contain the %prec symbol directive in the right-hand side part, to override the default precedence and associativity of the rule with the precedence and associativity of the given symbol.

Semantic actions are arbitrary Caml expressions, that are evaluated to produce the semantic attribute attached to the defined nonterminal. The semantic actions can access the semantic attributes of the symbols in the right-hand side of the rule with the $ notation: $1 is the attribute for the first (leftmost) symbol, $2 is the attribute for the second symbol, etc.

The rules may contain the special symbol error to indicate resynchronization points, as in

yacc.

Actions occurring in the middle of rules are not supported.

10.4.4 Error handling

Error recovery is supported as follows: when the parser reaches an error state (no grammar rules can apply), it calls a function named parse_error with the string syntax error as argument. The default parse_error function does nothing and returns, thus initiating error recovery (see below). The user can define a customized parse_error function in the header section of the grammar file.

The parser also enters error recovery mode if one of the grammar actions raise the

Parsing.Parse_error exception.

In error recovery mode, the parser discards states from the stack until it reaches a place where the error token can be shifted. It then discards tokens from the input until it finds three successive tokens that can be accepted, and starts processing with the first of these. If no state can be uncovered where the error token can be shifted, then the parser aborts by raising the

Parsing.Parse_error exception.

Refer to documentation on yacc for more details and guidance in how to use error recovery.

10.5 Options

The ocamlyacc command recognizes the following options:

-v Generate a description of the parsing tables and a report on conflicts resulting from ambiguities in the grammar. The description is put in file grammar.output.

-bprefix Name the output files prefix.ml, prefix.mli, prefix.output, instead of the default naming convention.

10.6 A complete example

The all-time favorite: a desk calculator. This program reads arithmetic expressions on standard input, one per line, and prints their values. Here is the grammar definition:

        /* File parser.mly */
        %token <int> INT
        %token PLUS MINUS TIMES DIV
        %token LPAREN RPAREN

%token EOL
%left PLUS MINUS        /* lowest precedence */
%left TIMES DIV         /* medium precedence */
%nonassoc UMINUS        /* highest precedence */
%start main             /* the entry point */
%type <int> main
%%
main:
    expr EOL                { $1 }
;
expr:
    INT                     { $1 }
  | LPAREN expr RPAREN      { $2 }
  | expr PLUS expr          { $1 + $3 }
  | expr MINUS expr         { $1 - $3 }
  | expr TIMES expr         { $1 * $3 }
  | expr DIV expr           { $1 / $3 }
  | MINUS expr %prec UMINUS { - $2 }
;

Here is the definition for the corresponding lexer:

(* File lexer.mll *)
{
open Parser        (* The type token is defined in parser.mli *)
exception Eof
}
rule token = parse
    [' ' '\t']     { token lexbuf }     (* skip blanks *)
  | ['\n' ]        { EOL }
  | ['0'-'9']+     { INT(int_of_string(Lexing.lexeme lexbuf)) }
  | '+'            { PLUS }
  | '-'            { MINUS }
  | '*'            { TIMES }
  | '/'            { DIV }
  | '('            { LPAREN }
  | ')'            { RPAREN }
  | eof            { raise Eof }

Here is the main program, that combines the parser with the lexer:

(* File calc.ml *)
let _=
  try
    let lexbuf = Lexing.from_channel stdin in
    while true do
      let result = Parser.main Lexer.token lexbuf in

        print_int result; print_newline(); flush stdout
    done
  with Lexer.Eof ->
    exit 0

To compile everything, execute:

ocamllex lexer.mll       # generates lexer.ml
ocamlyacc parser.mly     # generates parser.ml and parser.mli
ocamlc -c parser.mli
ocamlc -c lexer.ml
ocamlc -c parser.ml
ocamlc -c calc.ml
ocamlc -o calc lexer.cmo parser.cmo calc.cmo

Dependency generator (ocamldep)

The ocamldep command scans a set of Objective Caml source files (.ml and .mli files) for references to external compilation units, and outputs dependency lines in a format suitable for the make utility. This ensures that make will compile the source files in the correct order, and recompile those files that need to when a source file is modified.

The typical usage is:

        ocamldep options *.mli *.ml > .depend

where *.mli *.ml expands to all source files in the current directory and .depend is the file that should contain the dependencies. (See below for a typical Makefile.)

Dependencies are generated both for compiling with the bytecode compiler ocamlc and with the native-code compiler ocamlopt.

11.1 Options

The following command-line option is recognized by ocamldep.

-I directory Add the given directory to the list of directories searched for source files. If a source file foo.ml mentions an external compilation unit Bar, a dependency on that unit’s interface bar.cmi is generated only if the source for bar is found in the current directory or in one of the directories specified with -I. Otherwise, Bar is assumed to be a module form the standard library, and no dependencies are generated. For programs that span multiple directories, it is recommended to pass ocamldep the same -I options that are passed to the compiler.

11.2 A typical Makefile

Here is a template Makefile for a Objective Caml program.

CSLC=ocamlc
CSLOPT=ocamlopt
CSLDEP=ocamldep

INCLUDES=               #all relevant -I options here
CSLFLAGS=$(INCLUDES)    #add other options for ocamlc here
CSLOPTFLAGS=$(INCLUDES) #add other options for ocamlopt here
# prog1 should be compiled to bytecode, and is composed of three
# units: mod1, mod2 and mod3.
# The list of object files for prog1
PROG1_OBJS=mod1.cmo mod2.cmo mod3.cmo
prog1: $(PROG1_OBJS)
$(CSLC) -o prog1 $(CSLFLAGS) $(PROG1_OBJS)
# prog2 should be compiled to native-code, and is composed of two
# units: mod4 and mod5.
# The list of object files for prog2
PROG2_OBJS=mod4.cmx mod5.cmx
prog2: $(PROG2_OBJS)
$(CSLOPT) -o prog2 $(CSLFLAGS) $(PROG2_OBJS)
# Common rules
.SUFFIXES: .ml .mli .cmo .cmi .cmx
.ml.cmo:
$(CSLC) $(CSLFLAGS) -c $<
.mli.cmi:
$(CSLC) $(CSLFLAGS) -c $<
.ml.cmx:
$(CSLOPT) $(CSLOPTFLAGS) -c $<
# Clean up
clean:
rm -f prog1 prog2
rm -f *.cm[iox]
# Dependencies
depend:
        $(CSLDEP) $(INCLUDES) *.mli *.ml > .depend
include .depend

The debugger (ocamldebug)

This chapter describes the Objective Caml source-level replay debugger ocamldebug.

Unix:

The debugger is available on Unix systems that provides BSD sockets.

Windows:

The debugger is not available.

Mac:

The debugger is not available.

12.1 Compiling for debugging

Before the debugger can be used, the program must be compiled and linked with the -g option: all .cmo and .cma files that are part of the program should have been created with ocamlc -g, and they must be linked together with ocamlc -g.

Compiling with -g entails no penalty on the running time of programs: object files and bytecode executable files are bigger and take longer to produce, but the executable files run at exactly the same speed as if they had been compiled without -g.

12.2 Invocation

12.2.1 Starting the debugger

The Objective Caml debugger is invoked by running the program ocamldebug with the name of the bytecode executable file as first argument:

ocamldebug [options] program [arguments]

The arguments following program are optional, and are passed as command-line arguments to the program being debugged. (See also the set arguments command.)

The following command-line options are recognized:

-I directory Add directory to the list of directories searched for source files and compiled files. (See also the directory command.)

-s socket Use socket for communicating with the debugged program. See the description of the command set socket (section 12.8.6) for the format of socket.

-c count Set the maximum number of simultaneously live checkpoints to count.

-cd directory Run the debugger program from the working directory directory, instead of the current directory. (See also the cd command.)

-emacs

Tell the debugger it is executed under Emacs. (See section 12.10 for information on how to run the debugger under Emacs.)

12.2.2 Exiting the debugger

The command quit exits the debugger. You can also exit the debugger by typing an end-of-file character (usually ctrl-D).

Typing an interrupt character (usually ctrl-C) will not exit the debugger, but will terminate the action of any debugger command that is in progress and return to the debugger command level.

12.3 Commands

A debugger command is a single line of input. It starts with a command name, which is followed by arguments depending on this name. Examples:

        run
        goto 1000
        set arguments arg1 arg2

A command name can be truncated as long as there is no ambiguity. For instance, go 1000 is understood as goto 1000, since there are no other commands whose name starts with go. For the most frequently used commands, ambiguous abbreviations are allowed. For instance, r stands for run even though there are others commands starting with r. You can test the validity of an abbreviation using the help command.

If the previous command has been successful, a blank line (typing just RET) will repeat it.

12.3.1 Getting help

The Objective Caml debugger has a simple on-line help system, which gives a brief description of each command and variable.

help

Print the list of commands.

help command Give help about the command command.

help set variable, help show variable Give help about the variable variable. The list of all debugger variables can be obtained with

     help set.
help info topic

Give help about topic. Use help info to get a list of known topics.

12.3.2 Accessing the debugger state

set variable value Set the debugger variable variable to the value value.

show variable Print the value of the debugger variable variable.

info subject Give information about the given subject. For instance, info breakpoints will print the list of all breakpoints.

12.4 Executing a program

12.4.1 Events

Events are “interesting” locations in the source code, corresponding to the beginning or end of evaluation of “interesting” sub-expressions. Events are the unit of single-stepping (stepping goes to the next or previous event encountered in the program execution). Also, breakpoints can only be set at events. Thus, events play the role of line numbers in debuggers for conventional languages.

During program execution, a counter is incremented at each event encountered. The value of this counter is referred as the current time. Thanks to reverse execution, it is possible to jump back and forth to any time of the execution.

Here is where the debugger events (written ▷◁) are located in the source code:

• Following a function application:

     (f arg)▷◁

• On entrance to a function:

     fun x y z -> ▷◁ ...

• On each case of a pattern-matching definition (function, match… with construct, try… with construct):

     function pat1 -> ▷◁ expr1
            | ...
            | patN -> ▷◁ exprN

• Between subexpressions of a sequence:

     expr1; ▷◁ expr2; ▷◁ ...; ▷◁ exprN

• In the two branches of a conditional expression:

     if cond then ▷◁ expr1 else ▷◁ expr2

• At the beginning of each iteration of a loop:

     while cond do ▷◁ body done
     for i = a to b do ▷◁ body done

Exceptions: A function application followed by a function return is replaced by the compiler by a jump (tail-call optimization). In this case, no event is put after the function application.

12.4.2 Starting the debugged program

The debugger starts executing the debugged program only when needed. This allows setting breapoints or assigning debugger variables before execution starts. There are several ways to start execution:

run Run the program until a breakpoint is hit, or the program terminates.

step 0

Load the program and stop on the first event.

goto time

Load the program and execute it until the given time. Useful when you already know approximately at what time the problem appears. Also useful to set breakpoints on function values that have not been computed at time 0 (see section 12.5).

The execution of a program is affected by certain information it receives when the debugger starts it, such as the command-line arguments to the program and its working directory. The debugger provides commands to specify this information (set arguments and cd). These commands must be used before program execution starts. If you try to change the arguments or the working directory after starting your program, the debugger will kill the program (after asking for confirmation).

12.4.3 Running the program

The following commands execute the program forward or backward, starting at the current time. The execution will stop either when specified by the command or when a breakpoint is encountered.

run Execute the program forward from current time. Stops at next breakpoint or when the program terminates.

reverse

Execute the program backward from current time. Mostly useful to go to the last breakpoint encountered before the current time.

step [count] Run the program and stop at the next event. With an argument, do it count times.

backstep [count]

Run the program backward and stop at the previous event. With an argument, do it count times.

next [count] Run the program and stop at the next event, skipping over function calls. With an argument, do it count times.

previous [count]

Run the program backward and stop at the previous event, skipping over function calls. With an argument, do it count times.

finish

Run the program until the current function returns.

start

Run the program backward and stop at the first event before the current function invocation.

12.4.4 Time travel

You can jump directly to a given time, without stopping on breakpoints, using the goto command.

As you move through the program, the debugger maintains an history of the successive times you stop at. The last command can be used to revisit these times: each last command moves one step back through the history. That is useful mainly to undo commands such as step and next.

goto time

Jump to the given time.

last [count] Go back to the latest time recorded in the execution history. With an argument, do it count times.

set history size

Set the size of the execution history.

12.4.5 Killing the program

kill

Kill the program being executed. This command is mainly useful if you wish to recompile the program without leaving the debugger.

12.5 Breakpoints

A breakpoint causes the program to stop whenever a certain point in the program is reached. It can be set in several ways using the break command. Breakpoints are assigned numbers when set, for further reference. The most comfortable way to set breakpoints is through the Emacs interface (see section 12.10).

break

Set a breakpoint at the current position in the program execution. The current position must be on an event (i.e., neither at the beginning, nor at the end of the program).

break function Set a breakpoint at the beginning of function. This works only when the functional value of the identifier function has been computed and assigned to the identifier. Hence this command cannot be used at the very beginning of the program execution, when all identifiers are still undefined; use goto time to advance execution until the functional value is available.

break @ [module] line Set a breakpoint in module module (or in the current module if module is not given), at the first event of line line.

break @ [module] line column Set a breakpoint in module module (or in the current module if module is not given), at the event closest to line line, column column.

break @ [module] # character Set a breakpoint in module module at the event closest to character number character.

break address Set a breakpoint at the code address address.

delete [breakpoint-numbers] Delete the specified breakpoints. Without argument, all breakpoints are deleted (after asking for confirmation).

info breakpoints

Print the list of all breakpoints.

12.6 The call stack

Each time the program performs a function application, it saves the location of the application (the return address) in a block of data called a stack frame. The frame also contains the local variables of the caller function. All the frames are allocated in a region of memory called the call stack. The command backtrace (or bt) displays parts of the call stack.

At any time, one of the stack frames is “selected” by the debugger; several debugger commands refer implicitly to the selected frame. In particular, whenever you ask the debugger for the value of a local variable, the value is found in the selected frame. The commands frame, up and down select whichever frame you are interested in.

When the program stops, the debugger automatically selects the currently executing frame and describes it briefly as the frame command does.

frame

Describe the currently selected stack frame.

frame frame-number Select a stack frame by number and describe it. The frame currently executing when the program stopped has number 0; its caller has number 1; and so on up the call stack.

backtrace [count], bt [count] Print the call stack. This is useful to see which sequence of function calls led to the currently executing frame. With a positive argument, print only the innermost count frames. With a negative argument, print only the outermost -count frames.

up [count] Select and display the stack frame just “above” the selected frame, that is, the frame that called the selected frame. An argument says how many frames to go up.

down [count] Select and display the stack frame just “below” the selected frame, that is, the frame that was called by the selected frame. An argument says how many frames to go down.

12.7 Examining variable values

The debugger can print the current value of simple expressions. The expressions can involve program variables: all the identifiers that are in scope at the selected program point can be accessed.

Expressions that can be printed are a subset of Objective Caml expressions, as described by the following grammar:

expr ::= lowercase-ident
| {uppercase-ident .} lowercase-ident

                             | *

| $ integer
| expr . lowercase-ident
| expr .( integer )
| expr .[ integer ]
| ! expr
| ( expr )

The first two cases refer to a value identifier, either unqualified or qualified by the path to the structure that define it. * refers to the result just computed (typically, the value of a function application), and is valid only if the selected event is an “after” event (typically, a function application). $ integer refer to a previously printed value. The remaining four forms select part of an expression: respectively, a record field, an array element, a string element, and the current contents of a reference.

print variables Print the values of the given variables. print can be abbreviated as p.

display variables Same as print, but limit the depth of printing to 1. Useful to browse large data structures without printing them in full. display can be abbreviated as d.

When printing a complex expression, a name of the form $integer is automatically assigned to its value. Such names are also assigned to parts of the value that cannot be printed because the maximal printing depth is exceeded. Named values can be printed later on with the commands p $integer or d $integer. Named values are valid only as long as the program is stopped. They are forgotten as soon as the program resumes execution.

set print_depth d

Limit the printing of values to a maximal depth of d.

set print_length l

Limit the printing of values to at most l nodes printed.

12.8 Controlling the debugger

12.8.1 Setting the program name and arguments

set program file

Set the program name to file.

set arguments arguments

Give arguments as command-line arguments for the program.

A shell is used to pass the arguments to the debugged program. You can therefore use wildcards, shell variables, and file redirections inside the arguments. To debug programs that read from standard input, it is recommended to redirect their input from a file (using set arguments < input-file), otherwise input to the program and input to the debugger are not properly separated, and inputs are not properly replayed when running the program backwards.

12.8.2 How programs are loaded

The loadingmode variable controls how the program is executed.

set loadingmode direct

The program is run directly by the debugger. This is the default mode.

set loadingmode runtime

The debugger execute the Objective Caml runtime camlrun on the program. Rarely useful; moreover it prevents the debugging of programs compiled in “custom runtime” mode.

set loadingmode manual

The user starts manually the program, when asked by the debugger. Allows remote debugging (see section 12.8.6).

12.8.3 Search path for files

The debugger searches for source files and compiled interface files in a list of directories, the search path. The search path initially contains the current directory . and the standard library directory. The directory command adds directories to the path.

Whenever the search path is modified, the debugger will clear any information it may have cached about the files.

directory directorynames Add the given directories to the search path. These directories are added at the front, and will therefore be searched first.

directory

Reset the search path. This requires confirmation.

12.8.4 Working directory

Each time a program is started in the debugger, it inherits its working directory from the current working directory of the debugger. This working directory is initially whatever it inherited from its parent process (typically the shell), but you can specify a new working directory in the debugger with the cd command or the -cd command-line option.

cd directory Set the working directory for camldebug to directory.

pwd Print the working directory for camldebug.

12.8.5 Turning reverse execution on and off

In some cases, you may want to turn reverse execution off. This speeds up the program execution, and is also sometimes useful for interactive programs.

Normally, the debugger takes checkpoints of the program state from time to time. That is, it makes a copy of the current state of the program (using the Unix system call fork). If the variable checkpoints is set to off, the debugger will not take any checkpoints.

set checkpoints on/off

Select whether the debugger makes checkpoints or not.

12.8.6 Communication between the debugger and the program

The debugger communicate with the program being debugged through a Unix socket. You may need to change the socket name, for example if you need to run the debugger on a machine and your program on another.

set socket socket

Use socket for communication with the program. socket can be either a file name, or an Internet port specification host:port, where host is a host name or an Internet address in dot notation, and port is a port number on the host.

On the debugged program side, the socket name is passed either by the -D command line option to camlrun, or through the CAML_DEBUG_SOCKET environment variable.

12.8.7 Fine-tuning the debugger

Several variables enables to fine-tune the debugger. Reasonable defaults are provided, and you should normally not have to change them.

set processcount count

Set the maximum number of checkpoints to count. More checkpoints facilitate going far back in time, but use more memory and create more Unix processes.

As checkpointing is quite expensive, it must not be done too often. On the other hand, backward execution is faster when checkpoints are taken more often. In particular, backward single-stepping is more responsive when many checkpoints have been taken just before the current time. To finetune the checkpointing strategy, the debugger does not take checkpoints at the same frequency for long displacements (e.g. run) and small ones (e.g. step). The two variables bigstep and smallstep contain the number of events between two checkpoints in each case.

set bigstep count

Set the number of events between two checkpoints for long displacements.

set smallstep count

Set the number of events between two checkpoints for small displacements.

The following commands display information on checkpoints and events:

info checkpoints

Print a list of checkpoints.

info events [module]

Print the list of events in the given module (the current module, by default).

12.8.8 User-defined printers

Just as in the toplevel system (section 7.2), the user can register functions for printing values of certain types. For technical reasons, the debugger cannot call printing functions that reside in the program being debugged. The code for the printing functions must therefore be loaded explicitly in the debugger.

load_printer "file-name"

Load in the debugger the indicated .cmo or .cma object file. The file is loaded in an environment consisting only of the Objective Caml standard library plus the definitions provided by object files previously loaded using load_printer. If this file depends on other object files not yet loaded, the debugger automatically loads them if it is able to find them in the search path. The loaded file does not have direct access to the modules of the program being debugged.

install_printer printer-name

Register the function named printer-name (a value path) as a printer for objects whose types match the argument type of the function. That is, the debugger will call printer-name when it has such an object to print. The printing function printer-name must use the Format library module to produce its output, otherwise its output will not be correctly located in the values printed by the toplevel loop.

The value path printer-name must refer to one of the functions defined by the object files loaded using load_printer. It cannot reference the functions of the program being debugged.

remove_printer printer-name

Remove the named function from the table of value printers.

12.9 Miscellaneous commands

list [module] [beginning] [end] List the source of module module, from line number beginning to line number end. By default, 20 lines of the current module are displayed, starting 10 lines before the current position.

source filename Read debugger commands from the script filename.

12.10 Running the debugger under Emacs

The most user-friendly way to use the debugger is to run it under Emacs. See the file emacs/README in the distribution for information on how to load the Emacs Lisp files for Caml support.

The Caml debugger is started under Emacs by the command M-x camldebug, with argument the name of the executable file progname to debug. Communication with the debugger takes place in an Emacs buffer named *camldebug-progname*. The editing and history facilities of Shell mode are available for interacting with the debugger.

In addition, Emacs displays the source files containing the current event (the current position in the program execution) and highlights the location of the event. This display is updated synchronously with the debugger action.

The following bindings for the most common debugger commands are available in the *camldebug-progname* buffer:

C-c C-s

(command step): execute the program one step forward.

C-c C-k

(command backstep): execute the program one step backward.

C-c C-n

(command next): execute the program one step forward, skipping over function calls.

Middle mouse button (command display): display named value. $n under mouse cursor (support incremental browsing of large data structures).

C-c C-p

(command print): print value of identifier at point.

C-c C-d

(command display): display value of identifier at point.

C-c C-r

(command run): execute the program forward to next breakpoint.

C-c C-v

(command reverse): execute the program backward to latest breakpoint.

C-c C-l

(command last): go back one step in the command history.

C-c C-t

(command backtrace): display backtrace of function calls.

C-c C-f

(command finish): run forward till the current function returns.

C-c <

(command up): select the stack frame below the current frame.

C-c >

(command down): select the stack frame above the current frame.

In all buffers in Caml editing mode, the following debugger commands are also available:

C-x C-a C-b

(command break): set a breakpoint at event closest to point

C-x C-a C-p

(command print): print value of identifier at point

C-x C-a C-d

(command display): display value of identifier at point 136

Profiling (ocamlprof)

This chapter describes how the execution of Objective Caml programs can be profiled, by recording how many times functions are called, branches of conditionals are taken, …

13.1 Compiling for profiling

Before profiling an execution, the program must be compiled in profiling mode, using the ocamlcp front-end to the ocamlc compiler (see chapter 6). When compiling modules separately, ocamlcp must be used when compiling the modules (production of .cmo files), and can also be used (though this is not strictly necessary) when linking them together.

The amount of profiling information can be controlled through the -p option to ocamlcp, followed by one or several letters indicating which parts of the program should be profiled:

a all options

f function calls : a count point is set at the beginning of function bodies

i if … then … else …: count points are set in both then branch and else branch

l while, for loops: a count point is set at the beginning of the loop body

m match branches: a count point is set at the beginning of the body of each branch

t try … with … branches: a count point is set at the beginning of the body of each branch

For instance, compiling with ocamlcp -pfilm profiles function calls, if… then… else…, loops and pattern matching.

Calling ocamlcp without the -p option defaults to -p fm, meaning that only function calls and pattern matching are profiled.

13.2 Profiling an execution

Running a bytecode executable file that has been compiled with ocamlcp records the execution counts for the specified parts of the program and saves them in a file called ocamlprof.dump in the current directory.

The ocamlprof.dump file is written only if the program terminates normally (by calling exit or by falling through). It is not written if the program terminates with an uncaught exception.

If a compatible dump file already exists in the current directory, then the profiling information is accumulated in this dump file. This allows, for instance, the profiling of several executions of a program on different inputs.

13.3 Printing profiling information

The ocamlprof command produces a source listing of the program modules where execution counts have been inserted as comments. For instance,

        ocamlprof foo.ml

prints the source code for the foo module, with comments indicating how many times the functions in this module have been called. Naturally, this information is accurate only if the source file has not been modified since the profiling execution took place.

The following options are recognized by ocamlprof:

-f dumpfile Specifies an alternate dump file of profiling information

-F string Specifies an additional string to be output with profiling information. By default, ocamlprof will annotate programs with comments of the form (* n *) where n is the counter value for a profiling point. With option -F s, the annotation will be (* s n *).

13.4 Time profiling

Profiling with ocamlprof only records execution counts, not the actual time spent into each function. There is currently no way to perform time profiling on bytecode programs generated by ocamlc. On native-code programs generated by ocamlopt, the standard Unix profiler prof can be used; just add the -ccopt -p option when linking the program:

        ocamlopt -o myprog -ccopt -p other-options files
        ./myprog
        prof myprog

Function names in the output of prof have the following format:

Module-name_function-name_unique-number

Interfacing C with Objective Caml

This chapter describes how user-defined primitives, written in C, can be linked with Caml code and called from Caml functions.

14.1 Overview and compilation information

14.1.1 Declaring primitives

User primitives are declared in an implementation file or struct… end module expression using the external keyword:

external name : type = C-function-name

This defines the value name name as a function with type type that executes by calling the given C function. For instance, here is how the input primitive is declared in the standard library module

Pervasives:
        external input : in_channel -> string -> int -> int -> int
                       = "input"

Primitives with several arguments are always curried. The C function does not necessarily have the same name as the ML function.

External functions thus defined can be specified in interface files or sig… end signatures either as regular values

val name : type

thus hiding their implementation as a C function, or explicitly as “manifest” external functions

external name : type = C-function-name

The latter is slightly more efficient, as it allows clients of the module to call directly the C function instead of going through the corresponding Caml function.

14.1.2 Implementing primitives

User primitives with arity n ≤ 5 are implemented by C functions that take n arguments of type value, and return a result of type value. The type value is the type of the representations for Caml values. It encodes objects of several base types (integers, floating-point numbers, strings, …), as well as Caml data structures. The type value and the associated conversion functions and macros are described in details below. For instance, here is the declaration for the C function implementing the input primitive:

        value input(channel, buffer, offset, length)
                value channel, buffer, offset, length;
        {
         ...
        }

When the primitive function is applied in a Caml program, the C function is called with the values of the expressions to which the primitive is applied as arguments. The value returned by the function is passed back to the Caml program as the result of the function application.

User primitives with arity greater than 5 should be implemented by two C functions. The first function, to be used in conjunction with the bytecode compiler ocamlc, receives two arguments: a pointer to an array of Caml values (the values for the arguments), and an integer which is the number of arguments provided. The other function, to be used in conjunction with the native-code compiler ocamlopt, takes its arguments directly. For instance, here are the two C functions for the 7-argument primitive Nat.add_nat:

        value add_nat_native(nat1, ofs1, len1, nat2, ofs2, len2, carry_in)
             value nat1, ofs1, len1, nat2, ofs2, len2, carry_in;
        {
          ...
        }
        value add_nat_bytecode(argv, argn)
             value * argv;
             int argn;
        {
          return add_nat_native(argv[0], argv[1], argv[2], argv[3],
                                argv[4], argv[5], argv[6]);
        }

The names of the two C functions must be given in the primitive declaration, as follows:

        external name : type =

bytecode-C-function-name native-code-C-function-name

For instance, in the case of add_nat, the declaration is:

        external add_nat: nat -> int -> int -> nat -> int -> int -> int -> int
                        = "add_nat_bytecode" "add_nat_native"

Implementing a user primitive is actually two separate tasks: on the one hand, decoding the arguments to extract C values from the given Caml values, and encoding the return value as a Caml value; on the other hand, actually computing the result from the arguments. Except for very simple primitives, it is often preferable to have two distinct C functions to implement these two tasks. The first function actually implements the primitive, taking native C values as arguments and returning a native C value. The second function, often called the “stub code”, is a simple wrapper around the first function that converts its arguments from Caml values to C values, call the first function, and convert the returned C value to Caml value. For instance, here is the stub code for the input primitive:

        value input(channel, buffer, offset, length)
                value channel, buffer, offset, length;
        {
          return Val_long(getblock((struct channel *) channel,
                                   &Byte(buffer, Long_val(offset)),
                                   Long_val(length)));
        }

(Here, Val_long, Long_val and so on are conversion macros for the type value, that will be described later.) The hard work is performed by the function getblock, which is declared as:

        long getblock(channel, p, n)
             struct channel * channel;
             char * p;
             long n;
        {
          ...
        }

To write C code that operates on Objective Caml values, the following include files are provided:

Include file  Provides

caml/mlvalues.h definition of the value type, and conversion macros

caml/alloc.h allocation functions (to create structured Caml objects)

caml/memory.h miscellaneous memory-related functions (for in-place modification of

structures, etc).

caml/callback.h callback from C to Caml (see section 14.6).

These files reside in the caml/ subdirectory of the Objective Caml standard library directory

(usually /usr/local/lib/ocaml).

14.1.3 Linking C code with Caml code

The Objective Caml runtime system comprises three main parts: the bytecode interpreter, the memory manager, and a set of C functions that implement the primitive operations. Some bytecode instructions are provided to call these C functions, designated by their offset in a table of functions (the table of primitives).

In the default mode, the Caml linker produces bytecode for the standard runtime system, with a standard set of primitives. References to primitives that are not in this standard set result in the “unavailable C primitive” error.

In the “custom runtime” mode, the Caml linker scans the object files and determines the set of required primitives. Then, it builds a suitable runtime system, by calling the native code linker with:

• the table of the required primitives

• a library that provides the bytecode interpreter, the memory manager, and the standard primitives

• libraries and object code files (.o files) mentioned on the command line for the Caml linker, that provide implementations for the user’s primitives.

This builds a runtime system with the required primitives. The Caml linker generates bytecode for this custom runtime system. The bytecode is appended to the end of the custom runtime system, so that it will be automatically executed when the output file (custom runtime + bytecode) is launched.

To link in “custom runtime” mode, execute the ocamlc command with:

• the -custom option

• the names of the desired Caml object files (.cmo files)

• the names of the C object files and libraries (.o and .a files) that implement the required primitives. (Under Unix, a library named libname.a residing in one of the standard library directories can also be specified as -cclib -lname.)

14.2 The value type

All Caml objects are represented by the C type value, defined in the include file caml/mlvalues.h, along with macros to manipulate values of that type. An object of type value is either:

• an unboxed integer

• a pointer to a block inside the heap (such as the blocks allocated through one of the alloc_* functions below)

• a pointer to an object outside the heap (e.g., a pointer to a block allocated by malloc, or to a C variable).

14.2.1 Integer values

Integer values encode 31-bit signed integers (63-bit on 64-bit architectures). They are unboxed (unallocated).

14.2.2 Blocks

Blocks in the heap are garbage-collected, and therefore have strict structure constraints. Each block includes a header containing the size of the block (in words), and the tag of the block. The tag governs how the contents of the blocks are structured. A tag lower than No_scan_tag indicates a structured block, containing well-formed values, which is recursively traversed by the garbage collector. A tag greater than or equal to No_scan_tag indicates a raw block, whose contents are not scanned by the garbage collector. For the benefits of ad-hoc polymorphic primitives such as equality and structured input-output, structured and raw blocks are further classified according to their tags as follows:

Tag Contents of the block

0 to No_scan_tag − 1 A structured block (an array of Caml objects). Each field is

                          a value.

Closure_tag A closure representing a functional value. The first word is a pointer to a piece of bytecode, the remaining words are value containing the environment.

String_tag A character string.

Double_tag A double-precision floating-point number.

Double_array_tag An array of double-precision floating-point numbers (for the

native-code compiler only).

Abstract_tag A block representing an abstract datatype.

Final_tag A block representing an abstract datatype with a “finalization” function, to be called when the block is deallocated.

14.2.3 Pointers to outside the heap

Any word-aligned pointer to outside the heap can be safely cast to and from the type value. This includes pointers returned by malloc, and pointers to C variables (of size at least one word) obtained with the & operator.

14.3 Representation of Caml data types

This section describes how Caml data types are encoded in the value type.

14.3.1 Atomic types

Caml type Encoding
int  Unboxed integer values.
char  Unboxed integer values (ASCII code).

                    float  Blocks with tag Double_tag.
                    string  Blocks with tag String_tag.

14.3.2 Tuples and records

Tuples are represented by pointers to blocks, with tag 0.

Records are also represented by zero-tagged blocks. The ordering of labels in the record type declaration determines the layout of the record fields: the value associated to the label declared first is stored in field 0 of the block, the value associated to the label declared next goes in field 1, and so on.

The native-code compiler represents specially records whose fields all have type float. These are represented as arrays of floating-point numbers, with tag Double_array_tag. (See the section below on arrays.)

14.3.3 Arrays

In the bytecode compiler, all arrays are represented like tuples, that is, as pointers to blocks tagged 0.

The native-code compiler has a special, unboxed, more efficient representation for arrays of floating-point numbers (type float array). These arrays are represented by pointers to blocks with tag Double_array_tag. They should be accessed with the Double_field and

Store_double_field macros.

Arrays of floating-point numbers in the bytecode compiler should be accessed with

Double_val(Field(v,  n)) for reading and modify(&Field(v,  n),  copy_double(d)) for

writing.

14.3.4 Concrete types

Constructed terms are represented either by unboxed integers (for constant constructors) or by blocks whose tag encode the constructor (for non-constant constructors). The constant constructors and the non-constant constructors for a given concrete type are numbered separately, starting from 0, in the order in which they appear in the concrete type declaration. Constant constructors are represented by unboxed integers equal to the constructor number. Non-constant constructors declared with a n-tuple as argument are represented by a block of size n, tagged with the constructor number; the n fields contain the components of its tuple argument. Other non-constant constructors are represented by a block of size 1, tagged with the constructor number; the field 0 contains the value of the constructor argument. Example:

Constructed term Representation

             ()               Val_int(0)
             false            Val_int(0)
             true             Val_int(1)
             []               Val_int(0)

h::t Block with size = 2 and tag = 0; first field contains h, second field t

   As a convenience, caml/mlvalues.h defines the macros Val_unit, Val_false and Val_true

to refer to (), false and true.

14.3.5 Objects

Objects are represented as zero-tagged blocks. The first field of the block refers to the object class and associated method suite, in a format that cannot easily be exploited from C. The remaining Chapter 14. Interfacing C with Objective Caml 145

fields of the object contain the values of the instance variables of the object. Instance variables are stored in the order in which they appear in the class definition (taking inherited classes into account).

14.4 Operations on values

14.4.1 Kind tests

• Is_long(v) is true if value v is an immediate integer, false otherwise

• Is_block(v) is true if value v is a pointer to a block, and false if it is an immediate integer.

14.4.2 Operations on integers

• Val_long(l) returns the value encoding the long int l

• Long_val(v) returns the long int encoded in value v

• Val_int(i) returns the value encoding the int i

• Int_val(v) returns the int encoded in value v

14.4.3 Accessing blocks

• Wosize_val(v) returns the size of value v, in words, excluding the header.

• Tag_val(v) returns the tag of value v.

• Field(v, n) returns the value contained in the n^{th} field of the structured block v. Fields are numbered from 0 to Wosize_val(v) − 1.

• Code_val(v) returns the code part of the closure v.

• string_length(v) returns the length (number of characters) of the string v.

• Byte(v, n) returns the n^{th} character of the string v, with type char. Characters are numbered from 0 to string_length(v) − 1.

• Byte_u(v, n) returns the n^{th} character of the string v, with type unsigned char. Characters are numbered from 0 to string_length(v) − 1.

• String_val(v) returns a pointer to the first byte of the string v, with type char *. This pointer is a valid C string: there is a null character after the last character in the string. However, Caml strings can contain embedded null characters, that will confuse the usual C functions over strings.

• Double_val(v) returns the floating-point number contained in value v, with type double.

• (Native-code only.) Double_field(v, n) returns the n^{th} element of the array of floatingpoint numbers v (a block tagged Double_array_tag).

• (Native-code only.) Store_double_field(v, n, d) stores the double precision floating-point number d in the n^{th} element of the array of floating-point numbers v.

The expressions Field(v, n), Byte(v, n) and Byte_u(v, n) are valid l-values. Hence, they can be assigned to, resulting in an in-place modification of value v. Assigning directly to Field(v, n) must be done with care to avoid confusing the garbage collector (see below).

14.4.4 Allocating blocks

From the standpoint of the allocation functions, blocks are divided according to their size as zerosized blocks, small blocks (with size less than or equal to Max_young_wosize), and large blocks (with size greater than to Max_young_wosize). The constant Max_young_wosize is declared in the include file mlvalues.h. It is guaranteed to be at least 64 (words), so that any block with constant size less than or equal to 64 can be assumed to be small. For blocks whose size is computed at run-time, the size must be compared against Max_young_wosize to determine the correct allocation procedure.

• Atom(t) returns an “atom” (zero-sized block) with tag t. Zero-sized blocks are preallocated outside of the heap. It is incorrect to try and allocate a zero-sized block using the functions below. For instance, Atom(0) represents the empty array.

• alloc(n, t) returns a fresh small block of size n ≤ Max_young_wosize words, with tag t. If this block is a structured block (i.e. if t < No_scan_tag), then the fields of the block (initially containing garbage) must be initialized with legal values (using direct assignment to the fields of the block) before the next allocation.

• alloc_tuple(n) returns a fresh small block of size n ≤ Max_young_wosize words, with tag 0. The fields of this block must be filled with legal values before the next allocation or modification.

• alloc_shr(n, t) returns a fresh block of size n, with tag t. The size of the block can be greater than Max_young_wosize. (It can also be smaller, but in this case it is more efficient to call alloc instead of alloc_shr.) If this block is a structured block (i.e. if t < No_scan_tag), then the fields of the block (initially containing garbage) must be initialized with legal values (using the initialize function described below) before the next allocation.

• alloc_string(n) returns a string value of length n characters. The string initially contains garbage.

• copy_string(s) returns a string value containing a copy of the null-terminated C string s (a

char *).

• copy_double(d) returns a floating-point value initialized with the double d.

• alloc_array(f, a) allocates an array of values, calling function f over each element of the input array a to transform it into a value. The array a is an array of pointers terminated by the null pointer. The function f receives each pointer as argument, and returns a value. The zero-tagged block returned by alloc_array(f, a) is filled with the values returned by the successive calls to f.

• copy_string_array(p) allocates an array of strings, copied from the pointer to a string array

     p (a char **).

14.4.5 Raising exceptions

Two functions are provided to raise two standard exceptions:

• failwith(s), where s is a null-terminated C string (with type char *), raises exception Failure with argument s.

• invalid_argument(s), where s is a null-terminated C string (with type char *), raises exception Invalid_argument with argument s.

Raising arbitrary exceptions from C is more delicate: the exception identifier is dynamically allocated by the Caml program, and therefore must be communicated to the C function using the registration facility described below in section 14.6.3. Once the exception identifier is recovered in C, the following functions actually raise the exception:

• raise_constant(id) raises the exception id with no argument;

• raise_with_arg(id, v) raises the exception id with the Caml value v as argument;

• raise_with_string(id, s), where s is a null-terminated C string, raises the exception id with a copy of the C string s as argument.

14.5 Living in harmony with the garbage collector

Unused blocks in the heap are automatically reclaimed by the garbage collector. This requires some cooperation from C code that manipulates heap-allocated blocks.

Rule 1 After a structured block (a block with tag less than No_scan_tag) is allocated, all fields of this block must be filled with well-formed values before the next allocation operation. If the block has been allocated with alloc or alloc_tuple, filling is performed by direct assignment to the fields of the block:

        Field(v, n) = v_{n};

If the block has been allocated with alloc_shr, filling is performed through the initialize function:

        initialize(&Field(v, n), v_{n});

The next allocation can trigger a garbage collection. The garbage collector assumes that all structured blocks contain well-formed values. Newly created blocks contain random data, which generally do not represent well-formed values.

If you really need to allocate before the fields can receive their final value, first initialize with a constant value (e.g. Val_long(0)), then allocate, then modify the fields with the correct value (see rule 3).

Rule 2 Local variables containing values must be registered with the garbage collector (using the Begin_roots and End_roots macros), if they are to survive a call to an allocation function.

Registration is performed with the Begin_roots set of macros. Begin_roots1(v) registers variable v with the garbage collector. Generally, v will be a local variable or a parameter of your function. It must be initialized to a valid value (e.g. Val_unit) before the first allocation. Likewise, Begin_roots2, …, Begin_roots5 let you register up to 5 variables at the same time. Begin_root is the same as Begin_roots1. Begin_roots_block(ptr,size) allows you to register an array of roots. ptr is a pointer to the first element, and size is the number of elements in the array.

Once registered, each of your variables (or array element) has the following properties: if it points to a heap-allocated block, this block (and its contents) will not be reclaimed; moreover, if this block is relocated by the garbage collector, the variable is updated to point to the new location for the block.

Each of the Begin_roots macros open a C block that must be closed with a matching End_roots at the same nesting level. The block must be exited normally (i.e. not with return or goto). However, the roots are automatically un-registered if a Caml exception is raised, so you can exit the block with failwith, invalid_argument, or one of the raise functions.

Note: The Begin_roots macros use a local variable and a structure tag named caml__roots_block. Do not use this identifier in your programs.

Rule 3 Global variables containing values must be registered with the garbage collector using the

register_global_root function.

Registration of a global variable v is achieved by calling register_global_root(&v) just before a valid value is stored in v for the first time.

A registered global variable v can be un-registered by calling remove_global_root(&v).

Rule 4 Direct assignment to a field of a block, as in

        Field(v, n) = w;

is safe only if v is a block newly allocated by alloc or alloc_tuple; that is, if no allocation took place between the allocation of v and the assignment to the field. In all other cases, never assign directly. If the block has just been allocated by alloc_shr, use initialize to assign a value to a field for the first time:

        initialize(&Field(v, n), w);

Otherwise, you are updating a field that previously contained a well-formed value; then, call the modify function:

        modify(&Field(v, n), w);

To illustrate the rules above, here is a C function that builds and returns a list containing the two integers given as parameters:

value alloc_list_int(i1, i2)
        int i1, i2;
{
  value result = Val_unit;
  value r = Val_unit;

  Begin_roots2 (r, result);
    r = alloc(2, 0);                        /* Allocate a cons cell */
    Field(r, 0) = Val_int(i2);              /* car = the integer i2 */
    Field(r, 1) = Val_int(0);               /* cdr = the empty list [] */
    result = alloc(2, 0);                   /* Allocate the other cons cell */
    Field(result, 0) = Val_int(i1);         /* car = the integer i1 */
    Field(result, 1) = r;                   /* cdr = the first cons cell */
  End_roots ();
  return result;
}

Here, the registering of result is not strictly needed, because no allocation takes place after it gets its value, but it’s easier and safer to simply register all the local variables that have type value.

In the example above, the list is built bottom-up. Here is an alternate way, that proceeds top-down. It is less efficient, but illustrates the use of modify.

value alloc_list_int(i1, i2)
        int i1, i2;
{
  value tail;
  value r = Val_unit;
  Begin_root (r);
    r = alloc(2, 0);                        /* Allocate a cons cell */
    Field(r, 0) = Val_int(i1);              /* car = the integer i1 */
    Field(r, 1) = Val_int(0);               /* A dummy value
    tail = alloc(2, 0);                     /* Allocate the other cons cell */
    Field(tail, 0) = Val_int(i2);           /* car = the integer i2 */
    Field(tail, 1) = Val_int(0);            /* cdr = the empty list [] */
    modify(&Field(r, 1), tail);             /* cdr of the result = tail */
  End_roots ();
  return r;
}

It would be incorrect to perform Field(r, 1) = tail directly, because the allocation of tail has taken place since r was allocated. tail is not registered as a root because there is no allocation between the assignment where it takes its value and the modify statement that uses the value.

14.6 Callbacks from C to Caml

So far, we have described how to call C functions from Caml. In this section, we show how C functions can call Caml functions, either as callbacks (Caml calls C which calls Caml), or because the main program is written in C.

14.6.1 Applying Caml closures from C

C functions can apply Caml functional values (closures) to Caml values. The following functions are provided to perform the applications:

• callback(f, a) applies the functional value f to the value a and return the value returned by f.

• callback2(f, a, b) applies the functional value f (which is assumed to be a curried Caml function with two arguments) to a and b.

• callback3(f, a, b, c) applies the functional value f (a curried Caml function with three arguments) to a, b and c.

If the function f does not return, but raises an exception that escapes the scope of the application, then this exception is propagated to the next enclosing Caml code, skipping over the C code. That is, if a Caml function f calls a C function g that calls back a Caml function h that raises a stray exception, then the execution of g is interrupted and the exception is propagated back into f.

14.6.2 Registering Caml closures for use in C functions

The main difficulty with the callback functions described above is obtaining a closure to the Caml function to be called. For this purpose, Objective Caml provides a simple registration mechanism, by which Caml code can register Caml functions under some global name, and then C code can retrieve the corresponding closure by this global name.

On the Caml side, registration is performed by evaluating Callback.register n v. Here, n is the global name (an arbitrary string) and v the Caml value. For instance:

    let f x = print_string "f is applied to "; print_int n; print_newline()
    let _= Callback.register "test function" f

On the C side, a pointer to the value registered under name n is obtained by calling caml_named_value(n). The returned pointer must then be dereferenced to recover the actual Caml value. If no value is registered under the name n, the null pointer is returned. For example, here is a C wrapper that calls the Caml function f above:

    void call_caml_f(int arg)
    {
        callback(*caml_named_value("test function"), Val_int(arg));
    }

The pointer returned by caml_named_value is constant and can safely be cached in a C variable to avoid repeated name lookups. On the other hand, the value pointed to can change during garbage collection and must always be recomputed at the point of use. Here is a more efficient variant of

call_caml_f above that calls caml_named_value only once:
    void call_caml_f(int arg)
    {
        static value * closure_f = NULL;

        if (closure_f == NULL) {
            /* First time around, look up by name */
            closure_f = caml_named_value("test function");
        }
        callback(*closure_f, Val_int(arg));
    }

14.6.3 Registering Caml exceptions for use in C functions

The registration mechanism described above can also be used to communicate exception identifiers from Caml to C. The Caml code registers the exception by evaluating Callback.register_exception n exn, where n is an arbitrary name and exn is an exception value of the exception to register. For example:

    exception Error of string
    let _= Callback.register_exception "test exception" (Error "any string")

The C code can then recover the exception identifier using caml_named_value and pass it as first

argument to the functions raise_constant, raise_with_arg, and raise_with_string (described

in section 14.4.5) to actually raise the exception. For example, here is a C function that raises the Error exception with the given argument:

    void raise_error(char * msg)
    {
        raise_with_string(*caml_named_value("test exception"), msg);
    }

14.6.4 Main program in C

In normal operation, a mixed Caml/C program starts by executing the Caml initialization code, which then can proceed to call C functions. We say that the main program is the Caml code. In some applications, it is desirable that the C code plays the role of the main program, calling Caml functions when needed. This can be achieved as follows:

• The C part of the program must provide a main function, which will override the default main function provided by the Caml runtime system. Execution will start in the user-defined main function just like for a regular C program.

• At some point, the C code must call caml_main(argv) to initialize the Caml code. The argv argument is a C array of strings (type char **) which represents the command-line arguments, as passed as second argument to main. The Caml array Sys.argv will be initialized from this parameter. For the bytecode compiler, argv[0] and argv[1] are also consulted to find the file containing the bytecode.

• The call to caml_main initializes the Caml runtime system, loads the bytecode (in the case of the bytecode compiler), and executes the initialization code of the Caml program. Typically, this initialization code registers callback functions using Callback.register. Once the Caml initialization code is complete, control returns to the C code that called caml_main.

• The C code can then invoke Caml functions using the callback mechanism (see section 14.6.1).

14.6.5 Embedding the Caml code in the C code

The bytecode compiler in custom runtime mode (ocamlc -custom) normally appends the bytecode to the executable file containing the custom runtime. This has two consequences. First, the final linking step must be performed by ocamlc. Second, the Caml runtime library must be able to find the name of the executable file from the command-line arguments. When using caml_main(argv) as in section 14.6.4, this means that argv[0] or argv[1] must contain the executable file name.

An alternative is to embed the bytecode in the C code. The -output-obj option to ocamlc is provided for this purpose. It causes the ocamlc compiler to output a C object file (.o file) containing the bytecode for the Caml part of the program, as well as a caml_startup function. The C object file produced by ocamlc -output-obj can then be linked with C code using the standard C compiler, or stored in a C library.

The caml_startup function must be called from the main C program in order to initialize the Caml runtime and execute the Caml initialization code. Just like caml_main, it takes one argv parameter containing the command-line parameters. Unlike caml_main, this argv parameter is used only to initialize Sys.argv, but not for finding the name of the executable file.

The native-code compiler ocamlopt also supports the -output-obj option, causing it to output a C object file containing the native code for all Caml modules on the command-line, as well as the Caml startup code. Initialization is performed by calling caml_startup as in the case of the bytecode compiler.

Warning: On some ports, special options are required on the final linking phase that links together the object file produced by the -output-obj option and the remainder of the program. Those options are shows in the configuration file config/Makefile generated during compilation of Objective Caml, as the variables BYTECCLINKOPTS (for object files produced by ocamlc -output-obj) and NATIVECCLINKOPTS (for object files produced by ocamlopt -output-obj). Currently, the only ports that require special attention are:

• Digital Unix on the Alpha: object files produced by ocamlc -output-obj must be linked with the -taso option. This is not necessary for object files produced by ocamlopt -output-obj.

• Windows NT: the object file produced by Objective Caml have been compiled with the /MT flag, and therefore all other object files linked with it should also be compiled with /MT.

14.7 A complete example

This section outlines how the functions from the Unix curses library can be made available to Objective Caml programs. First of all, here is the interface curses.mli that declares the curses primitives and data types:

type window                   (* The type "window" remains abstract *)
external initscr: unit -> window = "curses_initscr"
external endwin: unit -> unit = "curses_endwin"
external refresh: unit -> unit = "curses_refresh"
external wrefresh : window -> unit = "curses_wrefresh"
external newwin: int -> int -> int -> int -> window = "curses_newwin"

external mvwin: window -> int -> int -> unit = "curses_mvwin"
external addch: char -> unit = "curses_addch"
external mvwaddch: window -> int -> int -> char -> unit = "curses_mvwaddch"
external addstr: string -> unit = "curses_addstr"
external mvwaddstr: window -> int -> int -> string -> unit = "curses_mvwaddstr"
(* lots more omitted *)

To compile this interface:

        ocamlc -c curses.mli

To implement these functions, we just have to provide the stub code; the core functions are already implemented in the curses library. The stub code file, curses.o, looks like:

#include <curses.h>
#include <mlvalues.h>
value curses_initscr(unit)
        value unit;
{
  return (value) initscr();     /* OK to coerce directly from WINDOW * to value
                                   since that's a block created by malloc() */
}
value curses_wrefresh(win)
        value win;
{
  wrefresh((WINDOW *) win);
  return Val_unit;
}
value curses_newwin(nlines, ncols, x0, y0)
        value nlines, ncols, x0, y0;
{
  return (value) newwin(Int_val(nlines), Int_val(ncols),
                        Int_val(x0), Int_val(y0));
}
value curses_addch(c)
        value c;
{
  addch(Int_val(c));            /* Characters are encoded like integers */
  return Val_unit;
}
value curses_addstr(s)
        value s;

{
  addstr(String_val(s));
  return Val_unit;
}
/* This goes on for pages. */

(Actually, it would be better to create a library for the stub code, with each stub code function in a separate file, so that linking would pick only those functions from the curses library that are actually used.)

The file curses.c can be compiled with:

        cc -c -I/usr/local/lib/ocaml curses.c

or, even simpler,

        ocamlc -c curses.c

(When passed a .c file, the ocamlc command simply calls the C compiler on that file, with the right -I option.)

Now, here is a sample Caml program test.ml that uses the curses module:

open Curses
let main_window = initscr () in
let small_window = newwin 10 5 20 10 in
  mvwaddstr main_window 10 2 "Hello";
  mvwaddstr small_window 4 3 "world";
  refresh();
  for i = 1 to 100000 do () done;
  endwin()

To compile this program, run:

        ocamlc -c test.ml

Finally, to link everything together:

        ocamlc -custom -o test test.cmo curses.o -cclib -lcurses

14.8 Advanced example with callbacks

This section illustrates the callback facilities described in section 14.6. We are going to package some Caml functions in such a way that they can be linked with C code and called from C just like any C functions. The Caml functions are defined in the following mod.ml Caml source:

(* File mod.ml -- some ``useful'' Caml functions *)
let rec fib n = if n < 2 then 1 else fib(n-1) + fib(n-2)

let format_result n = Printf.sprintf "Result is: %d\n" n
(* Export those two functions to C *)
let _= Callback.register "fib" fib
let _= Callback.register "format_result" format_result

Here is the C stub code for calling these functions from C:

/* File modwrap.c -- wrappers around the Caml functions */
#include <stdio.h>
#include <string.h>
#include <caml/mlvalues.h>
#include <caml/callback.h>
int fib(int n)
{
  static value * fib_closure = NULL;
  if (fib_closure == NULL) fib_closure = caml_named_value("fib");
  return Int_val(callback(*fib_closure, Val_int(n)));
}
char * format_result(int n)
{
  static value * format_result_closure = NULL;
  if (format_result_closure == NULL)
    format_result_closure = caml_named_value("format_result");
  return strdup(String_val(callback(*format_result_closure, Val_int(n))));
  /* We copy the C string returned by String_val to the C heap
     so that it remains valid after garbage collection. */
}

We now compile the Caml code to a C object file and put it in a C library along with the stub code in modwrap.c and the Caml runtime system:

        ocamlc -custom -output-obj -o modcaml.o mod.ml
        ocamlc -c modwrap.c
        cp /usr/local/lib/ocaml/libcamlrun.a mod.a
        ar r mod.a modcaml.o modwrap.o

Now, we can use the two fonctions fib and format_result in any C program, just like regular C functions. Just remember to call caml_startup once before.

/* File main.c -- a sample client for the Caml functions */
#include <stdio.h>

int main(int argc, char ** argv)
{
int result;
/* Initialize Caml code */
caml_startup(argv);
/* Do some computation */
result = fib(10);
printf("fib(10) = %s\n", format_result(result));
return 0;
}

To build the whole program, just invoke the C compiler as follows:

      cc -o prog main.c mod.a

Part IV

The Objective Caml library The core library

This chapter describes the functions provided by the Caml Light core library module: module Pervasives. This module is special in two ways:

• It is automatically linked with the user’s object code files by the ocamlc command (chapter 6).

• It is automatically “opened” when a compilation starts, or when the toplevel system is launched. Hence, it is possible to use unqualified identifiers to refer to the functions provided by the Pervasives module, without adding a open Pervasives directive.

Conventions

The declarations from the signature of the Pervasives module are printed one by one in typewriter font, followed by a short comment. All modules and the identifiers they export are indexed at the end of this report.

15.1 Module Pervasives: the initially opened module

This module provides the built-in types (numbers, booleans, strings, exceptions, references, lists, arrays, input-output channels, ...) and the basic operations over these types.

This module is automatically opened at the beginning of each compilation. All components of this module can therefore be referred by their short name, without prefixing them by

Pervasives.

Predefined types

type int

The type of integer numbers.

type char

The type of characters.

type string

The type of character strings.

type float

The type of floating-point numbers.

type bool

The type of booleans (truth values).

type unit = ()

The type of the unit value.

type exn

The type of exception values.

type 'a array

The type of arrays whose elements have type 'a.

type 'a list = [] | :: of 'a * 'a list

The type of lists whose elements have type 'a.

type 'a option = None | Some of 'a

The type of optional values.

type ('a, 'b, 'c) format

The type of format strings. 'a is the type of the parameters of the format, 'c is the result type for the printf-style function, and 'b is the type of the first argument given to %a and %t printing functions (see module Printf).

Exceptions

val raise : exn -> 'a

Raise the given exception value

exception Match_failure of (string * int * int)

Exception raised when none of the cases of a pattern-matching apply. The arguments are the location of the pattern-matching in the source code (file name, position of first character, position of last character).

Chapter 15. The core library 161

exception Assert_failure of (string * int * int)

Exception raised when an assertion fails. The arguments are the location of the pattern-matching in the source code (file name, position of first character, position of last character).

exception Invalid_argument of string

Exception raised by library functions to signal that the given arguments do not make sense.

exception Failure of string

Exception raised by library functions to signal that they are undefined on the given arguments.

exception Not_found

Exception raised by search functions when the desired object could not be found.

exception Out_of_memory

Exception raised by the garbage collector when there is insufficient memory to complete the computation.

exception Stack_overflow

Exception raised by the bytecode interpreter when the evaluation stack reaches its maximal size. This often indicates infinite or excessively deep recursion in the user’s program.

exception Sys_error of string

Exception raised by the input/output functions to report an operating system error.

exception End_of_file

Exception raised by input functions to signal that the end of file has been reached.

exception Division_by_zero

Exception raised by division and remainder operations when their second argument is null.

exception Exit

This exception is not raised by any library function. It is provided for use in your programs.

val invalid_arg: string -> 'a

Raise exception Invalid_argument with the given string.

val failwith: string -> 'a

Raise exception Failure with the given string.

Comparisons

val (=) : 'a -> 'a -> bool

e1 = e2 tests for structural equality of e1 and e2. Mutable structures (e.g. references and arrays) are equal if and only if their current contents are structurally equal, even if the two mutable objects are not the same physical object. Equality between functional values raises Invalid_argument. Equality between cyclic data structures may not terminate.

val (<>) : 'a -> 'a -> bool

Negation of (=).

val (<) : 'a -> 'a -> bool
val (>) : 'a -> 'a -> bool
val (<=) : 'a -> 'a -> bool
val (>=) : 'a -> 'a -> bool

Structural ordering functions. These functions coincide with the usual orderings over integer, string and floating-point numbers, and extend them to a total ordering over all types. The ordering is compatible with (=). As in the case of (=), mutable structures are compared by contents. Comparison between functional values raises Invalid_argument. Comparison between cyclic structures may not terminate.

val compare: 'a -> 'a -> int

compare x y returns 0 if x=y, a negative integer if x<y, and a positive integer if x>y. The same restrictions as for = apply. compare can be used as the comparison function required by the Set and Map modules.

val min: 'a -> 'a -> 'a

Return the smaller of the two arguments.

val max: 'a -> 'a -> 'a

Return the greater of the two arguments.

val (==) : 'a -> 'a -> bool

e1 == e2 tests for physical equality of e1 and e2. On integers and characters, it is the same as structural equality. On mutable structures, e1 == e2 is true if and only if physical modification of e1 also affects e2. On non-mutable structures, the behavior of (==) is implementation-dependent, except that e1 == e2 implies e1 = e2.

val (!=) : 'a -> 'a -> bool

Negation of (==).

Boolean operations

val not : bool -> bool

The boolean negation.

val (&) : bool -> bool -> bool
val (&&) : bool -> bool -> bool

The boolean “and”. Evaluation is sequential, left-to-right: in e1 & e2, e1 is evaluated first, and if it returns false, e2 is not evaluated at all.

val (or) : bool -> bool -> bool
val (||) : bool -> bool -> bool

The boolean “or”. Evaluation is sequential, left-to-right: in e1 or e2, e1 is evaluated first, and if it returns true, e2 is not evaluated at all.

Integer arithmetic

Integers are 31 bits wide (or 63 bits on 64-bit processors). All operations are taken modulo 2^{31} (or 2^{63}). They do not fail on overflow.

val (~-) : int -> int

Unary negation. You can also write -e instead of ~-e.

val succ : int -> int
     succ x is x+1.
val pred : int -> int
     pred x is x-1.
val (+) : int -> int -> int

Integer addition.

val (-) : int -> int -> int

Integer subtraction.

val (*) : int -> int -> int

Integer multiplication.

val (/) : int -> int -> int
val (mod) : int -> int -> int

Integer division and remainder. Raise Division_by_zero if the second argument is 0. If one of the arguments is negative, the result is platform-dependent.

val abs : int -> int

Return the absolute value of the argument.

val max_int: int
val min_int: int

The greatest and smallest representable integers.

Bitwise operations

val (land) : int -> int -> int

Bitwise logical and.

val (lor) : int -> int -> int

Bitwise logical or.

val (lxor) : int -> int -> int

Bitwise logical exclusive or.

val lnot: int -> int

Bitwise logical negation.

val (lsl) : int -> int -> int

n lsl m shifts n to the left by m bits.

val (lsr) : int -> int -> int

n lsr m shifts n to the right by m bits. This is a logical shift: zeroes are inserted regardless of the sign of n.

val (asr) : int -> int -> int

n asr m shifts n to the right by m bits. This is an arithmetic shift: the sign bit of n is replicated.

Floating-point arithmetic

On most platforms, Caml’s floating-point numbers follow the IEEE 754 standard, using double precision (64 bits) numbers. Floating-point operations do not fail on overflow or underflow, but return denormal numbers.

val (~-.) : float -> float

Unary negation. You can also write -.e instead of ~-.e.

val (+.) : float -> float -> float

Floating-point addition

val (-.) : float -> float -> float

Floating-point subtraction

val (*.) : float -> float -> float

Floating-point multiplication Chapter 15. The core library 165

val (/.) : float -> float -> float

Floating-point division.

val (**) : float -> float -> float

Exponentiation

val exp : float -> float
val acos : float -> float
val asin : float -> float
val atan : float -> float
val atan2 : float -> float -> float
val cos : float -> float
val cosh : float -> float
val log : float -> float
val log10 : float -> float
val sin : float -> float
val sinh : float -> float
val sqrt : float -> float
val tan : float -> float
val tanh : float -> float

Usual transcendental functions on floating-point numbers.

val ceil : float -> float
val floor : float -> float

Round the given float to an integer value. floor f returns the greatest integer value less than or equal to f. ceil f returns the least integer value greater than or equal to f.

val abs_float : float -> float

Return the absolute value of the argument.

val mod_float : float -> float -> float

fmod a b returns the remainder of a with respect to b.

val frexp : float -> float * int

frexp f returns the pair of the significant and the exponent of f (when f is zero, the significant x and the exponent n of f are equal to zero; when f is non-zero, they are defined

     by f = x *. 2 ** n).
val ldexp : float -> int -> float
     ldexp x n returns x *. 2 ** n.
val modf : float -> float * float

modf f returns the pair of the fractional and integral part of f.

val float : int -> float

Convert an integer to floating-point.

val truncate : float -> int

Truncate the given floating-point number to an integer. The result is unspecified if it falls outside the range of representable integers.

String operations

More string operations are provided in module String.

val (^) : string -> string -> string

String concatenation.

String conversion functions

val string_of_bool : bool -> string

Return the string representation of a boolean.

val string_of_int : int -> string

Return the string representation of an integer, in decimal.

val int_of_string : string -> int

Convert the given string to an integer. The string is read in decimal (by default) or in hexadecimal, octal or binary if the string begins with 0x, 0o or 0b respectively. Raise Failure "int_of_string" if the given string is not a valid representation of an integer.

val string_of_float : float -> string

Return the string representation of a floating-point number.

val float_of_string : string -> float

Convert the given string to a float. The result is unspecified if the given string is not a valid representation of a float.

Pair operations

val fst : 'a * 'b -> 'a

Return the first component of a pair.

val snd : 'a * 'b -> 'b

Return the second component of a pair.

Chapter 15. The core library 167

List operations

More list operations are provided in module List.

val (@) : 'a list -> 'a list -> 'a list

List concatenation.

Input/output

type in_channel
type out_channel

The types of input channels and output channels.

val stdin : in_channel
val stdout : out_channel
val stderr : out_channel

The standard input, standard output, and standard error output for the process.

Output functions on standard output

val print_char : char -> unit

Print a character on standard output.

val print_string : string -> unit

Print a string on standard output.

val print_int : int -> unit

Print an integer, in decimal, on standard output.

val print_float : float -> unit

Print a floating-point number, in decimal, on standard output.

val print_endline : string -> unit

Print a string, followed by a newline character, on standard output.

val print_newline : unit -> unit

Print a newline character on standard output, and flush standard output. This can be used to simulate line buffering of standard output.

Output functions on standard error

val prerr_char : char -> unit

Print a character on standard error.

val prerr_string : string -> unit

Print a string on standard error.

val prerr_int : int -> unit

Print an integer, in decimal, on standard error.

val prerr_float : float -> unit

Print a floating-point number, in decimal, on standard error.

val prerr_endline : string -> unit

Print a string, followed by a newline character on standard error and flush standard error.

val prerr_newline : unit -> unit

Print a newline character on standard error, and flush standard error.

Input functions on standard input

val read_line : unit -> string

Flush standard output, then read characters from standard input until a newline character is encountered. Return the string of all characters read, without the newline character at the end.

val read_int : unit -> int

Flush standard output, then read one line from standard input and convert it to an integer. Raise Failure "int_of_string" if the line read is not a valid representation of an integer.

val read_float : unit -> float

Flush standard output, then read one line from standard input and convert it to a floating-point number. The result is unspecified if the line read is not a valid representation of a floating-point number.

General output functions

type open_flag =
Open_rdonly | Open_wronly | Open_append
| Open_creat | Open_trunc | Open_excl
| Open_binary | Open_text | Open_nonblock
Opening modes for open_out_gen and open_in_gen.

Open_rdonly: open for reading. Open_wronly: open for writing. Open_append: open for appending. Open_creat: create the file if it does not exist. Open_trunc: empty the file if it already exists. Open_excl: fail if the file already exists. Open_binary: open in binary mode (no conversion). Open_text: open in text mode (may perform conversions). Open_nonblock: open in non-blocking mode.

val open_out : string -> out_channel

Open the named file for writing, and return a new output channel on that file, positionned at the beginning of the file. The file is truncated to zero length if it already exists. It is created if it does not already exists. Raise Sys_error if the file could not be opened.

val open_out_bin : string -> out_channel

Same as open_out, but the file is opened in binary mode, so that no translation takes place during writes. On operating systems that do not distinguish between text mode and binary mode, this function behaves like open_out.

val open_out_gen : open_flag list -> int -> string -> out_channel
open_out_gen mode rights filename opens the file named filename for writing, as

above. The extra argument mode specify the opening mode. The extra argument rights specifies the file permissions, in case the file must be created. open_out and open_out_bin are special cases of this function.

val flush : out_channel -> unit

Flush the buffer associated with the given output channel, performing all pending writes on that channel. Interactive programs must be careful about flushing standard output and standard error at the right time.

val output_char : out_channel -> char -> unit

Write the character on the given output channel.

val output_string : out_channel -> string -> unit

Write the string on the given output channel.

val output : out_channel -> string -> int -> int -> unit

output chan buff ofs len writes len characters from string buff, starting at offset ofs,

to the output channel chan. Raise Invalid_argument "output" if ofs and len do not

designate a valid substring of buff.

val output_byte : out_channel -> int -> unit

Write one 8-bit integer (as the single character with that code) on the given output channel. The given integer is taken modulo 256.

val output_binary_int : out_channel -> int -> unit

Write one integer in binary format on the given output channel. The only reliable way to read it back is through the input_binary_int function. The format is compatible across all machines for a given version of Objective Caml.

val output_value : out_channel -> 'a -> unit

Write the representation of a structured value of any type to a channel. Circularities and sharing inside the value are detected and preserved. The object can be read back, by the function input_value. See the description of module Marshal for more information. output_value is equivalent to Marshal.to_channel with an empty list of flags.

val seek_out : out_channel -> int -> unit

seek_out chan pos sets the current writing position to pos for channel chan. This works only for regular files. On files of other kinds (such as terminals, pipes and sockets), the behavior is unspecified.

val pos_out : out_channel -> int

Return the current writing position for the given channel.

val out_channel_length : out_channel -> int

Return the total length (number of characters) of the given channel. This works only for regular files. On files of other kinds, the result is meaningless.

val close_out : out_channel -> unit

Close the given channel, flushing all buffered write operations. The behavior is unspecified if any of the functions above is called on a closed channel.

General input functions

val open_in : string -> in_channel

Open the named file for reading, and return a new input channel on that file, positionned at the beginning of the file. Raise Sys_error if the file could not be opened.

val open_in_bin : string -> in_channel

Same as open_in, but the file is opened in binary mode, so that no translation takes place during reads. On operating systems that do not distinguish between text mode and binary mode, this function behaves like open_in.

val open_in_gen : open_flag list -> int -> string -> in_channel
open_in_gen mode rights filename opens the file named filename for reading, as above.

The extra arguments mode and rights specify the opening mode and file permissions. open_in and open_in_bin are special cases of this function.

val input_char : in_channel -> char

Read one character from the given input channel. Raise End_of_file if there are no more characters to read.

val input_line : in_channel -> string

Read characters from the given input channel, until a newline character is encountered. Return the string of all characters read, without the newline character at the end. Raise End_of_file if the end of the file is reached at the beginning of line.

val input : in_channel -> string -> int -> int -> int

input chan buff ofs len attempts to read len characters from channel chan, storing them in string buff, starting at character number ofs. It returns the actual number of characters read, between 0 and len (inclusive). A return value of 0 means that the end of file was reached. A return value between 0 and len exclusive means that no more characters were available at that time; input must be called again to read the remaining characters, if desired. Exception Invalid_argument "input" is raised if ofs and len do not designate a valid substring of buff.

val really_input : in_channel -> string -> int -> int -> unit

really_input chan buff ofs len reads len characters from channel chan, storing them in string buff, starting at character number ofs. Raise End_of_file if the end of file is reached before len characters have been read. Raise Invalid_argument "really_input" if ofs and len do not designate a valid substring of buff.

val input_byte : in_channel -> int

Same as input_char, but return the 8-bit integer representing the character. Raise End_of_file if an end of file was reached.

val input_binary_int : in_channel -> int

Read an integer encoded in binary format from the given input channel. See output_binary_int. Raise End_of_file if an end of file was reached while reading the integer.

val input_value : in_channel -> 'a

Read the representation of a structured value, as produced by output_value, and return the corresponding value. This function is identical to Marshal.from_channel; see the description of module Marshal for more information, in particular concerning the lack of type safety.

val seek_in : in_channel -> int -> unit

seek_in chan pos sets the current reading position to pos for channel chan. This works only for regular files. On files of other kinds, the behavior is unspecified.

val pos_in : in_channel -> int

Return the current reading position for the given channel.

val in_channel_length : in_channel -> int

Return the total length (number of characters) of the given channel. This works only for regular files. On files of other kinds, the result is meaningless.

val close_in : in_channel -> unit

Close the given channel. Anything can happen if any of the functions above is called on a closed channel.

References

type 'a ref = { mutable contents: 'a }

The type of references (mutable indirection cells) containing a value of type 'a.

val ref : 'a -> 'a ref

Return a fresh reference containing the given value.

val (!) : 'a ref -> 'a

!r returns the current contents of reference r. Could be defined as fun r -> r.contents.

val (:=) : 'a ref -> 'a -> unit

r := a stores the value of a in reference r. Could be defined as

fun r v -> r.contents <- v.
val incr : int ref -> unit

Increment the integer contained in the given reference. Could be defined as

fun r -> r := succ !r.
val decr : int ref -> unit

Decrement the integer contained in the given reference. Could be defined as

fun r -> r := pred !r.

Program termination

val exit : int -> 'a

Flush all pending writes on stdout and stderr, and terminate the process, returning the given status code to the operating system (usually 0 to indicate no errors, and a small positive integer to indicate failure.) An implicit exit 0 is performed each time a program terminates normally (but not if it terminates because of an uncaught exception).

val at_exit: (unit -> unit) -> unit

Register the given function to be called at program termination time. The functions registered with at_exit will be called in some unspecified order when the program executes exit. They will not be called if the program terminates because of an uncaught exception.

174

The standard library

This chapter describes the functions provided by the Caml Light standard library. The modules from the standard library are automatically linked with the user’s object code files by the ocamlc command. Hence, these modules can be used in standalone programs without having to add any .cmo file on the command line for the linking phase. Similarly, in interactive use, these globals can be used in toplevel phrases without having to load any .cmo file in memory.

Unlike the Pervasive module from the core library, the modules from the standard library are not automatically “opened” when a compilation starts, or when the toplevel system is launched. Hence it is necessary to use qualified identifiers to refer to the functions provided by these modules, or to add open directives.

Conventions

For easy reference, the modules are listed below in alphabetical order of module names. For each module, the declarations from its signature are printed one by one in typewriter font, followed by a short comment. All modules and the identifiers they export are indexed at the end of this report.

Overview

Here is a short listing, by theme, of the standard library modules.

Data structures:

Char  p. 180 character operations
String p. 210 string operations
Array  p. 178 array operations
List  p. 197 list operations
Sort  p. 208 sorting and merging lists
Hashtbl p. 193 hash tables and hash functions
Random p. 206 pseudo-random number generator
Set  p. 206 sets over ordered types
Map  p. 199 association tables over ordered types
Oo  p. 203 useful functions on objects
Stack  p. 208 last-in first-out stacks
Queue  p. 205 first-in first-out queues
Lazy  p. 195 delayed evaluation
Weak  p. 215 references that don’t prevent objects from being garbage-collected

Input/output:

Format p. 183 pretty printing
Marshal p. 201 marshaling of data structures
Printf p. 204 formatting printing functions
Digest p. 181 MD5 message digest

Parsing:

Genlex p. 192 a generic lexer over streams
Lexing p. 195 the run-time library for lexers generated by camllex
Parsing p. 203 the run-time library for parsers generated by camlyacc
Stream p. 209 basic functions over streams

System interface:

Arg  p. 176 parsing of command line arguments
Callback p. 180 registering Caml functions to be called from C
Filename p. 182 operations on file names
Gc  p. 189 memory management control and statistics
Printexc p. 203 a catch-all exception handler
Sys  p. 212 system interface

16.1 Module Arg: parsing of command line arguments

This module provides a general mechanism for extracting options and arguments from the command line to the program.

Syntax of command lines: A keyword is a character string starting with a -. An option is a keyword alone or followed by an argument. There are six types of keywords: Unit, Set,

Clear, String, Int, and Float. Unit, Set and Clear keywords take no argument. String,

Int, and Float keywords take the following word on the command line as an argument. Arguments not preceded by a keyword are called anonymous arguments.

Examples (cmd is assumed to be the command name): cmd -flag (a unit option) cmd -int 1 (an int option with argument 1) cmd -string foobar (a string option with argument "foobar") cmd -float 12.34 (a float option with argument 12.34) cmd a b c (three anonymous arguments: "a", "b", and "c")

type spec =
| Unit of (unit -> unit)     (* Call the function with unit argument *)
| Set of bool ref            (* Set the reference to true *)
| Clear of bool ref          (* Set the reference to false *)
| String of (string -> unit) (* Call the function with a string argument *)
| Int of (int -> unit)       (* Call the function with an int argument *)
| Float of (float -> unit)   (* Call the function with a float argument *)

The concrete type describing the behavior associated with a keyword.

val parse : (string * spec * string) list -> (string -> unit) -> string -> unit
parse speclist anonfun usage_msg parses the command line. speclist is a list of

triples (key, spec, doc). key is the option keyword, it must start with a '-' character. spec gives the option type and the function to call when this option is found on the command line. doc is a one-line description of this option. anonfun is called on anonymous arguments. The functions in spec and anonfun are called in the same order as their arguments appear on the command line.

If an error occurs, parse exits the program, after printing an error message as follows: The reason for the error: unknown option, invalid or missing argument, etc.

usage_msg

The list of options, each followed by the corresponding doc string.

For the user to be able to specify anonymous arguments starting with a -, include for

example ("--", String anonfun, doc) in speclist.

By default, parse recognizes a unit option -help, which will display usage_msg and the list of options, and exit the program. You can override this behaviour by specifying your own

-help option in speclist.
exception Bad of string

Functions in spec or anonfun can raise Bad with an error message to reject invalid arguments.

val usage: (string * spec * string) list -> string -> unit
usage speclist usage_msg speclist and usage_msg are the same as for parse. usage

prints the same error message that parse prints in case of error.

val current: int ref;;

Position (in Sys.argv) of the argument being processed. You can change this value, e.g. to force parse to skip some arguments.

16.2 Module Array: array operations

val length : 'a array -> int

Return the length (number of elements) of the given array.

val get: 'a array -> int -> 'a

Array.get a n returns the element number n of array a. The first element has number 0. The last element has number Array.length a - 1. Raise

Invalid_argument "Array.get" if n is outside the range 0 to (Array.length a - 1).

You can also write a.(n) instead of Array.get a n.

val set: 'a array -> int -> 'a -> unit

Array.set a n x modifies array a in place, replacing element number n with x. Raise

Invalid_argument "Array.set" if n is outside the range 0 to Array.length a - 1. You

can also write a.(n) <- x instead of Array.set a n x.

val make: int -> 'a -> 'a array
val create: int -> 'a -> 'a array

Array.make n x returns a fresh array of length n, initialized with x. All the elements of this new array are initially physically equal to x (in the sense of the == predicate). Consequently, if x is mutable, it is shared among all elements of the array, and modifying x through one of the array entries will modify all other entries at the same time.

val init: int -> (int -> 'a) -> 'a array

Array.init n f returns a fresh array of length n, with element number i equal to f i.

val make_matrix: int -> int -> 'a -> 'a array array
val create_matrix: int -> int -> 'a -> 'a array array

Array.make_matrix dimx dimy e returns a two-dimensional array (an array of arrays) with first dimension dimx and second dimension dimy. All the elements of this new matrix are initially physically equal to e. The element (x,y) of a matrix m is accessed with the

notation m.(x).(y).

Chapter 16. The standard library 179

val append: 'a array -> 'a array -> 'a array

Array.append v1 v2 returns a fresh array containing the concatenation of arrays v1 and

v2.
val concat: 'a array list -> 'a array

Same as Array.append, but catenates a list of arrays.

val sub: 'a array -> int -> int -> 'a array

Array.sub a start len returns a fresh array of length len, containing the elements

number start to start + len - 1 of array a. Raise Invalid_argument "Array.sub" if

start and len do not designate a valid subarray of a; that is, if start < 0, or len < 0, or

start + len > Array.length a.
val copy: 'a array -> 'a array

Array.copy a returns a copy of a, that is, a fresh array containing the same elements as a.

val fill: 'a array -> int -> int -> 'a -> unit

Array.fill a ofs len x modifies the array a in place, storing x in elements number ofs

to ofs + len - 1. Raise Invalid_argument "Array.fill" if ofs and len do not

designate a valid subarray of a.

val blit: 'a array -> int -> 'a array -> int -> int -> unit

Array.blit v1 o1 v2 o2 len copies len elements from array v1, starting at element number o1, to array v2, starting at element number o2. It works correctly even if v1 and v2 are the same array, and the source and destination chunks overlap. Raise Invalid_argument "Array.blit" if o1 and len do not designate a valid subarray of v1, or if o2 and len do not designate a valid subarray of v2.

val to_list: 'a array -> 'a list

Array.to_list a returns the list of all the elements of a.

val of_list: 'a list -> 'a array

Array.of_list l returns a fresh array containing the elements of l.

val iter: ('a -> unit) -> 'a array -> unit

Array.iter f a applies function f in turn to all the elements of a. It is equivalent to

f a.(0); f a.(1); ...; f a.(Array.length a - 1); ().
val map: ('a -> 'b) -> 'a array -> 'b array

Array.map f a applies function f to all the elements of a, and builds an array with the

results returned by f: [| f a.(0); f a.(1); ...; f a.(Array.length a - 1) |].

val iteri: (int -> 'a -> unit) -> 'a array -> unit
val mapi: (int -> 'a -> 'b) -> 'a array -> 'b array

Same as Array.iter and Array.map respectively, but the function is applied to the index of the element as first argument, and the element itself as second argument.

val fold_left: ('a -> 'b -> 'a) -> 'a -> 'b array -> 'a
Array.fold_left f x a computes f (... (f (f x a.(0)) a.(1)) ...) a.(n-1),

where n is the length of the array a.

val fold_right: ('b -> 'a -> 'a) -> 'b array -> 'a -> 'a
Array.fold_right f a x computes f a.(0) (f a.(1) ( ... (f a.(n-1) x) ...)),

where n is the length of the array a.

16.3 Module Callback: registering Caml values with the C

runtime

This module allows Caml values to be registered with the C runtime under a symbolic name, so that C code can later call back registered Caml functions, or raise registered Caml exceptions.

val register: string -> 'a -> unit

Callback.register n v registers the value v under the name n. C code can later retrieve a

handle to v by calling caml_named_value(n).
val register_exception: string -> exn -> unit

Callback.register_exception n exn registers the exception contained in the exception value exn under the name n. C code can later retrieve a handle to the exception by calling caml_named_value(n). The exception value thus obtained is suitable for passign as first

argument to raise_constant or raise_with_arg.

16.4 Module Char: character operations

val code: char -> int

Return the ASCII code of the argument.

val chr: int -> char

Return the character with the given ASCII code. Raise Invalid_argument "Char.chr" if the argument is outside the range 0–255.

Chapter 16. The standard library 181

val escaped : char -> string

Return a string representing the given character, with special characters escaped following the lexical conventions of Objective Caml.

val lowercase: char -> char
val uppercase: char -> char

Convert the given character to its equivalent lowercase or uppercase character, respectively.

16.5 Module Digest: MD5 message digest

This module provides functions to compute 128-bit “digests” of arbitrary-length strings or files. The digests are cryptographic quality: it is very hard, given a digest, to forge a string having that digest. The algorithm used is MD5.

type t = string

The type of digests: 16-character strings.

val string: string -> t

Return the digest of the given string.

val substring: string -> int -> int -> t

Digest.substring s ofs len returns the digest of the substring of s starting at character number ofs and containing len characters.

val channel: in_channel -> int -> t

Digest.channel ic len reads len characters from channel ic and returns their digest.

val file: string -> t

Return the digest of the file whose name is given.

val output: out_channel -> t -> unit

Write a digest on the given output channel.

val input: in_channel -> t

Read a digest from the given input channel.

16.6 Module Filename: operations on file names

val current_dir_name : string

The conventional name for the current directory (e.g. . in Unix).

val concat : string -> string -> string

concat dir file returns a file name that designates file file in directory dir.

val is_relative : string -> bool

Return true if the file name is relative to the current directory, false if it is absolute (i.e. in Unix, starts with /.

val is_implicit : string -> bool

Return true if the file name is relative and does not start with an explicit reference to the current directory (./ or ../ in Unix), false if it starts with an explicit reference to the root directory or the current directory.

val check_suffix : string -> string -> bool

check_suffix name suff returns true if the filename name ends with the suffix suff.

val chop_suffix : string -> string -> string

chop_suffix name suff removes the suffix suff from the filename name. The behavior is undefined if name does not end with the suffix suff.

val chop_extension : string -> string

Return the given file name without its extension. The extension is the shortest suffix starting with a period, .xyz for instance. Raise Invalid_argument if the given name does not contain a period.

val basename : string -> string
val dirname : string -> string

Split a file name into directory name / base file name.

concat (dirname name) (basename name) returns a file name which is equivalent to name.

Moreover, after setting the current directory to dirname name (with Sys.chdir), references to basename name (which is a relative file name) designate the same file as name before the call to chdir.

val temp_file: string -> string -> string

temp_file prefix suffix returns the name of a non-existent temporary file in the temporary directory. The base name of the temporary file is formed by concatenating prefix, then a suitably chosen integer number, then suffix. Under Unix, the temporary directory is /tmp by default; if set, the value of the environment variable TMPDIR is used instead. Under Windows, the name of the temporary directory is the value of the environment variable TEMP, or C:\temp by default.

16.7 Module Format: pretty printing

This module implements a pretty-printing facility to format text within “pretty-printing boxes”. The pretty-printer breaks lines at specified break hints, and indents lines according to the box structure.

Rule of thumb for casual users: use simple boxes (as obtained by open_box 0); use simple break hints (as obtained by print_cut () that outputs a simple break hint, or by print_space () that ouputs a space indicating a break hint); once a box is opened, display its material with basic printing functions (e. g. print_int

and print_string);

when the material for a box has been printed, call close_box () to close the box; at the end of your routine, evaluate print_newline () to close all remaining boxes and flush the pretty-printer.

You may alternatively consider this module as providing an extension to the printf facility: you can simply add pretty-printing annotations to your regular printf formats, as explained below in the documentation of the function fprintf.

The behaviour of pretty-printing commands is unspecified if there is no opened pretty-printing box. Each box opened via one of the open_functions below must be closed using close_box for proper formatting. Otherwise, some of the material printed in the boxes may not be output, or may be formatted incorrectly.

In case of interactive use, the system closes all opened boxes and flushes all pending text (as with the print_newline function) after each phrase. Each phrase is therefore executed in the initial state of the pretty-printer.

Boxes

val open_box : int -> unit;;

open_box d opens a new pretty-printing box with offset d. This box is the general purpose pretty-printing box. Material in this box is displayed “horizontal or vertical”: break hints inside the box may lead to a new line, if there is no more room on the line to print the remainder of the box, or if a new line may lead to a new indentation (demonstrating the indentation of the box). When a new line is printed in the box, d is added to the current indentation.

val close_box : unit -> unit;;

Close the most recently opened pretty-printing box.

Formatting functions

val print_string : string -> unit;;

print_string str prints str in the current box.

val print_as : int -> string -> unit;;

print_as len str prints str in the current box. The pretty-printer formats str as if it were of length len.

val print_int : int -> unit;;

Print an integer in the current box.

val print_float : float -> unit;;

Print a floating point number in the current box.

val print_char : char -> unit;;

Print a character in the current box.

val print_bool : bool -> unit;;

Print an boolean in the current box.

Break hints

val print_space : unit -> unit;;

print_space () is used to separate items (typically to print a space between two words). It indicates that the line may be split at this point. It either prints one space or splits the line. It is equivalent to print_break 1 0.

val print_cut : unit -> unit;;

print_cut () is used to mark a good break position. It indicates that the line may be split at this point. It either prints nothing or splits the line. This allows line splitting at the current point, without printing spaces or adding indentation. It is equivalent to

print_break 0 0.
val print_break : int -> int -> unit;;

Insert a break hint in a pretty-printing box. print_break nspaces offset indicates that the line may be split (a newline character is printed) at this point, if the contents of the current box does not fit on one line. If the line is split at that point, offset is added to the current indentation. If the line is not split, nspaces spaces are printed.

val print_flush : unit -> unit;;

Flush the pretty printer: all opened boxes are closed, and all pending text is displayed.

Chapter 16. The standard library 185

val print_newline : unit -> unit;;

Equivalent to print_flush followed by a new line.

val force_newline : unit -> unit;;

Force a newline in the current box. Not the normal way of pretty-printing, you should prefer break hints.

val print_if_newline : unit -> unit;;

Execute the next formatting command if the preceding line has just been split. Otherwise, ignore the next formatting command.

Margin

val set_margin : int -> unit;;

set_margin d sets the value of the right margin to d (in characters): this value is used to detect line overflows that leads to split lines. Nothing happens if d is smaller than 2 or bigger than 999999999.

val get_margin : unit -> int;;

Return the position of the right margin.

Maximum indentation limit

val set_max_indent : int -> unit;;

set_max_indent d sets the value of the maximum indentation limit to d (in characters): once this limit is reached, boxes are rejected to the left, if they do not fit on the current line. Nothing happens if d is smaller than 2 or bigger than 999999999.

val get_max_indent : unit -> int;;

Return the value of the maximum indentation limit (in characters).

Formatting depth: maximum number of boxes allowed before ellipsis

val set_max_boxes : int -> unit;;

set_max_boxes max sets the maximum number of boxes simultaneously opened. Material inside boxes nested deeper is printed as an ellipsis (more precisely as the text returned by get_ellipsis_text ()). Nothing happens if max is not greater than 1.

val get_max_boxes : unit -> int;;

Return the maximum number of boxes allowed before ellipsis.

val over_max_boxes : unit -> bool;;

Test the maximum number of boxes allowed have already been opened.

Advanced formatting

val open_hbox : unit -> unit;;

open_hbox () opens a new pretty-printing box. This box is “horizontal”: the line is not split in this box (new lines may still occur inside boxes nested deeper).

val open_vbox : int -> unit;;

open_vbox d opens a new pretty-printing box with offset d. This box is “vertical”: every break hint inside this box leads to a new line. When a new line is printed in the box, d is added to the current indentation.

val open_hvbox : int -> unit;;

open_hvbox d opens a new pretty-printing box with offset d. This box is “horizontal-vertical”: it behaves as an “horizontal” box if it fits on a single line, otherwise it behaves as a “vertical” box. When a new line is printed in the box, d is added to the current indentation.

val open_hovbox : int -> unit;;

open_hovbox d opens a new pretty-printing box with offset d. This box is “horizontal or vertical”: break hints inside this box may lead to a new line, if there is no more room on the line to print the remainder of the box. When a new line is printed in the box, d is added to the current indentation.

Tabulations

val open_tbox : unit -> unit;;

Open a tabulation box.

val close_tbox : unit -> unit;;

Close the most recently opened tabulation box.

val print_tbreak : int -> int -> unit;;

Break hint in a tabulation box. print_tbreak spaces offset moves the insertion point to the next tabulation (spaces being added to this position). Nothing occurs if insertion point is already on a tabulation mark. If there is no next tabulation on the line, then a newline is printed and the insertion point moves to the first tabulation of the box. If a new line is printed, offset is added to the current indentation.

val set_tab : unit -> unit;;

Set a tabulation mark at the current insertion point.

val print_tab : unit -> unit;;
print_tab () is equivalent to print_tbreak (0,0).

Ellipsis

val set_ellipsis_text : string -> unit;;

Set the text of the ellipsis printed when too many boxes are opened (a single dot, ., by default).

val get_ellipsis_text : unit -> string;;

Return the text of the ellipsis.

Redirecting formatter output

val set_formatter_out_channel : out_channel -> unit;;

Redirect the pretty-printer output to the given channel.

val set_formatter_output_functions :
 (string -> int -> int -> unit) -> (unit -> unit) -> unit;;
set_formatter_output_functions out flush redirects the pretty-printer output to the

functions out and flush. The out function performs the pretty-printer output. It is called with a string s, a start position p, and a number of characters n; it is supposed to output characters p to p+n-1 of s. The flush function is called whenever the pretty-printer is

flushed using print_flush or print_newline.
val get_formatter_output_functions :
   unit -> (string -> int -> int -> unit) * (unit -> unit);;

Return the current output functions of the pretty-printer.

Multiple formatted output

type formatter;;

Abstract data type corresponding to a pretty-printer and all its machinery. Defining new pretty-printers permits the output of material in parallel on several channels. Parameters of the pretty-printer are local to the pretty-printer: margin, maximum indentation limit, maximum number of boxes simultaneously opened, ellipsis, and so on, are specific to each pretty-printer and may be fixed independently. A new formatter is obtained by calling the

make_formatter function.
val std_formatter : formatter;;

The standard formatter used by the formatting functions above. It is defined using make_formatter with output function output stdout and flushing function

fun () -> flush stdout.

val err_formatter : formatter;;

A formatter to use with formatting functions below for output to standard error. It is defined using make_formatter with output function output stderr and flushing function

     fun () -> flush stderr.
val make_formatter :
        (string -> int -> int -> unit) -> (unit -> unit) -> formatter;;

make_formatter out flush returns a new formatter that writes according to the output function out, and flushing function flush. Hence, a formatter to out channel oc is returned

     by make_formatter (output oc) (fun () -> flush oc).
val pp_open_hbox : formatter -> unit -> unit;;
val pp_open_vbox : formatter -> int -> unit;;
val pp_open_hvbox : formatter -> int -> unit;;
val pp_open_hovbox : formatter -> int -> unit;;
val pp_open_box : formatter -> int -> unit;;
val pp_close_box : formatter -> unit -> unit;;
val pp_print_string : formatter -> string -> unit;;
val pp_print_as : formatter -> int -> string -> unit;;
val pp_print_int : formatter -> int -> unit;;
val pp_print_float : formatter -> float -> unit;;
val pp_print_char : formatter -> char -> unit;;
val pp_print_bool : formatter -> bool -> unit;;
val pp_print_break : formatter -> int -> int -> unit;;
val pp_print_cut : formatter -> unit -> unit;;
val pp_print_space : formatter -> unit -> unit;;
val pp_force_newline : formatter -> unit -> unit;;
val pp_print_flush : formatter -> unit -> unit;;
val pp_print_newline : formatter -> unit -> unit;;
val pp_print_if_newline : formatter -> unit -> unit;;
val pp_open_tbox : formatter -> unit -> unit;;
val pp_close_tbox : formatter -> unit -> unit;;
val pp_print_tbreak : formatter -> int -> int -> unit;;
val pp_set_tab : formatter -> unit -> unit;;
val pp_print_tab : formatter -> unit -> unit;;
val pp_set_margin : formatter -> int -> unit;;
val pp_get_margin : formatter -> unit -> int;;
val pp_set_max_indent : formatter -> int -> unit;;
val pp_get_max_indent : formatter -> unit -> int;;
val pp_set_max_boxes : formatter -> int -> unit;;
val pp_get_max_boxes : formatter -> unit -> int;;
val pp_over_max_boxes : formatter -> unit -> bool;;
val pp_set_ellipsis_text : formatter -> string -> unit;;
val pp_get_ellipsis_text : formatter -> unit -> string;;
val pp_set_formatter_out_channel : formatter -> out_channel -> unit;;

val pp_set_formatter_output_functions : formatter ->
   (string -> int -> int -> unit) -> (unit -> unit) -> unit;;
val pp_get_formatter_output_functions :
   formatter -> unit -> (string -> int -> int -> unit) * (unit -> unit);;

The basic functions to use with formatters. These functions are the basic ones: usual functions operating on the standard formatter are defined via partial evaluation of these

primitives. For instance, print_string is equal to pp_print_string std_formatter.
val fprintf : formatter -> ('a, formatter, unit) format -> 'a;;

fprintf ff format arg1 ... argN formats the arguments arg1 to argN according to the format string format, and outputs the resulting string on the formatter ff. The format is a character string which contains three types of objects: plain characters and conversion specifications as specified in the printf module, and pretty-printing indications. The pretty-printing indication characters are introduced by a @ character, and their meanings are: [: open a pretty-printing box. The type and offset of the box may be optionally specified with the following syntax: the < character, followed by an optional box type indication, then an optional integer offset, and the closing > character. Box type is one of h, v, hv, or hov, which stand respectively for an horizontal, vertical, “horizontal-vertical” and “horizontal or vertical” box. ]: close the most recently opened pretty-printing box. ,: output a good break as with print_cut (). : output a space, as with print_space (). \n: force a newline, as with force_newline (). ;: output a good break as with print_break. The nspaces and offset parameters of the break may be optionally specified with the following syntax: the < character, followed by an integer nspaces value, then an integer offset, and a closing > character. .: flush the pretty printer as with print_newline (). @: a plain @ character.

val printf : ('a, formatter, unit) format -> 'a;;
Same as fprintf, but output on std_formatter.
val eprintf: ('a, formatter, unit) format -> 'a;;
Same as fprintf, but output on err_formatter.

16.8 Module Gc: memory management control and statistics

type stat = {
minor_words : int;
promoted_words : int;

major_words : int;
minor_collections : int;
major_collections : int;
heap_words : int;
heap_chunks : int;
live_words : int;
live_blocks : int;
free_words : int;
free_blocks : int;
largest_free : int;
fragments : int;
compactions : int
}

The memory management counters are returned in a stat record. All the numbers are computed since the start of the program. The fields of this record are: minor_words Number of words allocated in the minor heap. promoted_words Number of words allocated in the minor heap that survived a minor collection and were moved to the major heap. major_words Number of words allocated in the major heap, including the promoted words. minor_collections Number of minor collections. major_collections Number of major collection cycles, not counting the current cycle. heap_words Total number of words in the major heap. heap_chunks Number of times the major heap size was increased. live_words Number of words of live data in the major heap, including the header words. live_blocks Number of live objects in the major heap. free_words Number of words in the free list. free_blocks Number of objects in the free list. largest_free Size (in words) of the largest object in the free list. fragments Number of wasted words due to fragmentation. These are 1-words free blocks placed between two live objects. They cannot be inserted in the free list, thus they are not available for allocation. compactions Number of heap compactions.

The total amount of memory allocated by the program is (in words)

minor_words + major_words - promoted_words. Multiply by the word size (4 on a 32-bit

machine, 8 on a 64-bit machine) to get the number of bytes.

type control = {
mutable minor_heap_size : int;
mutable major_heap_increment : int;
mutable space_overhead : int;
mutable verbose : bool;
mutable max_overhead : int;

mutable stack_limit : int
}

The GC parameters are given as a control record. The fields are: minor_heap_size The size (in words) of the minor heap. Changing this parameter will trigger a minor collection. Default: 32k. major_heap_increment The minimum number of words to add to the major heap when increasing it. Default: 62k. space_overhead The major GC speed is computed from this parameter. This is the memory that will be ”wasted” because the GC does not immediatly collect unreachable objects. It is expressed as a percentage of the memory used for live data. The GC will work more (use more CPU time and collect objects more eagerly) if space_overhead is smaller. The computation of the GC speed assumes that the amount of live data is constant. Default: 42. max_overhead Heap compaction is triggered when the estimated amount of free memory is more than max_overhead percent of the amount of live data. If max_overhead is set to 0, heap compaction is triggered at the end of each major GC cycle (this setting is intended for testing purposes only). If max_overhead >= 1000000, compaction is never triggered. Default: 1000000. verbose This flag controls the GC messages on standard error output. Default: false. stack_limit The maximum size of the stack (in words). This is only relevant to the byte-code runtime, as the native code runtime uses the operating system’s stack. Default: 256k.

val stat : unit -> stat

Return the current values of the memory management counters in a stat record.

val print_stat : out_channel -> unit

Print the current values of the memory management counters (in human-readable form) into the channel argument.

val get : unit -> control

Return the current values of the GC parameters in a control record.

val set : control -> unit

set r changes the GC parameters according to the control record r. The normal usage is:

       let r = Gc.get () in    (* Get the current parameters. *)
         r.verbose <- true;    (* Change some of them. *)
         Gc.set r              (* Set the new values. *)
val minor : unit -> unit

Trigger a minor collection.

val major : unit -> unit

Finish the current major collection cycle.

val full_major : unit -> unit

Finish the current major collection cycle and perform a complete new cycle. This will collect all currently unreachable objects.

val compact : unit -> unit = "gc_compaction";;

Perform a full major collection and compact the heap. Note that heap compaction is a lengthy operation.

16.9 Module Genlex: a generic lexical analyzer

This module implements a simple “standard” lexical analyzer, presented as a function from character streams to token streams. It implements roughly the lexical conventions of Caml, but is parameterized by the set of keywords of your language.

type token =
Kwd of string
| Ident of string
| Int of int
| Float of float
| String of string
| Char of char

The type of tokens. The lexical classes are: Int and Float for integer and floating-point numbers; String for string literals, enclosed in double quotes; Char for character literals, enclosed in single quotes; Ident for identifiers (either sequences of letters, digits, underscores and quotes, or sequences of “operator characters” such as +, *, etc); and Kwd for keywords (either identifiers or single “special characters” such as (, }, etc).

val make_lexer: string list -> (char Stream.t -> token Stream.t)

Construct the lexer function. The first argument is the list of keywords. An identifier s is returned as Kwd s if s belongs to this list, and as Ident s otherwise. A special character s is returned as Kwd s if s belongs to this list, and cause a lexical error (exception Parse_error) otherwise. Blanks and newlines are skipped. Comments delimited by (* and *) are skipped as well, and can be nested.

Example: a lexer suitable for a desk calculator is obtained by

           let lexer = make_lexer ["+";"-";"*";"/";"let";"="; "("; ")"]

The associated parser would be a function from token stream to, for instance, int, and would have rules such as:

           let parse_expr = parser
                  [< 'Int n >] -> n
                | [< 'Kwd "("; n = parse_expr; 'Kwd ")" >] -> n
                | [< n1 = parse_expr; n2 = parse_remainder n1 >] -> n2
           and parse_remainder n1 = parser
                  [< 'Kwd "+"; n2 = parse_expr >] -> n1+n2
                | ...

16.10 Module Hashtbl: hash tables and hash functions

Hash tables are hashed association tables, with in-place modification.

Generic interface

type ('a, 'b) t

The type of hash tables from type 'a to type 'b.

val create : int -> ('a,'b) t

Hashtbl.create n creates a new, empty hash table, with initial size n. The table grows as needed, so n is just an initial guess. Better results are said to be achieved when n is a prime

number. Raise Invalid_argument "hashtbl__new" if n is less than 1.
val clear : ('a, 'b) t -> unit

Empty a hash table.

val add : ('a, 'b) t -> 'a -> 'b -> unit

Hashtbl.add tbl x y adds a binding of x to y in table tbl. Previous bindings for x are not removed, but simply hidden. That is, after performing remove tbl x, the previous binding for x, if any, is restored. (This is the semantics of association lists.)

val find : ('a, 'b) t -> 'a -> 'b

Hashtbl.find tbl x returns the current binding of x in tbl, or raises Not_found if no such binding exists.

val find_all : ('a, 'b) t -> 'a -> 'b list

Hashtbl.find_all tbl x returns the list of all data associated with x in tbl. The current binding is returned first, then the previous bindings, in reverse order of introduction in the table.

val remove : ('a, 'b) t -> 'a -> unit

Hashtbl.remove tbl x removes the current binding of x in tbl, restoring the previous binding if it exists. It does nothing if x is not bound in tbl.

val iter : ('a -> 'b -> unit) -> ('a, 'b) t -> unit

Hashtbl.iter f tbl applies f to all bindings in table tbl. f receives the key as first argument, and the associated value as second argument. The order in which the bindings are passed to f is unspecified. Each binding is presented exactly once to f.

Functorial interface

module type HashedType =
sig
type t
val equal: t -> t -> bool
val hash: t -> int
end

The input signature of the functor Hashtbl.Make. t is the type of keys. equal is the equality predicate used to compare keys. hash is a hashing function on keys, returning a non-negative integer. It must be such that if two keys are equal according to equal, then they must have identical hash values as computed by hash. Examples: suitable (equal, hash) pairs for arbitrary key types include ((=), Hashtbl.hash) for comparing objects by structure, ((==), Hashtbl.hash) for comparing objects by addresses (e.g. for mutable or cyclic keys).

module type S =
sig
type key
type 'a t
val create: int -> 'a t
val clear: 'a t -> unit
val add: 'a t -> key -> 'a -> unit
val remove: 'a t -> key -> unit
val find: 'a t -> key -> 'a
val find_all: 'a t -> key -> 'a list
val iter: (key -> 'a -> unit) -> 'a t -> unit
end
module Make(H: HashedType): (S with type key = H.t)

The functor Hashtbl.Make returns a structure containing a type key of keys and a type 'a t of hash tables associating data of type 'a to keys of type key. The operations perform similarly to those of the generic interface, but use the hashing and equality functions specified in the functor argument H instead of generic equality and hashing.

The polymorphic hash primitive

val hash : 'a -> int

Hashtbl.hash x associates a positive integer to any value of any type. It is guaranteed that if x = y, then hash x = hash y. Moreover, hash always terminates, even on cyclic structures.

val hash_param : int -> int -> 'a -> int

Hashtbl.hash_param n m x computes a hash value for x, with the same properties as for hash. The two extra parameters n and m give more precise control over hashing. Hashing performs a depth-first, right-to-left traversal of the structure x, stopping after n meaningful nodes were encountered, or m nodes, meaningful or not, were encountered. Meaningful nodes are: integers; floating-point numbers; strings; characters; booleans; and constant constructors. Larger values of m and n means that more nodes are taken into account to compute the final hash value, and therefore collisions are less likely to happen. However, hashing takes longer. The parameters m and n govern the tradeoff between accuracy and speed.

16.11 Module Lazy: deferred computations.

type 'a status =
| Delayed of (unit -> 'a)
| Value of 'a
| Exception of exn
;;
type 'a t = 'a status ref;;

A value of type 'a Lazy.t is a deferred computation (called a suspension) that computes a result of type 'a. The expression lazy (expr) returns a suspension that computes expr.

val force: 'a t -> 'a;;

Lazy.force x computes the suspension x and returns its result. If the suspension was already computed, Lazy.force x returns the same value again. If it raised an exception, the same exception is raised again.

16.12 Module Lexing: the run-time library for lexers generated

    by camllex

Lexer buffers

type lexbuf =
{ refill_buff : lexbuf -> unit;
mutable lex_buffer : string;

mutable lex_buffer_len : int;
mutable lex_abs_pos : int;
mutable lex_start_pos : int;
mutable lex_curr_pos : int;
mutable lex_last_pos : int;
mutable lex_last_action : int;
mutable lex_eof_reached : bool }

The type of lexer buffers. A lexer buffer is the argument passed to the scanning functions defined by the generated scanners. The lexer buffer holds the current state of the scanner, plus a function to refill the buffer from the input.

val from_channel : in_channel -> lexbuf

Create a lexer buffer on the given input channel. Lexing.from_channel inchan returns a lexer buffer which reads from the input channel inchan, at the current reading position.

val from_string : string -> lexbuf

Create a lexer buffer which reads from the given string. Reading starts from the first character in the string. An end-of-input condition is generated when the end of the string is reached.

val from_function : (string -> int -> int) -> lexbuf

Create a lexer buffer with the given function as its reading method. When the scanner needs more characters, it will call the given function, giving it a character string s and a character count n. The function should put n characters or less in s, starting at character number 0, and return the number of characters provided. A return value of 0 means end of input.

Functions for lexer semantic actions

The following functions can be called from the semantic actions of lexer definitions (the ML code enclosed in braces that computes the value returned by lexing functions). They give access to the character string matched by the regular expression associated with the semantic action. These functions must be applied to the argument lexbuf, which, in the code generated by camllex, is bound to the lexer buffer passed to the parsing function.

val lexeme : lexbuf -> string

Lexing.lexeme lexbuf returns the string matched by the regular expression.

val lexeme_char : lexbuf -> int -> char

Lexing.lexeme_char lexbuf i returns character number i in the matched string.

val lexeme_start : lexbuf -> int

Lexing.lexeme_start lexbuf returns the position in the input stream of the first character of the matched string. The first character of the stream has position 0.

Chapter 16. The standard library 197

val lexeme_end : lexbuf -> int

Lexing.lexeme_end lexbuf returns the position in the input stream of the character following the last character of the matched string. The first character of the stream has position 0.

16.13 Module List: list operations

val length : 'a list -> int

Return the length (number of elements) of the given list.

val hd : 'a list -> 'a

Return the first element of the given list. Raise Failure "hd" if the list is empty.

val tl : 'a list -> 'a list

Return the given list without its first element. Raise Failure "tl" if the list is empty.

val nth : 'a list -> int -> 'a

Return the n-th element of the given list. The first element (head of the list) is at position 0. Raise Failure "nth" if the list is too short.

val rev : 'a list -> 'a list

List reversal.

val concat  : 'a list list -> 'a list
val flatten : 'a list list -> 'a list

Catenate (flatten) a list of lists.

Iterators

val iter : ('a -> unit) -> 'a list -> unit

List.iter f [a1; ...; an] applies function f in turn to a1; ...; an. It is equivalent to

begin f a1; f a2; ...; f an; () end.
val map : ('a -> 'b) -> 'a list -> 'b list

List.map f [a1; ...; an] applies function f to a1, ..., an, and builds the list [f a1; ...; f an] with the results returned by f.

val fold_left : ('a -> 'b -> 'a) -> 'a -> 'b list -> 'a
List.fold_left f a [b1; ...; bn] is f (... (f (f a b1) b2) ...) bn.
val fold_right : ('a -> 'b -> 'b) -> 'a list -> 'b -> 'b
List.fold_right f [a1; ...; an] b is f a1 (f a2 (... (f an b) ...)).

Iterators on two lists

val iter2 : ('a -> 'b -> unit) -> 'a list -> 'b list -> unit
List.iter2 f [a1; ...; an] [b1; ...; bn] calls in turn f a1 b1; ...; f an bn.

Raise Invalid_argument if the two lists have different lengths.

val map2 : ('a -> 'b -> 'c) -> 'a list -> 'b list -> 'c list
List.map2 f [a1; ...; an] [b1; ...; bn] is [f a1 b1; ...; f an bn]. Raise

Invalid_argument if the two lists have different lengths.

val fold_left2 : ('a -> 'b -> 'c -> 'a) -> 'a -> 'b list -> 'c list -> 'a
List.fold_left2 f a [b1; ...; bn] [c1; ...; cn] is
f (... (f (f a b1 c1) b2 c2) ...) bn cn. Raise Invalid_argument if the two lists

have different lengths.

val fold_right2 : ('a -> 'b -> 'c -> 'c) -> 'a list -> 'b list -> 'c -> 'c
List.fold_right2 f [a1; ...; an] [b1; ...; bn] c is
f a1 b1 (f a2 b2 (... (f an bn c) ...)). Raise Invalid_argument if the two lists

have different lengths.

List scanning

val for_all : ('a -> bool) -> 'a list -> bool

for_all p [a1; ...; an] checks if all elements of the list satisfy the predicate p. That is,

it returns (p a1) & (p a2) & ... & (p an).
val exists : ('a -> bool) -> 'a list -> bool

exists p [a1; ...; an] checks if at least one element of the list satisfies the predicate p.

That is, it returns (p a1) or (p a2) or ... or (p an).
val for_all2 : ('a -> 'b -> bool) -> 'a list -> 'b list -> bool
val exists2 : ('a -> 'b -> bool) -> 'a list -> 'b list -> bool

Same as for_all and exists, but for a two-argument predicate. Raise Invalid_argument if the two lists have different lengths.

val mem : 'a -> 'a list -> bool

mem a l is true if and only if a is equal to an element of l.

val memq : 'a -> 'a list -> bool

Same as mem, but uses physical equality instead of structural equality to compare list elements.

Association lists

val assoc : 'a -> ('a * 'b) list -> 'b

assoc a l returns the value associated with key a in the list of pairs l. That is, assoc a [ ...; (a,b); ...] = b if (a,b) is the leftmost binding of a in list l. Raise Not_found if there is no value associated with a in the list l.

val mem_assoc : 'a -> ('a * 'b) list -> bool

Same as assoc, but simply return true if a binding exists, and false if no bindings exist for the given key.

val assq : 'a -> ('a * 'b) list -> 'b

Same as assoc, but uses physical equality instead of structural equality to compare keys.

Lists of pairs

val split : ('a * 'b) list -> 'a list * 'b list

Transform a list of pairs into a pair of lists: split [(a1,b1); ...; (an,bn)] is

([a1; ...; an], [b1; ...; bn])
val combine : 'a list -> 'b list -> ('a * 'b) list

Transform a pair of lists into a list of pairs: combine ([a1; ...; an], [b1; ...; bn]) is [(a1,b1); ...; (an,bn)]. Raise Invalid_argument if the two lists have different lengths.

16.14 Module Map: association tables over ordered types

This module implements applicative association tables, also known as finite maps or dictionaries, given a total ordering function over the keys. All operations over maps are purely applicative (no side-effects). The implementation uses balanced binary trees, and therefore searching and insertion take time logarithmic in the size of the map.

module type OrderedType =
sig
type t
val compare: t -> t -> int
end

The input signature of the functor Map.Make. t is the type of the map keys. compare is a total ordering function over the keys. This is a two-argument function f such that f e1 e2 is zero if the keys e1 and e2 are equal, f e1 e2 is strictly negative if e1 is smaller than e2, and f e1 e2 is strictly positive if e1 is greater than e2. Example: a suitable ordering function is the generic structural comparison function compare.

module type S =
sig
type key

The type of the map keys.

type 'a t

The type of maps from type key to type 'a.

val empty: 'a t

The empty map.

val add: key -> 'a -> 'a t -> 'a t

add x y m returns a map containing the same bindings as m, plus a binding of x to y. If x was already bound in m, its previous binding disappears.

val find: key -> 'a t -> 'a

find x m returns the current binding of x in m, or raises Not_found if no such binding exists.

val remove: key -> 'a t -> 'a t

remove x m returns a map containing the same bindings as m, except for x which is unbound in the returned map.

val iter: (key -> 'a -> unit) -> 'a t -> unit

iter f m applies f to all bindings in map m. f receives the key as first argument, and the associated value as second argument. The order in which the bindings are passed to f is unspecified. Only current bindings are presented to f: bindings hidden by more recent bindings are not passed to f.

val map: ('a -> 'b) -> 'a t -> 'b t

map f m returns a map with same domain as m, where the associated value a of all bindings of m has been replaced by the result of the application of f to a. The order in which the associated values are passed to f is unspecified.

val fold: (key -> 'a -> 'b -> 'b) -> 'a t -> 'b -> 'b
fold f m a computes (f kN dN ... (f k1 d1 a)...), where k1 ... kN are the keys of

all bindings in m, and d1 ... dN are the associated data. The order in which the bindings are presented to f is unspecified.

end
module Make(Ord: OrderedType): (S with type key = Ord.t)

Functor building an implementation of the map structure given a totally ordered type.

16.15 Module Marshal: marshaling of data structures

This module provides functions to encode arbitrary data structures as sequences of bytes, which can then be written on a file or sent over a pipe or network connection. The bytes can then be read back later, possibly in another process, and decoded back into a data structure. The format for the byte sequences is compatible across all machines for a given version of Objective Caml.

Warning: marshaling is currently not type-safe. The type of marshaled data is not transmitted along the value of the data, making it impossible to check that the data read back possesses the type expected by the context. In particular, the result type of the Marshal.from_* functions is given as 'a, but this is misleading: the returned Caml value does not possess type 'a for all 'a; it has one, unique type which cannot be determined at compile-type. The programmer should explicitly give the expected type of the returned value, using the following syntax: (Marshal.from_channel chan : type). Anything can happen at run-time if the object in the file does not belong to the given type.

The representation of marshaled values is not human-readable, and uses bytes that are not printable characters. Therefore, input and output channels used in conjunction with

Marshal.to_channel and Marshal.from_channel must be opened in binary mode, using

e.g. open_out_bin or open_in_bin; channels opened in text mode will cause unmarshaling errors on platforms where text channels behave differently than binary channels, e.g. Windows.

type extern_flags =
No_sharing                          (* Don't preserve sharing *)
| Closures                            (* Send function closures *)

The flags to the Marshal.to_* functions below.

val to_channel: out_channel -> 'a -> extern_flags list -> unit

Marshal.to_channel chan v flags writes the representation of v on channel chan. The flags argument is a possibly empty list of flags that governs the marshaling behavior with respect to sharing and functional values.

If flags does not contain Marshal.No_sharing, circularities and sharing inside the value v are detected and preserved in the sequence of bytes produced. In particular, this guarantees that marshaling always terminates. Sharing between values marshaled by successive calls to

Marshal.to_channel is not detected, though. If flags contains Marshal.No_sharing,

sharing is ignored. This results in faster marshaling if v contains no shared substructures, but may cause slower marshaling and larger byte representations if v actually contains sharing, or even non-termination if v contains cycles.

If flags does not contain Marshal.Closures, marshaling fails when it encounters a functional value inside v: only “pure” data structures, containing neither functions nor objects, can safely be transmitted between different programs. If flags contains Marshal.Closures, functional values will be marshaled as a position in the code of the program. In this case, the output of marshaling can only be read back in processes that run exactly the same program, with exactly the same compiled code. (This is checked at un-marshaling time, using an MD5 digest of the code transmitted along with the code position.)

val to_string: 'a -> extern_flags list -> string

Marshal.to_string v flags returns a string containing the representation of v as a sequence of bytes. The flags argument has the same meaning as for Marshal.to_channel.

val to_buffer: string -> int -> int -> 'a -> extern_flags list -> int
Marshal.to_buffer buff ofs len v flags marshals the value v, storing its byte

representation in the string buff, starting at character number ofs, and writing at most len characters. It returns the number of characters actually written to the string. If the byte representation of v does not fit in len characters, the exception Failure is raised.

val from_channel: in_channel -> 'a

Marshal.from_channel chan reads from channel chan the byte representation of a structured value, as produced by one of the Marshal.to_* functions, and reconstructs and returns the corresponding value.

val from_string: string -> int -> 'a

Marshal.from_string buff ofs unmarshals a structured value like Marshal.from_channel does, except that the byte representation is not read from a channel, but taken from the string buff, starting at position ofs.

val header_size : int
val data_size : string -> int -> int
val total_size : string -> int -> int

The bytes representing a marshaled value are composed of a fixed-size header and a variable-sized data part, whose size can be determined from the header. Marshal.header_size is the size, in characters, of the header. Marshal.data_size buff ofs is the size, in characters, of the data part, assuming a valid header is stored in buff starting at position ofs. Finally, Marshal.total_size buff ofs is the total size, in characters, of the marshaled value. Both Marshal.data_size and Marshal.total_size raise Failure if buff, ofs does not contain a valid header.

To read the byte representation of a marshaled value into a string buffer, the program needs to read first Marshal.header_size characters into the buffer, then determine the length of the remainder of the representation using Marshal.data_size, make sure the buffer is large enough to hold the variable size, then read it, and finally call Marshal.from_string to unmarshal the value.

16.16 Module Oo: object-oriented extension

val copy : (< .. > as 'a) -> 'a

Oo.copy o returns a copy of object o, that is a fresh object with the same methods and instance variables as o

16.17 Module Parsing: the run-time library for parsers generated by camlyacc

val symbol_start : unit -> int
val symbol_end : unit -> int

symbol_start and symbol_end are to be called in the action part of a grammar rule only. They return the position of the string that matches the left-hand side of the rule: symbol_start() returns the position of the first character; symbol_end() returns the position of the last character, plus one. The first character in a file is at position 0.

val rhs_start: int -> int
val rhs_end: int -> int

Same as symbol_start and symbol_end, but return the position of the string matching the nth item on the right-hand side of the rule, where n is the integer parameter to lhs_start and lhs_end. n is 1 for the leftmost item.

val clear_parser : unit -> unit

Empty the parser stack. Call it just after a parsing function has returned, to remove all pointers from the parser stack to structures that were built by semantic actions during parsing. This is optional, but lowers the memory requirements of the programs.

exception Parse_error

Raised when a parser encounters a syntax error. Can also be raised from the action part of a grammar rule, to initiate error recovery.

16.18 Module Printexc: a catch-all exception handler

val catch: ('a -> 'b) -> 'a -> 'b

Printexc.catch fn x applies fn to x and returns the result. If the evaluation of fn x raises any exception, the name of the exception is printed on standard error output, and the programs aborts with exit code 2. Typical use is Printexc.catch main (), where main, with type unit->unit, is the entry point of a standalone program. This catches and reports any exception that escapes the program.

val print: ('a -> 'b) -> 'a -> 'b

Same as catch, but re-raise the stray exception after printing it, instead of aborting the program.

val to_string : exn -> string

Printexc.to_string e returns a string representation of e.

16.19 Module Printf: formatting printing functions

val fprintf: out_channel -> ('a, out_channel, unit) format -> 'a
fprintf outchan format arg1 ... argN formats the arguments arg1 to argN according

to the format string format, and outputs the resulting string on the channel outchan.

The format is a character string which contains two types of objects: plain characters, which are simply copied to the output channel, and conversion specifications, each of which causes conversion and printing of one argument.

Conversion specifications consist in the % character, followed by optional flags and field widths, followed by one conversion character. The conversion characters and their meanings are: d or i: convert an integer argument to signed decimal u: convert an integer argument to unsigned decimal x: convert an integer argument to unsigned hexadecimal, using lowercase letters. X: convert an integer argument to unsigned hexadecimal, using uppercase letters. s: insert a string argument c: insert a character argument f: convert a floating-point argument to decimal notation, in the style dddd.ddd e or E: convert a floating-point argument to decimal notation, in the style d.ddd e+-dd (mantissa and exponent) g or G: convert a floating-point argument to decimal notation, in style f or e, E (whichever is more compact) b: convert a boolean argument to the string true or false a: user-defined printer. Takes two arguments and apply the first one to outchan (the current output channel) and to the second argument. The first argument must therefore have type out_channel -> 'b -> unit and the second 'b. The output produced by the function is therefore inserted in the output of fprintf at the current point. t: same as %a, but takes only one argument (with type out_channel -> unit) and apply it

to outchan.

Refer to the C library printf function for the meaning of flags and field width specifiers.

If too few arguments are provided, printing stops just before converting the first missing argument.

Chapter 16. The standard library 205

val printf: ('a, out_channel, unit) format -> 'a

Same as fprintf, but output on stdout.

val eprintf: ('a, out_channel, unit) format -> 'a

Same as fprintf, but output on stderr.

val sprintf: ('a, unit, string) format -> 'a

Same as printf, but return the result of formatting in a string.

16.20 Module Queue: first-in first-out queues

This module implements queues (FIFOs), with in-place modification.

type 'a t

The type of queues containing elements of type 'a.

exception Empty

Raised when take is applied to an empty queue.

val create: unit -> 'a t

Return a new queue, initially empty.

val add: 'a -> 'a t -> unit

add x q adds the element x at the end of the queue q.

val take: 'a t -> 'a

take q removes and returns the first element in queue q, or raises Empty if the queue is empty.

val peek: 'a t -> 'a

peek q returns the first element in queue q, without removing it from the queue, or raises Empty if the queue is empty.

val clear : 'a t -> unit

Discard all elements from a queue.

val length: 'a t -> int

Return the number of elements in a queue.

val iter: ('a -> unit) -> 'a t -> unit

iter f q applies f in turn to all elements of q, from the least recently entered to the most recently entered. The queue itself is unchanged.

16.21 Module Random: pseudo-random number generator

val init : int -> unit

Initialize the generator, using the argument as a seed. The same seed will always yield the same sequence of numbers.

val full_init : int array -> unit

Same as init but takes more data as seed.

val bits : unit -> int

Return 30 random bits in a nonnegative integer.

val int : int -> int

Random.int bound returns a random integer between 0 (inclusive) and bound (exclusive). bound must be more than 0 and less than 2^{30}.

val float : float -> float

Random.float bound returns a random floating-point number between 0 (inclusive) and bound (exclusive). If bound is negative, the result is negative. If bound is 0, the result is 0.

16.22 Module Set: sets over ordered types

This module implements the set data structure, given a total ordering function over the set elements. All operations over sets are purely applicative (no side-effects). The implementation uses balanced binary trees, and is therefore reasonably efficient: insertion and membership take time logarithmic in the size of the set, for instance.

module type OrderedType =
sig
type t
val compare: t -> t -> int
end

The input signature of the functor Set.Make. t is the type of the set elements. compare is a total ordering function over the set elements. This is a two-argument function f such that f e1 e2 is zero if the elements e1 and e2 are equal, f e1 e2 is strictly negative if e1 is smaller than e2, and f e1 e2 is strictly positive if e1 is greater than e2. Example: a suitable ordering function is the generic structural comparison function compare.

module type S =
sig
type elt

The type of the set elements.

Chapter 16. The standard library 207

type t

The type of sets.

val empty: t

The empty set.

val is_empty: t -> bool

Test whether a set is empty or not.

val mem: elt -> t -> bool

mem x s tests whether x belongs to the set s.

val add: elt -> t -> t

add x s returns a set containing all elements of s, plus x. If x was already in s, s is returned unchanged.

val remove: elt -> t -> t

remove x s returns a set containing all elements of s, except x. If x was not in s, s is returned unchanged.

val union: t -> t -> t
val inter: t -> t -> t
val diff: t -> t -> t

Union, intersection and set difference.

val compare: t -> t -> int

Total ordering between sets. Can be used as the ordering function for doing sets of sets.

val equal: t -> t -> bool

equal s1 s2 tests whether the sets s1 and s2 are equal, that is, contain the same elements.

val subset: t -> t -> bool

subset s1 s2 tests whether the set s1 is a subset of the set s2.

val iter: (elt -> unit) -> t -> unit

iter f s applies f in turn to all elements of s. The order in which the elements of s are presented to f is unspecified.

val fold: (elt -> 'a -> 'a) -> t -> 'a -> 'a
fold f s a computes (f xN ... (f x2 (f x1 a))...), where x1 ... xN are the

elements of s. The order in which elements of s are presented to f is unspecified.

val cardinal: t -> int

Return the number of elements of a set.

val elements: t -> elt list

Return the list of all elements of the given set. The elements appear in the list in some unspecified order.

val choose: t -> elt

Return one element of the given set, or raise Not_found if the set is empty. Which element is chosen is unspecified, but equal elements will be chosen for equal sets.

end
module Make(Ord: OrderedType): (S with type elt = Ord.t)

Functor building an implementation of the set structure given a totally ordered type.

16.23 Module Sort: sorting and merging lists

val list : ('a -> 'a -> bool) -> 'a list -> 'a list

Sort a list in increasing order according to an ordering predicate. The predicate should return true if its first argument is less than or equal to its second argument.

val merge : ('a -> 'a -> bool) -> 'a list -> 'a list -> 'a list

Merge two lists according to the given predicate. Assuming the two argument lists are sorted according to the predicate, merge returns a sorted list containing the elements from the two lists. The behavior is undefined if the two argument lists were not sorted.

16.24 Module Stack: last-in first-out stacks

This module implements stacks (LIFOs), with in-place modification.

type 'a t

The type of stacks containing elements of type 'a.

exception Empty

Raised when pop is applied to an empty stack.

val create: unit -> 'a t

Return a new stack, initially empty.

val push: 'a -> 'a t -> unit

push x s adds the element x at the top of stack s.

val pop: 'a t -> 'a

pop s removes and returns the topmost element in stack s, or raises Empty if the stack is empty.

val clear : 'a t -> unit

Discard all elements from a stack.

val length: 'a t -> int

Return the number of elements in a stack.

val iter: ('a -> unit) -> 'a t -> unit

iter f s applies f in turn to all elements of s, from the element at the top of the stack to the element at the bottom of the stack. The stack itself is unchanged.

16.25 Module Stream: streams and parsers

type 'a t

The type of streams holding values of type 'a.

exception Failure;;

Raised by parsers when none of the first components of the stream patterns is accepted.

exception Error of string;;

Raised by parsers when the first component of a stream pattern is accepted, but one of the following components is rejected.

Stream builders Warning: these functions create streams with fast access; it is illegal to mix them with streams built with [< >]; would raise Failure when accessing such mixed streams.

val from : (int -> 'a option) -> 'a t;;

Stream.from f returns a stream built from the function f. To create a new stream element, the function f is called with the current stream count. The user function f must return either Some <value> for a value or None to specify the end of the stream.

val of_list : 'a list -> 'a t;;

Return the stream holding the elements of the list in the same order.

val of_string : string -> char t;;

Return the stream of the characters of the string parameter.

val of_channel : in_channel -> char t;;

Return the stream of the characters read from the input channel.

Stream iterator

val iter : ('a -> unit) -> 'a t -> unit;;

Stream.iter f s scans the whole stream s, applying function f in turn to each stream element encountered.

Predefined parsers

val next : 'a t -> 'a;;

Return the first element of the stream and remove it from the stream. Raise Stream.Failure if the stream is empty.

val empty : 'a t -> unit;;

Return () if the stream is empty, else raise Stream.Failure.

Useful functions

val peek : 'a t -> 'a option;;

Return Some of ”the first element” of the stream, or None if the stream is empty.

val junk : 'a t -> unit;;

Remove the first element of the stream, possibly unfreezing it before.

val count : 'a t -> int = "%field0";;

Return the current count of the stream elements, i.e. the number of the stream elements discarded.

val npeek : int -> 'a t -> 'a list;;

npeek n returns the list of the n first elements of the stream, or all its remaining elements if less than n elements are available.

16.26 Module String: string operations

val length : string -> int

Return the length (number of characters) of the given string.

val get : string -> int -> char

String.get s n returns character number n in string s. The first character is character number 0. The last character is character number String.length s - 1. Raise

Invalid_argument if n is ouside the range 0 to (String.length s - 1). You can also
write s.[n] instead of String.get s n.

Chapter 16. The standard library 211

val set : string -> int -> char -> unit

String.set s n c modifies string s in place, replacing the character number n by c. Raise

Invalid_argument if n is ouside the range 0 to (String.length s - 1). You can also
write s.[n] <- c instead of String.set s n c.
val create : int -> string

String.create n returns a fresh string of length n. The string initially contains arbitrary characters.

val make : int -> char -> string

String.make n c returns a fresh string of length n, filled with the character c.

val copy : string -> string

Return a copy of the given string.

val sub : string -> int -> int -> string

String.sub s start len returns a fresh string of length len, containing the characters

number start to start + len - 1 of string s. Raise Invalid_argument if start and len

do not designate a valid substring of s; that is, if start < 0, or len < 0, or

start + len > String.length s.
val fill : string -> int -> int -> char -> unit

String.fill s start len c modifies string s in place, replacing the characters number

start to start + len - 1 by c. Raise Invalid_argument if start and len do not

designate a valid substring of s.

val blit : string -> int -> string -> int -> int -> unit
String.blit src srcoff dst dstoff len copies len characters from string src, starting

at character number srcoff, to string dst, starting at character number dstoff. It works correctly even if src and dst are the same string, and the source and destination chunks overlap. Raise Invalid_argument if srcoff and len do not designate a valid substring of src, or if dstoff and len do not designate a valid substring of dst.

val concat : string -> string list -> string

String.concat sep sl catenates the list of strings sl, inserting the separator string sep between each.

val escaped: string -> string

Return a copy of the argument, with special characters represented by escape sequences, following the lexical conventions of Objective Caml.

val index: string -> char -> int

index s c returns the position of the leftmost occurrence of character c in string s. Raise Not_found if c does not occur in s.

val rindex: string -> char -> int

rindex s c returns the position of the rightmost occurrence of character c in string s. Raise Not_found if c does not occur in s.

val index_from: string -> int -> char -> int
val rindex_from: string -> int -> char -> int

Same as index and rindex, but start searching at the character position given as second argument. index s c is equivalent to index_from s 0 c, and rindex s c to

rindex_from s (String.length s - 1) c.
val uppercase: string -> string

Return a copy of the argument, with all lowercase letters translated to uppercase, including accented letters of the ISO Latin-1 (8859-1) character set.

val lowercase: string -> string

Return a copy of the argument, with all uppercase letters translated to lowercase, including accented letters of the ISO Latin-1 (8859-1) character set.

val capitalize: string -> string

Return a copy of the argument, with the first letter set to uppercase.

val uncapitalize: string -> string

Return a copy of the argument, with the first letter set to lowercase.

16.27 Module Sys: system interface

val argv: string array

The command line arguments given to the process. The first element is the command name used to invoke the program. The following elements are the arguments given to the program.

val file_exists: string -> bool

Test if a file with the given name exists.

val remove: string -> unit

Remove the given file name from the file system.

Chapter 16. The standard library 213

val rename : string -> string -> unit

Rename a file. The first argument is the old name and the second is the new name.

val getenv: string -> string

Return the value associated to a variable in the process environment. Raise Not_found if the variable is unbound.

val command: string -> int

Execute the given shell command and return its exit code.

val chdir: string -> unit

Change the current working directory of the process.

val getcwd: unit -> string

Return the current working directory of the process.

val interactive: bool ref

This reference is initially set to false in standalone programs and to true if the code is being executed under the interactive toplevel csltop.

val os_type: string

Operating system currently executing the Caml program. One of "Unix", "Win32", or

"MacOS".
val word_size: int

Size of one word on the machine currently executing the Caml program, in bits: 32 or 64.

val max_string_length: int

Maximum length of a string.

val max_array_length: int

Maximum length of an array.

Signal handling

type signal_behavior =
    Signal_default
  | Signal_ignore
  | Signal_handle of (int -> unit)

What to do when receiving a signal:

Signal_default: take the default behavior

Signal_ignore: ignore the signal

Signal_handle f: call function f, giving it the signal number as argument.

val signal: int -> signal_behavior -> unit

Set the behavior of the system on receipt of a given signal. The first argument is the signal number.

val sigabrt: int   (* Abnormal termination *)
val sigalrm: int   (* Timeout *)
val sigfpe: int    (* Arithmetic exception *)
val sighup: int    (* Hangup on controlling terminal *)
val sigill: int    (* Invalid hardware instruction *)
val sigint: int    (* Interactive interrupt (ctrl-C) *)
val sigkill: int   (* Termination (cannot be ignored) *)
val sigpipe: int   (* Broken pipe *)
val sigquit: int   (* Interactive termination *)
val sigsegv: int   (* Invalid memory reference *)
val sigterm: int   (* Termination *)
val sigusr1: int   (* Application-defined signal 1 *)
val sigusr2: int   (* Application-defined signal 2 *)
val sigchld: int   (* Child process terminated *)
val sigcont: int   (* Continue *)
val sigstop: int   (* Stop *)
val sigtstp: int   (* Interactive stop *)
val sigttin: int   (* Terminal read from background process *)
val sigttou: int   (* Terminal write from background process *)
val sigvtalrm: int (* Timeout in virtual time *)
val sigprof: int   (* Profiling interrupt *)

Signal numbers for the standard POSIX signals.

exception Break

Exception raised on interactive interrupt if catch_break is on.

val catch_break: bool -> unit

catch_break governs whether interactive interrupt (ctrl-C) terminates the program or raises the Break exception. Call catch_break true to enable raising Break, and catch_break false to let the system terminate the program on user interrupt.

Chapter 16. The standard library 215

16.28 Module Weak: arrays of weak pointers

type 'a t;;

The type of arrays of weak pointers (weak arrays). A weak pointer is an object that the garbage collector may erase at any time. A weak pointer is said to be full if it points to an object, empty if the object was erased by the GC.

val create : int -> 'a t;;

Weak.create n returns a new weak array of length n. All the pointers are initially empty.

val length : 'a t -> int;;

Weak.length ar returns the length (number of elements) of ar.

val set : 'a t -> int -> 'a option -> unit;;

Weak.set ar n (Some el) sets the nth cell of ar to be a (full) pointer to el; Weak.set ar n None sets the nth cell of ar to empty. Raise

Invalid_argument "Weak.set" if n is not in the range 0 to Weak.length a - 1.
val get : 'a t -> int -> 'a option;;

Weak.get ar n returns None if the nth cell of ar is empty, Some x (where x is the object) if it is full. Raise Invalid_argument "Weak.get" if n is not in the range 0 to

Weak.length a - 1.
val fill: 'a t -> int -> int -> 'a option -> unit;;
Weak.fill ar ofs len el sets to el all pointers of ar from ofs to ofs + len - 1. Raise

Invalid_argument "Weak.fill" if ofs and len do not designate a valid subarray of a.

val blit : 'a t -> int -> 'a t -> int -> int -> unit;;
Weak.blit ar1 off1 ar2 off2 len copies len weak pointers from ar1 (starting at off1)

to ar2 (starting at off2). It works correctly even if ar1 and ar2 are the same. Raise

Invalid_argument "Weak.blit" if off1 and len do not designate a valid subarray of ar1,

or if off2 and len do not designate a valid subarray of ar2.

216

The unix library: Unix system calls

The unix library makes many Unix system calls and system-related library functions available to Objective Caml programs. This chapter describes briefly the functions provided. Refer to sections 2 and 3 of the Unix manual for more details on the behavior of these functions.

Not all functions are provided by all Unix variants. If some functions are not available, they will raise Invalid_arg when called.

Unix:

Programs that use the unix library must be linked in “custom runtime” mode, as follows:

ocamlc -custom other options unix.cma other files -cclib -lunix
ocamlopt other options unix.cmxa other files -cclib -lunix

For interactive use of the unix library, do:

ocamlmktop -custom -o mytop unix.cma -cclib -lunix
./mytop

Windows:

A fairly complete emulation of the Unix system calls is provided in the Windows version of Objective Caml. The end of this chapter gives more information on the functions that are not supported under Windows.

Programs that use the unix library must be linked in “custom runtime” mode, as follows:

ocamlc -custom other options unix.cma other files \
        %CAMLLIB%/libunix.lib wsock32.lib

ocamlopt other options unix.cmxa other files \

        %CAMLLIB%/libunix.lib wsock32.lib

For interactive use of the unix library, do:

ocamlmktop -custom -o mytop.exe unix.cma \
        %CAMLLIB%/libunix.lib wsock32.lib
mytop.exe

17.1 Module Unix: interface to the Unix system

Error report

type error =

Errors defined in the POSIX standard

  E2BIG               (* Argument list too long *)
| EACCES              (* Permission denied *)
| EAGAIN              (* Resource temporarily unavailable; try again *)
| EBADF               (* Bad file descriptor *)
| EBUSY               (* Resource unavailable *)
| ECHILD              (* No child process *)
| EDEADLK             (* Resource deadlock would occur *)
| EDOM                (* Domain error for math functions, etc. *)
| EEXIST              (* File exists *)
| EFAULT              (* Bad address *)
| EFBIG               (* File too large *)
| EINTR               (* Function interrupted by signal *)
| EINVAL              (* Invalid argument *)
| EIO                 (* Hardware I/O error *)
| EISDIR              (* Is a directory *)
| EMFILE              (* Too many open files by the process *)
| EMLINK              (* Too many links *)
| ENAMETOOLONG        (* Filename too long *)
| ENFILE              (* Too many open files in the system *)
| ENODEV              (* No such device *)
| ENOENT              (* No such file or directory *)
| ENOEXEC             (* Not an executable file *)
| ENOLCK              (* No locks available *)
| ENOMEM              (* Not enough memory *)
| ENOSPC              (* No space left on device *)
| ENOSYS              (* Function not supported *)
| ENOTDIR             (* Not a directory *)
| ENOTEMPTY           (* Directory not empty *)
| ENOTTY              (* Inappropriate I/O control operation *)
| ENXIO               (* No such device or address *)
| EPERM               (* Operation not permitted *)
| EPIPE               (* Broken pipe *)
| ERANGE              (* Result too large *)
| EROFS               (* Read-only file system *)
| ESPIPE              (* Invalid seek e.g. on a pipe *)
| ESRCH               (* No such process *)
| EXDEV               (* Invalid link *)

Additional errors, mostly BSD

| EWOULDBLOCK         (* Operation would block *)
| EINPROGRESS         (* Operation now in progress *)
| EALREADY            (* Operation already in progress *)
| ENOTSOCK            (* Socket operation on non-socket *)
| EDESTADDRREQ        (* Destination address required *)
| EMSGSIZE            (* Message too long *)
| EPROTOTYPE          (* Protocol wrong type for socket *)
| ENOPROTOOPT         (* Protocol not available *)
| EPROTONOSUPPORT     (* Protocol not supported *)
| ESOCKTNOSUPPORT     (* Socket type not supported *)
| EOPNOTSUPP          (* Operation not supported on socket *)
| EPFNOSUPPORT        (* Protocol family not supported *)
| EAFNOSUPPORT        (* Address family not supported by protocol family *)
| EADDRINUSE          (* Address already in use *)
| EADDRNOTAVAIL       (* Can't assign requested address *)
| ENETDOWN            (* Network is down *)
| ENETUNREACH         (* Network is unreachable *)
| ENETRESET           (* Network dropped connection on reset *)
| ECONNABORTED        (* Software caused connection abort *)
| ECONNRESET          (* Connection reset by peer *)
| ENOBUFS             (* No buffer space available *)
| EISCONN             (* Socket is already connected *)
| ENOTCONN            (* Socket is not connected *)
| ESHUTDOWN           (* Can't send after socket shutdown *)
| ETOOMANYREFS        (* Too many references: can't splice *)
| ETIMEDOUT           (* Connection timed out *)
| ECONNREFUSED        (* Connection refused *)
| EHOSTDOWN           (* Host is down *)
| EHOSTUNREACH        (* No route to host *)
| ELOOP               (* Too many levels of symbolic links *)

All other errors are mapped to EUNKNOWNERR

| EUNKNOWNERR         (* Unknown error *)

The type of error codes.

exception Unix_error of error * string * string

Raised by the system calls below when an error is encountered. The first component is the error code; the second component is the function name; the third component is the string parameter to the function, if it has one, or the empty string otherwise.

val error_message : error -> string

Return a string describing the given error code.

val handle_unix_error : ('a -> 'b) -> 'a -> 'b

handle_unix_error f x applies f to x and returns the result. If the exception Unix_error is raised, it prints a message describing the error and exits with code 2.

Interface with the parent process

val environment : unit -> string array

Return the process environment, as an array of strings with the format “variable=value”.

See also Sys.getenv.

Process handling

type process_status =
WEXITED of int
| WSIGNALED of int
| WSTOPPED of int

The termination status of a process. WEXITED means that the process terminated normally by exit; the argument is the return code. WSIGNALED means that the process was killed by a signal; the argument is the signal number. WSTOPPED means that the process was stopped by a signal; the argument is the signal number.

type wait_flag =
WNOHANG
| WUNTRACED

Flags for waitopt and waitpid. WNOHANG means do not block if no child has died yet, but immediately return with a pid equal to 0. WUNTRACED means report also the children that receive stop signals.

val execv : string -> string array -> unit

execv prog args execute the program in file prog, with the arguments args, and the current process environment.

val execve : string -> string array -> string array -> unit

Same as execv, except that the third argument provides the environment to the program executed.

val execvp : string -> string array -> unit
val execvpe : string -> string array -> string array -> unit

Same as execv and execvp respectively, except that the program is searched in the path.

val fork : unit -> int

Fork a new process. The returned integer is 0 for the child process, the pid of the child process for the parent process.

val wait : unit -> int * process_status

Wait until one of the children processes die, and return its pid and termination status.

val waitpid : wait_flag list -> int -> int * process_status

Same as wait, but waits for the process whose pid is given. A pid of -1 means wait for any child. A pid of 0 means wait for any child in the same process group as the current process. Negative pid arguments represent process groups. The list of options indicates whether waitpid should return immediately without waiting, or also report stopped children.

val system : string -> process_status

Execute the given command, wait until it terminates, and return its termination status. The string is interpreted by the shell /bin/sh and therefore can contain redirections, quotes, variables, etc. The result WEXITED 127 indicates that the shell couldn’t be executed.

val getpid : unit -> int

Return the pid of the process.

val getppid : unit -> int

Return the pid of the parent process.

val nice : int -> int

Change the process priority. The integer argument is added to the “nice” value. (Higher values of the “nice” value mean lower priorities.) Return the new nice value.

Basic file input/output

type file_descr

The abstract type of file descriptors.

val stdin : file_descr
val stdout : file_descr
val stderr : file_descr

File descriptors for standard input, standard output and standard error.

type open_flag =
O_RDONLY                            (* Open for reading *)
| O_WRONLY                            (* Open for writing *)
| O_RDWR                              (* Open for reading and writing *)
| O_NONBLOCK                          (* Open in non-blocking mode *)
| O_APPEND                            (* Open for append *)
| O_CREAT                             (* Create if nonexistent *)
| O_TRUNC                             (* Truncate to 0 length if existing *)
| O_EXCL                              (* Fail if existing *)

The flags to open.

type file_perm = int

The type of file access rights.

val openfile : string -> open_flag list -> file_perm -> file_descr

Open the named file with the given flags. Third argument is the permissions to give to the file if it is created. Return a file descriptor on the named file.

val close : file_descr -> unit

Close a file descriptor.

val read : file_descr -> string -> int -> int -> int

read fd buff ofs len reads len characters from descriptor fd, storing them in string buff, starting at position ofs in string buff. Return the number of characters actually read.

val write : file_descr -> string -> int -> int -> int

write fd buff ofs len writes len characters to descriptor fd, taking them from string buff, starting at position ofs in string buff. Return the number of characters actually written.

Interfacing with the standard input/output library.

val in_channel_of_descr : file_descr -> in_channel

Create an input channel reading from the given descriptor.

val out_channel_of_descr : file_descr -> out_channel

Create an output channel writing on the given descriptor.

val descr_of_in_channel : in_channel -> file_descr

Return the descriptor corresponding to an input channel.

val descr_of_out_channel : out_channel -> file_descr

Return the descriptor corresponding to an output channel.

Seeking and truncating

type seek_command =
SEEK_SET
| SEEK_CUR
| SEEK_END

Positioning modes for lseek. SEEK_SET indicates positions relative to the beginning of the file, SEEK_CUR relative to the current position, SEEK_END relative to the end of the file.

val lseek : file_descr -> int -> seek_command -> int

Set the current position for a file descriptor

val truncate : string -> int -> unit

Truncates the named file to the given size.

val ftruncate : file_descr -> int -> unit

Truncates the file corresponding to the given descriptor to the given size.

File statistics

type file_kind =
S_REG                               (* Regular file *)
| S_DIR                               (* Directory *)
| S_CHR                               (* Character device *)
| S_BLK                               (* Block device *)
| S_LNK                               (* Symbolic link *)
| S_FIFO                              (* Named pipe *)
| S_SOCK                              (* Socket *)
type stats =
{ st_dev : int;                       (* Device number *)
st_ino : int;                       (* Inode number *)
st_kind : file_kind;                (* Kind of the file *)
st_perm : file_perm;                (* Access rights *)
st_nlink : int;                     (* Number of links *)
st_uid : int;                       (* User id of the owner *)
st_gid : int;                       (* Group id of the owner *)
st_rdev : int;                      (* Device minor number *)
st_size : int;                      (* Size in bytes *)
st_atime : int;                     (* Last access time *)
st_mtime : int;                     (* Last modification time *)
st_ctime : int }                    (* Last status change time *)

The informations returned by the stat calls.

val stat : string -> stats

Return the information for the named file.

val lstat : string -> stats

Same as stat, but in case the file is a symbolic link, return the information for the link itself.

val fstat : file_descr -> stats

Return the information for the file associated with the given descriptor.

Operations on file names

val unlink : string -> unit

Removes the named file

val rename : string -> string -> unit

rename old new changes the name of a file from old to new.

val link : string -> string -> unit

link source dest creates a hard link named dest to the file named new.

File permissions and ownership

type access_permission =
    R_OK                                (* Read permission *)
  | W_OK                                (* Write permission *)
  | X_OK                                (* Execution permission *)
  | F_OK                                (* File exists *)

Flags for the access call.

val chmod : string -> file_perm -> unit

Change the permissions of the named file.

val fchmod : file_descr -> file_perm -> unit

Change the permissions of an opened file.

val chown : string -> int -> int -> unit

Change the owner uid and owner gid of the named file.

val fchown : file_descr -> int -> int -> unit

Change the owner uid and owner gid of an opened file.

val umask : int -> int

Set the process creation mask, and return the previous mask.

val access : string -> access_permission list -> unit

Check that the process has the given permissions over the named file. Raise Unix_error otherwise.

Chapter 17. The unix library: Unix system calls 225

Operations on file descriptors

val dup : file_descr -> file_descr

Return a new file descriptor referencing the same file as the given descriptor.

val dup2 : file_descr -> file_descr -> unit

dup2 fd1 fd2 duplicates fd1 to fd2, closing fd2 if already opened.

val set_nonblock : file_descr -> unit
val clear_nonblock : file_descr -> unit

Set or clear the “non-blocking” flag on the given descriptor. When the non-blocking flag is set, reading on a descriptor on which there is temporarily no data available raises the EAGAIN or EWOULDBLOCK error instead of blocking; writing on a descriptor on which there is temporarily no room for writing also raises EAGAIN or EWOULDBLOCK.

val set_close_on_exec : file_descr -> unit
val clear_close_on_exec : file_descr -> unit

Set or clear the “close-on-exec” flag on the given descriptor. A descriptor with the close-on-exec flag is automatically closed when the current process starts another program with one of the exec functions.

Directories

val mkdir : string -> file_perm -> unit

Create a directory with the given permissions.

val rmdir : string -> unit

Remove an empty directory.

val chdir : string -> unit

Change the process working directory.

val getcwd : unit -> string

Return the name of the current working directory.

type dir_handle

The type of descriptors over opened directories.

val opendir : string -> dir_handle

Open a descriptor on a directory

val readdir : dir_handle -> string

Return the next entry in a directory. Raise End_of_file when the end of the directory has been reached.

val rewinddir : dir_handle -> unit

Reposition the descriptor to the beginning of the directory

val closedir : dir_handle -> unit

Close a directory descriptor.

Pipes and redirections

val pipe : unit -> file_descr * file_descr

Create a pipe. The first component of the result is opened for reading, that’s the exit to the pipe. The second component is opened for writing, that’s the entrance to the pipe.

val mkfifo : string -> file_perm -> unit

Create a named pipe with the given permissions.

High-level process and redirection management

val create_process :
string -> string array -> file_descr -> file_descr -> file_descr -> int
create_process prog args new_stdin new_stdout new_stderr forks a new process that

executes the program in file prog, with arguments args. The pid of the new process is returned immediately; the new process executes concurrently with the current process. The standard input and outputs of the new process are connected to the descriptors new_stdin, new_stdout and new_stderr. Passing e.g. stdout for new_stdout prevents the redirection and causes the new process to have the same standard output as the current process. The executable file prog is searched in the path. The new process has the same environment as the current process. All file descriptors of the current process are closed in the new process, except those redirected to standard input and outputs.

val create_process_env :
string -> string array -> string array ->
file_descr -> file_descr -> file_descr -> int
create_process_env prog args env new_stdin new_stdout new_stderr works as

create_process, except that the extra argument env specifies the environment passed to the program.

val open_process_in: string -> in_channel
val open_process_out: string -> out_channel
val open_process: string -> in_channel * out_channel

High-level pipe and process management. These functions run the given command in parallel with the program, and return channels connected to the standard input and/or the standard output of the command. The command is interpreted by the shell /bin/sh (cf. system). Warning: writes on channels are buffered, hence be careful to call flush at the right times to ensure correct synchronization.

val close_process_in: in_channel -> process_status
val close_process_out: out_channel -> process_status
val close_process: in_channel * out_channel -> process_status
Close channels opened by open_process_in, open_process_out and open_process,

respectively, wait for the associated command to terminate, and return its termination status.

Symbolic links

val symlink : string -> string -> unit

symlink source dest creates the file dest as a symbolic link to the file source.

val readlink : string -> string

Read the contents of a link.

Polling

val select :
file_descr list -> file_descr list -> file_descr list -> float ->
   file_descr list * file_descr list * file_descr list

Wait until some input/output operations become possible on some channels. The three list arguments are, respectively, a set of descriptors to check for reading (first argument), for writing (second argument), or for exceptional conditions (third argument). The fourth argument is the maximal timeout, in seconds; a negative fourth argument means no timeout (unbounded wait). The result is composed of three sets of descriptors: those ready for reading (first component), ready for writing (second component), and over which an exceptional condition is pending (third component).

Locking

type lock_command =
F_ULOCK               (* Unlock a region *)
| F_LOCK                (* Lock a region, and block if already locked *)
| F_TLOCK               (* Lock a region, or fail if already locked *)
| F_TEST                (* Test a region for other process' locks *)

Commands for lockf.

val lockf : file_descr -> lock_command -> int -> unit

lockf fd cmd size puts a lock on a region of the file opened as fd. The region starts at the current read/write position for fd (as set by lseek), and extends size bytes forward if size is positive, size bytes backwards if size is negative, or to the end of the file if size is zero.

Signals

val kill : int -> int -> unit

kill pid sig sends signal number sig to the process with id pid.

val pause : unit -> unit

Wait until a non-ignored signal is delivered.

Time functions

type process_times =
{ tms_utime : float;          (* User time for the process *)
tms_stime : float;          (* System time for the process *)
tms_cutime : float;         (* User time for the children processes *)
tms_cstime : float }        (* System time for the children processes *)

The execution times (CPU times) of a process.

type tm =
{ tm_sec : int;                       (* Seconds 0..59 *)
tm_min : int;                       (* Minutes 0..59 *)
tm_hour : int;                      (* Hours 0..23 *)
tm_mday : int;                      (* Day of month 1..31 *)
tm_mon : int;                       (* Month of year 0..11 *)
tm_year : int;                      (* Year - 1900 *)
tm_wday : int;                      (* Day of week (Sunday is 0) *)
tm_yday : int;                      (* Day of year 0..365 *)
tm_isdst : bool }                   (* Daylight time savings in effect *)

The type representing wallclock time and calendar date.

Chapter 17. The unix library: Unix system calls 229

val time : unit -> int

Return the current time since 00:00:00 GMT, Jan. 1, 1970, in seconds.

val gettimeofday : unit -> float

Same as time, but with resolution better than 1 second.

val gmtime : int -> tm

Convert a time in seconds, as returned by time, into a date and a time. Assumes Greenwich meridian time zone.

val localtime : int -> tm

Convert a time in seconds, as returned by time, into a date and a time. Assumes the local time zone.

val mktime : tm -> int * tm

Convert a date and time, specified by the tm argument, into a time in seconds, as returned by time. Also return a normalized copy of the given tm record, with the tm_wday and tm_yday recomputed from the other fields.

val alarm : int -> int

Schedule a SIGALRM signals after the given number of seconds.

val sleep : int -> unit

Stop execution for the given number of seconds.

val times : unit -> process_times =
         "unix_times_bytecode" "unix_times_native"

Return the execution times of the process.

val utimes : string -> int -> int -> unit

Set the last access time (second arg) and last modification time (third arg) for a file. Times are expressed in seconds from 00:00:00 GMT, Jan. 1, 1970.

type interval_timer =
ITIMER_REAL
| ITIMER_VIRTUAL
| ITIMER_PROF

The three kinds of interval timers. ITIMER_REAL decrements in real time, and sends the signal SIGALRM when expired. ITIMER_VIRTUAL decrements in process virtual time, and sends SIGVTALRM when expired. ITIMER_PROF (for profiling) decrements both when the process is running and when the system is running on behalf of the process; it sends SIGPROF when expired.

type interval_timer_status =
{ it_interval: float;                 (* Period *)
it_value: float }                   (* Current value of the timer *)

The type describing the status of an interval timer

val getitimer: interval_timer -> interval_timer_status

Return the current status of the given interval timer.

val setitimer:
interval_timer -> interval_timer_status -> interval_timer_status

setitimer t s sets the interval timer t and returns its previous status. The s argument is interpreted as follows: s.it_value, if nonzero, is the time to the next timer expiration; s.it_interval, if nonzero, specifies a value to be used in reloading it_value when the timer expires. Setting s.it_value to zero disable the timer. Setting s.it_interval to zero causes the timer to be disabled after its next expiration.

User id, group id

val getuid : unit -> int

Return the user id of the user executing the process.

val geteuid : unit -> int

Return the effective user id under which the process runs.

val setuid : int -> unit

Set the real user id and effective user id for the process.

val getgid : unit -> int

Return the group id of the user executing the process.

val getegid : unit -> int

Return the effective group id under which the process runs.

val setgid : int -> unit

Set the real group id and effective group id for the process.

val getgroups : unit -> int array

Return the list of groups to which the user executing the process belongs.

type passwd_entry =
  { pw_name : string;
    pw_passwd : string;
    pw_uid : int;
    pw_gid : int;
    pw_gecos : string;
    pw_dir : string;
    pw_shell : string }

Structure of entries in the passwd database.

type group_entry =
  { gr_name : string;
    gr_passwd : string;
    gr_gid : int;
    gr_mem : string array }

Structure of entries in the groups database.

val getlogin : unit -> string

Return the login name of the user executing the process.

val getpwnam : string -> passwd_entry

Find an entry in passwd with the given name, or raise Not_found.

val getgrnam : string -> group_entry

Find an entry in group with the given name, or raise Not_found.

val getpwuid : int -> passwd_entry

Find an entry in passwd with the given user id, or raise Not_found.

val getgrgid : int -> group_entry

Find an entry in group with the given group id, or raise Not_found.

Internet addresses

type inet_addr

The abstract type of Internet addresses.

val inet_addr_of_string : string -> inet_addr
val string_of_inet_addr : inet_addr -> string

Conversions between string with the format XXX.YYY.ZZZ.TTT and Internet addresses.

inet_addr_of_string raises Failure when given a string that does not match this format.

val inet_addr_any : inet_addr

A special Internet address, for use only with bind, representing all the Internet addresses that the host machine possesses.

Sockets

type socket_domain =
PF_UNIX                             (* Unix domain *)
| PF_INET                             (* Internet domain *)

The type of socket domains.

type socket_type =
SOCK_STREAM                         (* Stream socket *)
| SOCK_DGRAM                          (* Datagram socket *)
| SOCK_RAW                            (* Raw socket *)
| SOCK_SEQPACKET                      (* Sequenced packets socket *)

The type of socket kinds, specifying the semantics of communications.

type sockaddr =
ADDR_UNIX of string
| ADDR_INET of inet_addr * int

The type of socket addresses. ADDR_UNIX name is a socket address in the Unix domain; name is a file name in the file system. ADDR_INET(addr,port) is a socket address in the Internet domain; addr is the Internet address of the machine, and port is the port number.

val socket : socket_domain -> socket_type -> int -> file_descr

Create a new socket in the given domain, and with the given kind. The third argument is the protocol type; 0 selects the default protocol for that kind of sockets.

val socketpair :
   socket_domain -> socket_type -> int -> file_descr * file_descr

Create a pair of unnamed sockets, connected together.

val accept : file_descr -> file_descr * sockaddr

Accept connections on the given socket. The returned descriptor is a socket connected to the client; the returned address is the address of the connecting client.

val bind : file_descr -> sockaddr -> unit

Bind a socket to an address.

val connect : file_descr -> sockaddr -> unit

Connect a socket to an address.

val listen : file_descr -> int -> unit

Set up a socket for receiving connection requests. The integer argument is the maximal number of pending requests.

type shutdown_command =
    SHUTDOWN_RECEIVE                    (* Close for receiving *)
  | SHUTDOWN_SEND                       (* Close for sending *)
  | SHUTDOWN_ALL                        (* Close both *)

The type of commands for shutdown.

val shutdown : file_descr -> shutdown_command -> unit

Shutdown a socket connection. SHUTDOWN_SEND as second argument causes reads on the other end of the connection to return an end-of-file condition. SHUTDOWN_RECEIVE causes writes on the other end of the connection to return a closed pipe condition (SIGPIPE signal).

val getsockname : file_descr -> sockaddr

Return the address of the given socket.

val getpeername : file_descr -> sockaddr

Return the address of the host connected to the given socket.

type msg_flag =
    MSG_OOB
  | MSG_DONTROUTE
  | MSG_PEEK
     The flags for recv, recvfrom, send and sendto.
val recv : file_descr -> string -> int -> int -> msg_flag list -> int
val recvfrom :
        file_descr -> string -> int -> int -> msg_flag list -> int * sockaddr

Receive data from an unconnected socket.

val send : file_descr -> string -> int -> int -> msg_flag list -> int
val sendto :
        file_descr -> string -> int -> int -> msg_flag list -> sockaddr -> int

Send data over an unconnected socket.

type socket_option =
    SO_DEBUG               (* Record debugging information *)
  | SO_BROADCAST           (* Permit sending of broadcast messages *)
  | SO_REUSEADDR           (* Allow reuse of local addresses for bind *)
  | SO_KEEPALIVE           (* Keep connection active *)
  | SO_DONTROUTE           (* Bypass the standard routing algorithms *)
  | SO_OOBINLINE           (* Leave out-of-band data in line *)

The socket options settable with setsockopt.

val getsockopt : file_descr -> socket_option -> bool

Return the current status of an option in the given socket.

val setsockopt : file_descr -> socket_option -> bool -> unit

Set or clear an option in the given socket.

High-level network connection functions

val open_connection : sockaddr -> in_channel * out_channel

Connect to a server at the given address. Return a pair of buffered channels connected to the server. Remember to call flush on the output channel at the right times to ensure correct synchronization.

val shutdown_connection : in_channel -> unit

“Shut down” a connection established with open_connection; that is, transmit an end-of-file condition to the server reading on the other side of the connection.

val establish_server : (in_channel -> out_channel -> 'a) -> sockaddr -> unit

Establish a server on the given address. The function given as first argument is called for each connection with two buffered channels connected to the client. A new process is created for each connection. The function establish_server never returns normally.

Host and protocol databases

type host_entry =
{ h_name : string;
h_aliases : string array;
h_addrtype : socket_domain;
h_addr_list : inet_addr array }

Structure of entries in the hosts database.

type protocol_entry =
{ p_name : string;
p_aliases : string array;
p_proto : int }

Structure of entries in the protocols database.

type service_entry =
{ s_name : string;
s_aliases : string array;
s_port : int;
s_proto : string }

Structure of entries in the services database.

val gethostname : unit -> string

Return the name of the local host.

val gethostbyname : string -> host_entry

Find an entry in hosts with the given name, or raise Not_found.

val gethostbyaddr : inet_addr -> host_entry

Find an entry in hosts with the given address, or raise Not_found.

val getprotobyname : string -> protocol_entry

Find an entry in protocols with the given name, or raise Not_found.

val getprotobynumber : int -> protocol_entry

Find an entry in protocols with the given protocol number, or raise Not_found.

val getservbyname : string -> string -> service_entry

Find an entry in services with the given name, or raise Not_found.

val getservbyport : int -> string -> service_entry

Find an entry in services with the given service number, or raise Not_found.

Terminal interface

The following functions implement the POSIX standard terminal interface. They provide control over asynchronous communication ports and pseudo-terminals. Refer to the termios man page for a complete description.

type terminal_io = {

Input modes:

mutable c_ignbrk: bool;  (* Ignore the break condition. *)
mutable c_brkint: bool;  (* Signal interrupt on break condition. *)
mutable c_ignpar: bool;  (* Ignore characters with parity errors. *)
mutable c_parmrk: bool;  (* Mark parity errors. *)
mutable c_inpck: bool;   (* Enable parity check on input. *)
mutable c_istrip: bool;  (* Strip 8th bit on input characters. *)
mutable c_inlcr: bool;   (* Map NL to CR on input. *)
mutable c_igncr: bool;   (* Ignore CR on input. *)
mutable c_icrnl: bool;   (* Map CR to NL on input. *)
mutable c_ixon: bool;    (* Recognize XON/XOFF characters on input. *)
mutable c_ixoff: bool;   (* Emit XON/XOFF chars to control input flow. *)

Output modes:

mutable c_opost: bool;   (* Enable output processing. *)

Control modes:

mutable c_obaud: int;    (* Output baud rate (0 means close connection).*)
mutable c_ibaud: int;    (* Input baud rate. *)
mutable c_csize: int;    (* Number of bits per character (5-8). *)
mutable c_cstopb: int;   (* Number of stop bits (1-2). *)
mutable c_cread: bool;   (* Reception is enabled. *)
mutable c_parenb: bool;  (* Enable parity generation and detection. *)
mutable c_parodd: bool;  (* Specify odd parity instead of even. *)
mutable c_hupcl: bool;   (* Hang up on last close. *)
mutable c_clocal: bool;  (* Ignore modem status lines. *)

Local modes:

mutable c_isig: bool;    (* Generate signal on INTR, QUIT, SUSP. *)
mutable c_icanon: bool;  (* Enable canonical processing
                            (line buffering and editing) *)
mutable c_noflsh: bool;  (* Disable flush after INTR, QUIT, SUSP. *)
mutable c_echo: bool;    (* Echo input characters. *)
mutable c_echoe: bool;   (* Echo ERASE (to erase previous character). *)
mutable c_echok: bool;   (* Echo KILL (to erase the current line). *)
mutable c_echonl: bool;  (* Echo NL even if c_echo is not set. *)

Control characters:

mutable c_vintr: char;   (* Interrupt character (usually ctrl-C). *)
mutable c_vquit: char;   (* Quit character (usually ctrl-\). *)
mutable c_verase: char;  (* Erase character (usually DEL or ctrl-H). *)
mutable c_vkill: char;   (* Kill line character (usually ctrl-U). *)
mutable c_veof: char;    (* End-of-file character (usually ctrl-D). *)
mutable c_veol: char;    (* Alternate end-of-line char. (usually none). *)
mutable c_vmin: int;     (* Minimum number of characters to read
                            before the read request is satisfied. *)
mutable c_vtime: int;    (* Maximum read wait (in 0.1s units). *)
mutable c_vstart: char;  (* Start character (usually ctrl-Q). *)
mutable c_vstop: char    (* Stop character (usually ctrl-S). *)
}
val tcgetattr: file_descr -> terminal_io

Return the status of the terminal referred to by the given file descriptor.

type setattr_when = TCSANOW | TCSADRAIN | TCSAFLUSH
val tcsetattr: file_descr -> setattr_when -> terminal_io -> unit

Set the status of the terminal referred to by the given file descriptor. The second argument indicates when the status change takes place: immediately (TCSANOW), when all pending output has been transmitted (TCSADRAIN), or after flushing all input that has been received but not read (TCSAFLUSH). TCSADRAIN is recommended when changing the output parameters; TCSAFLUSH, when changing the input parameters.

val tcsendbreak: file_descr -> int -> unit

Send a break condition on the given file descriptor. The second argument is the duration of the break, in 0.1s units; 0 means standard duration (0.25s).

val tcdrain: file_descr -> unit

Waits until all output written on the given file descriptor has been transmitted.

type flush_queue = TCIFLUSH | TCOFLUSH | TCIOFLUSH
val tcflush: file_descr -> flush_queue -> unit

Discard data written on the given file descriptor but not yet transmitted, or data received but not yet read, depending on the second argument: TCIFLUSH flushes data received but not read, TCOFLUSH flushes data written but not transmitted, and TCIOFLUSH flushes both.

type flow_action = TCOOFF | TCOON | TCIOFF | TCION
val tcflow: file_descr -> flow_action -> unit

Suspend or restart reception or transmission of data on the given file descriptor, depending on the second argument: TCOOFF suspends output, TCOON restarts output, TCIOFF transmits a STOP character to suspend input, and TCION transmits a START character to restart input.

val setsid : unit -> int

Put the calling process in a new session and detach it from its controlling terminal.

Windows:

Below is a list of the functions that are not implemented, or only partially implemented, under Windows. Functions not mentioned are fully implemented and behave as described previously in this chapter.

Functions  Comment

fork not implemented, use create_process or

threads

wait not implemented, use waitpid waitpid can only wait for a given PID, not any child

process

getppid not implemented (meaningless under Windows) nice not implemented

truncate, ftruncate  not implemented

lstat, fstat not implemented link, symlink, readlink not implemented (no links under Windows) chmod, fchmod not implemented chown, fchown not implemented (make no sense on a DOS file

system)

umask not implemented set_nonblock, clear_nonblock implemented as dummy functions; use threads

instead of non-blocking I/O

rewinddir not implemented; re-open the directory instead mkfifo not implemented ioctl_int, ioctl_ptr not implemented (no ioctl in Windows) select implemented, but works only for sockets; use

threads if you need to wait on other kinds of file descriptors

lockf not implemented kill, pause not implemented (no inter-process signals in

Windows)

alarm, times not implemented

getitimer, setitimer  not implemented

getuid, getgid always return 1

getgid, getegid, getgroups  not implemented

setuid, setgid not implemented

getpwnam, getpwuid  always raise Not_found
getgrnam, getgrgid  always raise Not_found

type socket_domain the domain PF_UNIX is not supported; PF_INET

is fully supported

establish_server not implemented; use threads terminal functions (tc*) not implemented

The num library: arbitrary-precision rational arithmetic

The num library implements exact-precision rational arithmetic. It is built upon the state-of-the-art BigNum arbitrary-precision integer arithmetic package, and therefore achieves very high performance.

The functions provided in this library are fully documented in The CAML Numbers Reference Manual by Valérie Ménissier-Morain, technical report 141, INRIA, july 1992 (available by

anonymous FTP from ftp.inria.fr, directory INRIA/publications/RT, file RT-0141.ps.Z). A

summary of the functions is given below.

Unix:

Programs that use the num library must be linked in “custom runtime” mode, as follows:

             ocamlc -custom other options nums.cma other files -cclib -lnums
             ocamlopt other options nums.cmxa other files -cclib -lnums

For interactive use of the nums library, do:

             ocamlmktop -custom -o mytop nums.cma -cclib -lnums
             ./mytop

Windows:

Programs that use the num library must be linked in “custom runtime” mode, as follows:

             ocamlc -custom other options nums.cma other files %CAMLLIB%/libnums.lib
             ocamlopt other options nums.cmxa other files %CAMLLIB%/libnums.lib

For interactive use of the nums library, do:

             ocamlmktop -custom -o mytop.exe nums.cma %CAMLLIB%/libnums.lib
             ./mytop.exe

18.1 Module Num: operation on arbitrary-precision numbers

open Nat
open Big_int
open Ratio

Numbers (type num) are arbitrary-precision rational numbers, plus the special elements 1/0 (infinity) and 0/0 (undefined).

type num = Int of int | Big_int of big_int | Ratio of ratio

The type of numbers.

Arithmetic operations

val (+/) : num -> num -> num
val add_num : num -> num -> num

Addition

val minus_num : num -> num

Unary negation.

val (-/) : num -> num -> num
val sub_num : num -> num -> num

Subtraction

val (*/) : num -> num -> num
val mult_num : num -> num -> num

Multiplication

val square_num : num -> num

Squaring

val (//) : num -> num -> num
val div_num : num -> num -> num

Division

val quo_num : num -> num -> num
val mod_num : num -> num -> num

Euclidean division: quotient and remainder

val (**/) : num -> num -> num
val power_num : num -> num -> num

Exponentiation

val is_integer_num : num -> bool

Test if a number is an integer

val integer_num : num -> num
val floor_num : num -> num
val round_num : num -> num
val ceiling_num : num -> num

Approximate a number by an integer. floor_num n returns the largest integer smaller or equal to n. ceiling_num n returns the smallest integer bigger or equal to n. integer_num n returns the integer closest to n. In case of ties, rounds towards zero. round_num n returns the integer closest to n. In case of ties, rounds off zero.

val sign_num : num -> int

Return -1, 0 or 1 according to the sign of the argument.

val (=/) : num -> num -> bool
val (</) : num -> num -> bool
val (>/) : num -> num -> bool
val (<=/) : num -> num -> bool
val (>=/) : num -> num -> bool
val (<>/) : num -> num -> bool
val eq_num : num -> num -> bool
val lt_num : num -> num -> bool
val le_num : num -> num -> bool
val gt_num : num -> num -> bool
val ge_num : num -> num -> bool

Usual comparisons between numbers

val compare_num : num -> num -> int

Return -1, 0 or 1 if the first argument is less than, equal to, or greater than the second argument.

val max_num : num -> num -> num
val min_num : num -> num -> num

Return the greater (resp. the smaller) of the two arguments.

val abs_num : num -> num

Absolute value.

val succ_num: num -> num
     succ n is n+1

val pred_num: num -> num
     pred n is n-1
val incr_num: num ref -> unit

incr r is r:=!r+1, where r is a reference to a number.

val decr_num: num ref -> unit

decr r is r:=!r-1, where r is a reference to a number.

Coercions with strings

val string_of_num : num -> string

Convert a number to a string, using fractional notation.

val approx_num_fix : int -> num -> string
val approx_num_exp : int -> num -> string

Approximate a number by a decimal. The first argument is the required precision. The second argument is the number to approximate. approx_fix uses decimal notation; the first argument is the number of digits after the decimal point. approx_exp uses scientific (exponential) notation; the first argument is the number of digits in the mantissa.

val num_of_string : string -> num

Convert a string to a number.

Coercions between numerical types

val int_of_num : num -> int
val num_of_int : int -> num
val nat_of_num : num -> nat
val num_of_nat : nat -> num
val num_of_big_int : big_int -> num
val big_int_of_num : num -> big_int
val ratio_of_num : num -> ratio
val num_of_ratio : ratio -> num
val float_of_num : num -> float

18.2 Module Arith_status: flags that control rational arithmetic

val arith_status: unit -> unit

Print the current status of the arithmetic flags.

val get_error_when_null_denominator : unit -> bool
val set_error_when_null_denominator : bool -> unit

Get or set the flag null_denominator. When on, attempting to create a rational with a null denominator raises an exception. When off, rationals with null denominators are accepted. Initially: on.

val get_normalize_ratio : unit -> bool
val set_normalize_ratio : bool -> unit

Get or set the flag normalize_ratio. When on, rational numbers are normalized after each operation. When off, rational numbers are not normalized until printed. Initially: off.

val get_normalize_ratio_when_printing : unit -> bool
val set_normalize_ratio_when_printing : bool -> unit

Get or set the flag normalize_ratio_when_printing. When on, rational numbers are normalized before being printed. When off, rational numbers are printed as is, without normalization. Initially: on.

val get_approx_printing : unit -> bool
val set_approx_printing : bool -> unit

Get or set the flag approx_printing. When on, rational numbers are printed as a decimal approximation. When off, rational numbers are printed as a fraction. Initially: off.

val get_floating_precision : unit -> int
val set_floating_precision : int -> unit

Get or set the parameter floating_precision. This parameter is the number of digits displayed when approx_printing is on. Initially: 12.

244

The str library: regular expressions and string processing

The str library provides high-level string processing functions, some based on regular expressions. It is intended to support the kind of file processing that is usually performed with scripting languages

such as awk, perl or sed.

Unix:

Programs that use the str library must be linked in “custom runtime” mode, as follows:

ocamlc -custom other options str.cma other files -cclib -lstr
ocamlopt other options str.cmxa other files -cclib -lstr

For interactive use of the str library, do:

ocamlmktop -custom -o mytop str.cma -cclib -lstr
./mytop

Windows:

Programs that use the str library must be linked in “custom runtime” mode, as follows:

ocamlc -custom other options str.cma other files %CAMLLIB%/libstr.lib
ocamlopt other options str.cmxa other files %CAMLLIB%/libstr.lib

For interactive use of the str library, do:

ocamlmktop -custom -o mytop.exe str.cma %CAMLLIB%/libstr.lib
./mytop.exe

19.1 Module Str: regular expressions and high-level string processing

Regular expressions

type regexp

The type of compiled regular expressions.

val regexp: string -> regexp

Compile a regular expression. The syntax for regular expressions is the same as in Gnu Emacs. The special characters are

$^.*+?[]]. The following constructs are recognized:

\\[\smallskipamount]\verb. ‘ matches any character except newline

* (postfix) matches the previous expression zero, one or several times + (postfix) matches the previous expression one or several times ? (postfix) matches the previous expression once or not at all [..] character set; ranges are denoted with -, as in a-z; an initial ^, as in ^0-9, complements the set ^matches at beginning of line $ matches at end of line \| (infix) alternative between two expressions \(..\) grouping and naming of the enclosed expression \1 the text matched by the first \(...\) expression (\2 for the second expression, etc) \b matches word boundaries \ quotes special characters.

val regexp_case_fold: string -> regexp

Same as regexp, but the compiled expression will match text in a case-insensitive way: uppercase and lowercase letters will be considered equivalent.

String matching and searching

val string_match: regexp -> string -> int -> bool

string_match r s start tests whether the characters in s starting at position start match the regular expression r. The first character of a string has position 0, as usual.

val search_forward: regexp -> string -> int -> int

search_forward r s start searchs the string s for a substring matching the regular expression r. The search starts at position start and proceeds towards the end of the string. Return the position of the first character of the matched substring, or raise Not_found if no substring matches.

val search_backward: regexp -> string -> int -> int

Same as search_forward, but the search proceeds towards the beginning of the string.

val matched_string: string -> string

matched_string s returns the substring of s that was matched by the latest

string_match, search_forward or search_backward. The user must make sure that the

parameter s is the same string that was passed to the matching or searching function.

val match_beginning: unit -> int
val match_end: unit -> int

match_beginning() returns the position of the first character of the substring that was

matched by string_match, search_forward or search_backward. match_end() returns

the position of the character following the last character of the matched substring.

val matched_group: int -> string -> string

matched_group n s returns the substring of s that was matched by the nth group \(...\) of the regular expression during the latest string_match, search_forward or search_backward. The user must make sure that the parameter s is the same string that was passed to the matching or searching function.

val group_beginning: int -> int
val group_end: int -> int

group_beginning n returns the position of the first character of the substring that was matched by the nth group of the regular expression. group_end n returns the position of the character following the last character of the matched substring.

Replacement

val global_replace: regexp -> string -> string -> string

global_replace regexp repl s returns a string identical to s, except that all substrings of s that match regexp have been replaced by repl. The replacement text repl can contain \1, \2, etc; these sequences will be replaced by the text matched by the corresponding group in the regular expression. \0 stands for the text matched by the whole regular expression.

val replace_first: regexp -> string -> string -> string

Same as global_replace, except that only the first substring matching the regular expression is replaced.

val global_substitute: regexp -> (string -> string) -> string -> string

global_substitute regexp subst s returns a string identical to s, except that all substrings of s that match regexp have been replaced by the result of function subst. The function subst is called once for each matching substring, and receives s (the whole text) as argument.

val substitute_first: regexp -> (string -> string) -> string -> string

Same as global_substitute, except that only the first substring matching the regular expression is replaced.

Splitting

val split: regexp -> string -> string list

split r s splits s into substrings, taking as delimiters the substrings that match r, and returns the list of substrings. For instance, split (regexp "[ \t]+") s splits s into blank-separated words.

val bounded_split: regexp -> string -> int -> string list

Same as split, but splits into at most n substrings, where n is the extra integer parameter.

Extracting substrings

val string_before: string -> int -> string

string_before s n returns the substring of all characters of s that precede position n (excluding the character at position n).

val string_after: string -> int -> string

string_after s n returns the substring of all characters of s that follow position n (including the character at position n).

val first_chars: string -> int -> string

first_chars s n returns the first n characters of s. This is the same function as

string_before.
val last_chars: string -> int -> string

last_chars s n returns the last n characters of s.

The threads library

The threads library allows concurrent programming in Objective Caml. It provides multiple threads of control (also called lightweight processes) that execute concurrently in the same memory space. Threads communicate by in-place modification of shared data structures, or by sending and receiving data on communication channels.

The threads library is implemented by time-sharing on a single processor. It will not take advantage of multi-processor machines. Using this library will therefore never make programs run faster. However, many programs are easier to write when structured as several communicating processes.

Unix:

Programs that use the threads library must be linked as follows:

        ocamlc -thread -custom other options threads.cma other files -cclib -lthreads

All object files on the command line must also have been compiled with the -thread option, which selects a special, thread-safe version of the standard library (see chapter 6).

The default thread implementation cannot be used in native-code programs compiled with ocamlopt. If your operating system provides POSIX 1003.1c compliant threads, you can select an alternate implementation when configuring Objective Caml (use the -with-pthread option to configure) which also supports native-code programs. Programs that use this alternate implementation of the threads library must be linked as follows:

        ocamlc -thread -custom other options threads.cma other files \
                 -cclib -lthreads -cclib -lunix -cclib -lpthread
        ocamlopt -thread other options threads.cmxa other files \
                -cclib -lthreadsnat -cclib -lunix -cclib -lpthread

Depending on the operating system, extra system libraries can be necessary. For instance, under Solaris 2.5, add -cclib -lposix4 at the end of the command line.

Windows:

Programs that use the threads library must be linked as follows:

        ocamlc -thread -custom other options threads.cma other files \
                %CAMLLIB%/libthreads.lib %CAMLLIB%/libunix.lib
        ocamlopt -thread -custom other options threads.cmxa other files \
                %CAMLLIB%/libthreadsnat.lib %CAMLLIB%/libunix.lib

All object files on the command line must also have been compiled with the -thread option, which selects a special, thread-safe version of the standard library (see chapter 6).

20.1 Module Thread: lightweight threads

type t

The type of thread handles.

Thread creation and termination

val create : ('a -> 'b) -> 'a -> t

Thread.create funct arg creates a new thread of control, in which the function application funct arg is executed concurrently with the other threads of the program. The application of Thread.create returns the handle of the newly created thread. The new thread terminates when the application funct arg returns, either normally or by raising an uncaught exception. In the latter case, the exception is printed on standard error, but not propagated back to the parent thread. Similarly, the result of the application funct arg is discarded and not directly accessible to the parent thread.

val self : unit -> t

Return the thread currently executing.

val id : t -> int

Return the identifier of the given thread. A thread identifier is an integer that identifies uniquely the thread. It can be used to build data structures indexed by threads.

val exit : unit -> unit

Terminate prematurely the currently executing thread.

val kill : t -> unit

Terminate prematurely the thread whose handle is given. This functionality is available only with bytecode-level threads.

Suspending threads

val delay: float -> unit

delay d suspends the execution of the calling thread for d seconds. The other program threads continue to run during this time.

val join : t -> unit

join th suspends the execution of the calling thread until the thread th has terminated.

val wait_read : Unix.file_descr -> unit
val wait_write : Unix.file_descr -> unit

Suspend the execution of the calling thread until at least one character is available for reading (wait_read) or one character can be written without blocking (wait_write) on the given Unix file descriptor.

val wait_timed_read : Unix.file_descr -> float -> bool
val wait_timed_write : Unix.file_descr -> float -> bool

Same as wait_read and wait_write, but wait for at most the amount of time given as second argument (in seconds). Return true if the file descriptor is ready for input/output and false if the timeout expired.

val select :
Unix.file_descr list -> Unix.file_descr list ->
Unix.file_descr list -> float ->
Unix.file_descr list * Unix.file_descr list * Unix.file_descr list

Suspend the execution of the calling thead until input/output becomes possible on the given Unix file descriptors. The arguments and results have the same meaning as for

Unix.select.
val wait_pid : int -> int * Unix.process_status

wait_pid p suspends the execution of the calling thread until the Unix process specified by the process identifier p terminates. A pid p of -1 means wait for any child. A pid of 0 means wait for any child in the same process group as the current process. Negative pid arguments represent process groups. Returns the pid of the child caught and its termination status, as per Unix.wait.

20.2 Module Mutex: locks for mutual exclusion

Mutexes (mutual-exclusion locks) are used to implement critical sections and protect shared mutable data structures against concurrent accesses. The typical use is (if m is the mutex associated with the data structure D):

     Mutex.lock m;
     (* Critical section that operates over D *);
     Mutex.unlock m

type t

The type of mutexes.

val create: unit -> t

Return a new mutex.

val lock: t -> unit

Lock the given mutex. Only one thread can have the mutex locked at any time. A thread that attempts to lock a mutex already locked by another thread will suspend until the other thread unlocks the mutex.

val try_lock: t -> bool

Same as try_lock, but does not suspend the calling thread if the mutex is already locked: just return false immediately in that case. If the mutex is unlocked, lock it and return

true.
val unlock: t -> unit

Unlock the given mutex. Other threads suspended trying to lock the mutex will restart.

20.3 Module Condition: condition variables to synchronize between threads

Condition variables are used when one thread wants to wait until another thread has finished doing something: the former thread “waits” on the condition variable, the latter thread “signals” the condition when it is done. Condition variables should always be protected by a mutex. The typical use is (if D is a shared data structure, m its mutex, and c is a condition variable):

     Mutex.lock m;
     while (* some predicate P over D is not satisfied *) do
       Condition.wait c m
     done;
     (* Modify D *)
     if (* the predicate P over D is now satified *) then Condition.signal c;
     Mutex.unlock m
type t

The type of condition variables.

val create: unit -> t

Return a new condition variable.

Chapter 20. The threads library 253

val wait: t -> Mutex.t -> unit

wait c m atomically unlocks the mutex m and suspends the calling process on the condition variable c. The process will restart after the condition variable c has been signalled. The mutex m is locked again before wait returns.

val signal: t -> unit

signal c restarts one of the processes waiting on the condition variable c.

val broadcast: t -> unit

broadcast c restarts all processes waiting on the condition variable c.

20.4 Module Event: first-class synchronous communication

This module implements synchronous interprocess communications over channels. As in John Reppy’s Concurrent ML system, the communication events are first-class values: they can be built and combined independently before being offered for communication.

type 'a channel

The type of communication channels carrying values of type 'a.

val new_channel: unit -> 'a channel

Return a new channel.

type 'a event

The type of communication events returning a result of type 'a.

val send: 'a channel -> 'a -> unit event

send ch v returns the event consisting in sending the value v over the channel ch. The result value of this event is ().

val receive: 'a channel -> 'a event

receive ch returns the event consisting in receiving a value from the channel ch. The result value of this event is the value received.

val always: 'a -> 'a event

always v returns an event that is always ready for synchronization. The result value of this event is v.

val choose: 'a event list -> 'a event

choose evl returns the event that is the alternative of all the events in the list evl.

val wrap: 'a event -> ('a -> 'b) -> 'b event

wrap ev fn returns the event that performs the same communications as ev, then applies the post-processing function fn on the return value.

val guard: (unit -> 'a event) -> 'a event

guard fn returns the event that, when synchronized, computes fn() and behaves as the resulting event. This allows to compute events with side-effects at the time of the synchronization operation.

val sync: 'a event -> 'a

“Synchronize” on an event: offer all the communication possibilities specified in the event to the outside world, and block until one of the communications succeed. The result value of that communication is returned.

val select: 'a event list -> 'a

“Synchronize” on an alternative of events. select evl is shorthand for sync(choose evl).

val poll: 'a event -> 'a option

Non-blocking version of sync: offer all the communication possibilities specified in the event to the outside world, and if one can take place immediately, perform it and return Some r where r is the result value of that communication. Otherwise, return None without blocking.

20.5 Module ThreadUnix: thread-compatible system calls

This module reimplements some of the functions from Unix so that they only block the calling thread, not all threads in the program, if they cannot complete immediately. See the documentation of the Unix module for more precise descriptions of the functions below.

Process handling

val execv : string -> string array -> unit
val execve : string -> string array -> string array -> unit
val execvp : string -> string array -> unit
val wait : unit -> int * Unix.process_status
val waitpid : Unix.wait_flag list -> int -> int * Unix.process_status
val system : string -> Unix.process_status

Basic input/output

val read : Unix.file_descr -> string -> int -> int -> int
val write : Unix.file_descr -> string -> int -> int -> int

Input/output with timeout

val timed_read : Unix.file_descr -> string -> int -> int -> float -> int
val timed_write : Unix.file_descr -> string -> int -> int -> float -> int

Behave as read and write, except that Unix_error(ETIMEDOUT,_,_) is raised if no data is available for reading or ready for writing after d seconds. The delay d is given in the fifth argument, in seconds.

Polling

val select :
  Unix.file_descr list -> Unix.file_descr list ->
  Unix.file_descr list -> float ->
        Unix.file_descr list * Unix.file_descr list * Unix.file_descr list

Pipes and redirections

val pipe : unit -> Unix.file_descr * Unix.file_descr
val open_process_out: string -> out_channel
val open_process: string -> in_channel * out_channel

Time

val sleep : int -> unit

Sockets

val socket : Unix.socket_domain -> Unix.socket_type -> int -> Unix.file_descr
val socketpair : Unix.socket_domain -> Unix.socket_type -> int ->
                 Unix.file_descr * Unix.file_descr
val accept : Unix.file_descr -> Unix.file_descr * Unix.sockaddr
val connect : Unix.file_descr -> Unix.sockaddr -> unit
val recv : Unix.file_descr -> string -> int -> int -> Unix.msg_flag list -> int
val recvfrom : Unix.file_descr -> string -> int -> int ->
               Unix.msg_flag list -> int * Unix.sockaddr
val send : Unix.file_descr -> string -> int -> int ->
           Unix.msg_flag list -> int
val sendto : Unix.file_descr -> string -> int -> int ->
             Unix.msg_flag list -> Unix.sockaddr -> int
val open_connection : Unix.sockaddr -> in_channel * out_channel
val establish_server :
      (in_channel -> out_channel -> 'a) -> Unix.sockaddr -> unit

256

The graphics library

The graphics library provides a set of portable drawing primitives. Drawing takes place in a separate window that is created when open_graph is called.

Unix:

This library is implemented under the X11 windows system. Programs that use the graphics library must be linked as follows:

        ocamlc -custom other options graphics.cma other files \
                -cclib -lgraphics -cclib -lX11

For interactive use of the graphics library, do:

        ocamlmktop -custom -o mytop graphics.cma -cclib -lgraphics -cclib -lX11
        ./mytop

Here are the graphics mode specifications supported by open_graph on the X11 implementation of this library: the argument to open_graph has the format "display-name geometry", where display-name is the name of the X-windows display to connect to, and geometry is a standard X-windows geometry specification. The two components are separated by a space. Either can be omitted, or both. Examples:

open_graph "foo:0"

connects to the display foo:0 and creates a window with the default geometry

open_graph "foo:0 300x100+50-0"

connects to the display foo:0 and creates a window 300 pixels wide by 100 pixels tall, at location (50, 0)

open_graph " 300x100+50-0"

connects to the default display and creates a window 300 pixels wide by 100 pixels tall, at location (50, 0)

open_graph ""

connects to the default display and creates a window with the default geometry.

Windows:

This library is available only under the toplevel application ocamlwin.exe. Before using it, the Caml part of this library must be loaded in-core, either by typing

        #load  graphics.cmo;;

in the input windows or by using the “Load” entry of the “File” menu.

The screen coordinates are interpreted as shown in the figure below. Notice that the coordinate system used is the same as in mathematics: y increases from the bottom of the screen to the top of the screen, and angles are measured counterclockwise (in degrees). Drawing is clipped to the screen.

[figure: x; y; Screen; size_x(); size_y(); pixel at (x, y); x; y]

21.1 Module Graphics: machine-independent graphics primitives

exception Graphic_failure of string

Raised by the functions below when they encounter an error.

Initializations

val open_graph: string -> unit

Show the graphics window or switch the screen to graphic mode. The graphics window is cleared. The string argument is used to pass optional information on the desired graphics mode, the graphics window size, and so on. Its interpretation is implementation-dependent. If the empty string is given, a sensible default is selected.

val close_graph: unit -> unit

Delete the graphics window or switch the screen back to text mode.

val clear_graph : unit -> unit

Erase the graphics window.

val size_x : unit -> int
val size_y : unit -> int

Return the size of the graphics window. Coordinates of the screen pixels range over 0 .. size_x()-1 and 0 .. size_y()-1. Drawings outside of this rectangle are clipped, without causing an error. The origin (0,0) is at the lower left corner.

Colors

type color = int

A color is specified by its R, G, B components. Each component is in the range 0..255. The three components are packed in an int: 0xRRGGBB, where RR are the two hexadecimal digits for the red component, GG for the green component, BB for the blue component.

val rgb: int -> int -> int -> color

rgb r g b returns the integer encoding the color with red component r, green component g, and blue component b. r, g and b are in the range 0..255.

val set_color : color -> unit

Set the current drawing color.

val black : color
val white : color
val red : color
val green : color
val blue : color
val yellow : color
val cyan : color
val magenta : color

Some predefined colors.

val background: color
val foreground: color

Default background and foreground colors (usually, either black foreground on a white background or white foreground on a black background). clear_graph fills the screen with the background color. The initial drawing color is foreground.

Point and line drawing

val plot : int -> int -> unit

Plot the given point with the current drawing color.

val point_color : int -> int -> color

Return the color of the given point.

val moveto : int -> int -> unit

Position the current point.

val current_point : unit -> int * int

Return the position of the current point.

val lineto : int -> int -> unit

Draw a line with endpoints the current point and the given point, and move the current point to the given point.

val draw_arc : int -> int -> int -> int -> int -> int -> unit

draw_arc x y rx ry a1 a2 draws an elliptical arc with center x,y, horizontal radius rx, vertical radius ry, from angle a1 to angle a2 (in degrees). The current point is unchanged.

val draw_ellipse : int -> int -> int -> int -> unit

draw_ellipse x y rx ry draws an ellipse with center x,y, horizontal radius rx and vertical radius ry. The current point is unchanged.

val draw_circle : int -> int -> int -> unit

draw_circle x y r draws a circle with center x,y and radius r. The current point is unchanged.

val set_line_width : int -> unit

Set the width of points and lines drawn with the functions above. Under X Windows, set_line_width 0 selects a width of 1 pixel and a faster, but less precise drawing algorithm than the one used when set_line_width 1 is specified.

Chapter 21. The graphics library 261

Text drawing

val draw_char : char -> unit
val draw_string : string -> unit

Draw a character or a character string with lower left corner at current position. After drawing, the current position is set to the lower right corner of the text drawn.

val set_font : string -> unit
val set_text_size : int -> unit

Set the font and character size used for drawing text. The interpretation of the arguments to set_font and set_text_size is implementation-dependent.

val text_size : string -> int * int

Return the dimensions of the given text, if it were drawn with the current font and size.

Filling

val fill_rect : int -> int -> int -> int -> unit

fill_rect x y w h fills the rectangle with lower left corner at x,y, width w and heigth h, with the current color.

val fill_poly : (int * int) array -> unit

Fill the given polygon with the current color. The array contains the coordinates of the vertices of the polygon.

val fill_arc : int -> int -> int -> int -> int -> int -> unit

Fill an elliptical pie slice with the current color. The parameters are the same as for

draw_arc.
val fill_ellipse : int -> int -> int -> int -> unit

Fill an ellipse with the current color. The parameters are the same as for draw_ellipse.

val fill_circle : int -> int -> int -> unit

Fill a circle with the current color. The parameters are the same as for draw_circle.

Images

type image

The abstract type for images, in internal representation. Externally, images are represented as matrices of colors.

val transp : color

In matrices of colors, this color represent a “transparent” point: when drawing the corresponding image, all pixels on the screen corresponding to a transparent pixel in the image will not be modified, while other points will be set to the color of the corresponding point in the image. This allows superimposing an image over an existing background.

val make_image : color array array -> image

Convert the given color matrix to an image. Each sub-array represents one horizontal line. All sub-arrays must have the same length; otherwise, exception Graphic_failure is raised.

val dump_image : image -> color array array

Convert an image to a color matrix.

val draw_image : image -> int -> int -> unit

Draw the given image with lower left corner at the given point.

val get_image : int -> int -> int -> int -> image

Capture the contents of a rectangle on the screen as an image. The parameters are the same

as for fill_rect.
val create_image : int -> int -> image

create_image w h returns a new image w pixels wide and h pixels tall, to be used in conjunction with blit_image. The initial image contents are random.

val blit_image : image -> int -> int -> unit

blit_image img x y copies screen pixels into the image img, modifying img in-place. The pixels copied are those inside the rectangle with lower left corner at x,y, and width and height equal to those of the image.

Mouse and keyboard events

type status =
{ mouse_x : int;              (* X coordinate of the mouse *)
mouse_y : int;              (* Y coordinate of the mouse *)
button : bool;              (* true if a mouse button is pressed *)
keypressed : bool;          (* true if a key has been pressed *)
key : char }                (* the character for the key pressed *)

To report events.

type event =
Button_down                 (* A mouse button is pressed *)
| Button_up                   (* A mouse button is released *)
| Key_pressed                 (* A key is pressed *)
| Mouse_motion                (* The mouse is moved *)
| Poll                        (* Don't wait; return immediately *)

To specify events to wait for.

val wait_next_event : event list -> status

Wait until one of the events specified in the given event list occurs, and return the status of the mouse and keyboard at that time. If Poll is given in the event list, return immediately with the current status. If the mouse cursor is outside of the graphics window, the mouse_x and mouse_y fields of the event are outside the range 0..size_x()-1, 0..size_y()-1. Keypresses are queued, and dequeued one by one when the Key_pressed event is specified.

Mouse and keyboard polling

val mouse_pos : unit -> int * int

Return the position of the mouse cursor, relative to the graphics window. If the mouse cursor is outside of the graphics window, mouse_pos() returns a point outside of the range

0..size_x()-1, 0..size_y()-1.
val button_down : unit -> bool

Return true if the mouse button is pressed, false otherwise.

val read_key : unit -> char

Wait for a key to be pressed, and return the corresponding character. Keypresses are queued.

val key_pressed : unit -> bool

Return true if a keypress is available; that is, if read_key would not block.

Sound

val sound : int -> int -> unit

sound freq dur plays a sound at frequency freq (in hertz) for a duration dur (in milliseconds).

The dbm library: access to NDBM databases

The dbm library provides access to NDBM databases under Unix. NDBM databases maintain key/data associations, where both the key and the data are arbitrary strings. They support fairly large databases (several gigabytes) and can retrieve a keyed item in one or two file system accesses. Refer to the Unix manual pages for more information.

Unix:

Programs that use the dbm library must be linked in “custom runtime” mode, as follows:

             ocamlc -custom other options dbm.cma other files -cclib -lmldbm -cclib -lndbm
             ocamlopt other options dbm.cmxa other files -cclib -lmldbm -cclib -lndbm

For interactive use of the dbm library, do:

             ocamlmktop -custom -o mytop dbm.cma -cclib -lmldbm -cclib -lndbm
             ./mytop

Depending on the Unix system used, the -cclib -lndbm option is not always necessary, or the library may have another name than -lndbm.

Windows:

This library is not available.

22.1 Module Dbm: interface to the NDBM databases

type t

The type of file descriptors opened on NDBM databases.

type open_flag =
   Dbm_rdonly | Dbm_wronly | Dbm_rdwr | Dbm_create

Flags for opening a database (see opendbm).

exception Dbm_error of string

Raised by the following functions when an error is encountered.

val opendbm : string -> open_flag list -> int -> t

Open a descriptor on an NDBM database. The first argument is the name of the database (without the .dir and .pag suffixes). The second argument is a list of flags: Dbm_rdonly opens the database for reading only, Dbm_wronly for writing only, Dbm_rdwr for reading and writing; Dbm_create causes the database to be created if it does not already exist. The third argument is the permissions to give to the database files, if the database is created.

val close : t -> unit

Close the given descriptor.

val find : t -> string -> string

find db key returns the data associated with the given key in the database opened for the descriptor db. Raise Not_found if the key has no associated data.

val add : t -> string -> string -> unit

add db key data inserts the pair (key, data) in the database db. If the database already contains data associated with key, raise Dbm_error "Entry already exists".

val replace : t -> string -> string -> unit

replace db key data inserts the pair (key, data) in the database db. If the database already contains data associated with key, that data is discarded and silently replaced by the new data.

val remove : t -> string -> unit

remove db key data removes the data associated with key in db. If key has no associated

data, raise Dbm_error "dbm_delete".
val firstkey : t -> string
val nextkey : t -> string

Enumerate all keys in the given database, in an unspecified order. firstkey db returns the first key, and repeated calls to nextkey db return the remaining keys. Not_found is raised when all keys have been enumerated.

val iter : (string -> string -> 'a) -> t -> unit

iter f db applies f to each (key, data) pair in the database db. f receives key as first argument and data as second argument.

Part V Appendix

267

Index to the library

! (operator), 172 != (operator), 162 & (operator), 163 && (operator), 163

* (operator), 163 ** (operator), 165 **/ (operator), 240 *. (operator), 164 */ (operator), 240 + (operator), 163 +. (operator), 164 +/ (operator), 240

- (operator), 163 -. (operator), 164 -/ (operator), 240 / (operator), 163 /. (operator), 165 // (operator), 240 := (operator), 172 < (operator), 162 </ (operator), 241 <= (operator), 162 <=/ (operator), 241 <> (operator), 162 <>/ (operator), 241 = (operator), 162 =/ (operator), 241 == (operator), 162 > (operator), 162 >/ (operator), 241 >= (operator), 162 >=/ (operator), 241 @ (operator), 167 ^(operator), 166 || (operator), 163 ~- (operator), 163 ~-. (operator), 164

abs, 163

                                          abs_float, 165
                                          abs_num, 241

accept, 232, 255

                                          access, 224
                                          acos, 165

add, 193, 205, 266

                                          add_num, 240
                                          alarm, 229
                                          always, 253
                                          append, 179
                                          approx_num_exp, 242
                                          approx_num_fix, 242

Arg (module), 176

                                          argv, 212
                                          arith_status, 243
                                          Arith_status (module), 243

Array (module), 178

                                          asin, 165

asr (operator), 164

Assert_failure (exception), 85, 161

                                          assoc, 199
                                          assq, 199
                                          at_exit, 173
                                          atan, 165
                                          atan2, 165
                                          background, 259

Bad (exception), 177

                                          basename, 182
                                          big_int_of_num, 242

bind, 232
bits, 206
black, 259

blit, 179, 211, 215

blit_image, 262
blue, 259
bounded_split, 248

Break (exception), 214

broadcast, 253
button_down, 263

Callback (module), 180

capitalize, 212
catch, 203
catch_break, 214
ceil, 165
ceiling_num, 241
channel, 181

Char (module), 180 chdir, 213, 225

check_suffix, 182
chmod, 224
choose, 253
chop_extension, 182
chop_suffix, 182
chown, 224

chr, 180 clear, 193, 205, 209

clear_close_on_exec, 225
clear_graph, 258
clear_nonblock, 225
clear_parser, 203

close, 222, 266

close_box, 183
close_graph, 258
close_in, 172
close_out, 170
close_process, 227
close_process_in, 227
close_process_out, 227
close_tbox, 186
closedir, 226
code, 180
combine, 199
command, 213

                                          compact, 192
                                          compare, 162
                                          compare_num, 241

concat, 179, 182, 197, 211 Condition (module), 252 connect, 232, 255 copy, 179, 203, 211 cos, 165

                                          cosh, 165
                                          count, 210

create, 178, 193, 205, 208, 211, 215, 250, 252

                                          create_image, 262
                                          create_matrix, 178
                                          create_process, 226
                                          create_process_env, 226
                                          current, 178
                                          current_dir_name, 182
                                          current_point, 260
                                          cyan, 259
                                          data_size, 202

Dbm (module), 265

Dbm_error (exception), 266

                                          decr, 172
                                          decr_num, 242
                                          delay, 251
                                          descr_of_in_channel, 222
                                          descr_of_out_channel, 222

Digest (module), 181

                                          dirname, 182
                                          div_num, 240
                                          Division_by_zero (exception), 161
                                          draw_arc, 260
                                          draw_char, 261
                                          draw_circle, 260
                                          draw_ellipse, 260
                                          draw_image, 262
                                          draw_string, 261
                                          dump_image, 262

dup, 225

                                          dup2, 225
                                          empty, 210

Empty (exception), 205, 208 End_of_file (exception), 161

                                          environment, 220

eprintf, 189, 205

eq_num, 241
err_formatter, 188

Error (exception), 209

error_message, 219

escaped, 181, 211

establish_server, 234, 255

Event (module), 253 execv, 220, 254 execve, 220, 254 execvp, 220, 254

execvpe, 220
exists, 198
exists2, 198

exit, 173, 250 Exit (exception), 161 exp, 165

Failure (exception), 161, 209

failwith, 161
fchmod, 224
fchown, 224
file, 181
file_exists, 212

Filename (module), 182 fill, 179, 211, 215

fill_arc, 261
fill_circle, 261
fill_ellipse, 261
fill_poly, 261
fill_rect, 261

find, 193, 266

find_all, 193
first_chars, 248
firstkey, 266
flatten, 197

float, 166, 206

float_of_num, 242
float_of_string, 166
floor, 165
floor_num, 241
flush, 169
fold_left, 180, 197
fold_left2, 198
fold_right, 180, 197

                                          fold_right2, 198
                                          for_all, 198
                                          for_all2, 198

force, 85, 195

                                          force_newline, 185
                                          foreground, 259
                                          fork, 220

Format (module), 183 fprintf, 189, 204

                                          frexp, 165
                                          from, 209
                                          from_channel, 196, 202
                                          from_function, 196
                                          from_string, 196, 202

fst, 166

                                          fstat, 224
                                          ftruncate, 223
                                          full_init, 206
                                          full_major, 192

Gc (module), 189

                                          ge_num, 241

Genlex (module), 192 get, 178, 191, 210, 215

                                          get_approx_printing, 243
                                          get_ellipsis_text, 187
                                          get_error_when_null_denominator, 243
                                          get_floating_precision, 243
                                          get_formatter_output_functions, 187
                                          get_image, 262
                                          get_margin, 185
                                          get_max_boxes, 185
                                          get_max_indent, 185
                                          get_normalize_ratio, 243
                                          get_normalize_ratio_when_printing, 243

getcwd, 213, 225

                                          getegid, 230
                                          getenv, 213
                                          geteuid, 230
                                          getgid, 230
                                          getgrgid, 231
                                          getgrnam, 231
                                          getgroups, 230
                                          gethostbyaddr, 235
                                          gethostbyname, 234

gethostname, 234
getitimer, 230
getlogin, 231
getpeername, 233
getpid, 221
getppid, 221
getprotobyname, 235
getprotobynumber, 235
getpwnam, 231
getpwuid, 231
getservbyname, 235
getservbyport, 235
getsockname, 233
getsockopt, 233
gettimeofday, 229
getuid, 230
global_replace, 247
global_substitute, 247
gmtime, 229
Graphic_failure (exception), 258

Graphics (module), 258

green, 259
group_beginning, 247
group_end, 247
gt_num, 241
guard, 254
handle_unix_error, 219
hash, 195
hash_param, 195

HashedType (module type), 194 Hashtbl (module), 193 hd, 197

header_size, 202

id, 250

in_channel_length, 172
in_channel_of_descr, 222
incr, 172
incr_num, 242
index, 212
index_from, 212
inet_addr_any, 231
inet_addr_of_string, 231

init, 178, 206 input, 171, 181

                                          input_binary_int, 172
                                          input_byte, 171
                                          input_char, 171
                                          input_line, 171
                                          input_value, 172

int, 206

                                          int_of_num, 242
                                          int_of_string, 166
                                          integer_num, 241
                                          interactive, 213
                                          invalid_arg, 161
                                          Invalid_argument (exception), 161
                                          is_implicit, 182
                                          is_integer_num, 241
                                          is_relative, 182

iter, 179, 194, 197, 205, 209, 210, 266

                                          iter2, 198
                                          iteri, 180
                                          join, 251
                                          junk, 210
                                          key_pressed, 263

kill, 228, 250

land (operator), 164

                                          last_chars, 248

Lazy (module), 85, 195

                                          ldexp, 165
                                          le_num, 241

length, 178, 197, 205, 209, 210, 215

                                          lexeme, 196
                                          lexeme_char, 196
                                          lexeme_end, 197
                                          lexeme_start, 196

Lexing (module), 195

                                          lineto, 260
                                          link, 224
                                          list, 208

List (module), 197

                                          listen, 232
                                          lnot, 164
                                          localtime, 229
                                          lock, 252
                                          lockf, 228

log, 165

log10, 165

lor (operator), 164

lowercase, 181, 212
lseek, 223

lsl (operator), 164 lsr (operator), 164

lstat, 224
lt_num, 241

lxor (operator), 164

magenta, 259
major, 192

make, 178, 211 Make (functor), 194, 200, 208

make_formatter, 188
make_image, 262
make_lexer, 192
make_matrix, 178

map, 179, 197 Map (module), 199

map2, 198
mapi, 180

Marshal (module), 201

match_beginning, 247
match_end, 247

Match_failure (exception), 65–67, 160

matched_group, 247
matched_string, 247

max, 162

max_array_length, 213
max_int, 163
max_num, 241
max_string_length, 213

mem, 198

mem_assoc, 199
memq, 198
merge, 208

min, 162

min_int, 163
min_num, 241
minor, 191
minus_num, 240
mkdir, 225
mkfifo, 226
mktime, 229

mod (operator), 163

                                          mod_float, 165
                                          mod_num, 240
                                          modf, 165
                                          mouse_pos, 263
                                          moveto, 260
                                          mult_num, 240

Mutex (module), 251

                                          nat_of_num, 242
                                          new_channel, 253
                                          next, 210
                                          nextkey, 266
                                          nice, 221

not, 163

Not_found (exception), 161

                                          npeek, 210

nth, 197

Num (module), 240

                                          num_of_big_int, 242
                                          num_of_int, 242
                                          num_of_nat, 242
                                          num_of_ratio, 242
                                          num_of_string, 242
                                          of_channel, 209

of_list, 179, 209

                                          of_string, 209

Oo (module), 203

                                          open_box, 183
                                          open_connection, 234, 255
                                          open_graph, 258
                                          open_hbox, 186
                                          open_hovbox, 186
                                          open_hvbox, 186
                                          open_in, 171
                                          open_in_bin, 171
                                          open_in_gen, 171
                                          open_out, 169
                                          open_out_bin, 169
                                          open_out_gen, 169
                                          open_process, 227, 255
                                          open_process_in, 227
                                          open_process_out, 227, 255
                                          open_tbox, 186
                                          open_vbox, 186

opendbm, 266
opendir, 225
openfile, 222

or (operator), 163 OrderedType (module type), 199, 206

os_type, 213
out_channel_length, 170
out_channel_of_descr, 222

Out_of_memory (exception), 161 output, 170, 181

output_binary_int, 170
output_byte, 170
output_char, 169
output_string, 170
output_value, 170
over_max_boxes, 185
parse, 177

Parse_error (exception), 203 Parsing (module), 203

pause, 228

peek, 205, 210 Pervasives (module), 159 pipe, 226, 255

plot, 260
point_color, 260
poll, 254

pop, 209

pos_in, 172
pos_out, 170
power_num, 240
pp_close_box, 189
pp_close_tbox, 189
pp_force_newline, 189
pp_get_ellipsis_text, 189
pp_get_formatter_output_functions, 189
pp_get_margin, 189
pp_get_max_boxes, 189
pp_get_max_indent, 189
pp_open_box, 189
pp_open_hbox, 189
pp_open_hovbox, 189
pp_open_hvbox, 189
pp_open_tbox, 189
pp_open_vbox, 189

                                          pp_over_max_boxes, 189
                                          pp_print_as, 189
                                          pp_print_bool, 189
                                          pp_print_break, 189
                                          pp_print_char, 189
                                          pp_print_cut, 189
                                          pp_print_float, 189
                                          pp_print_flush, 189
                                          pp_print_if_newline, 189
                                          pp_print_int, 189
                                          pp_print_newline, 189
                                          pp_print_space, 189
                                          pp_print_string, 189
                                          pp_print_tab, 189
                                          pp_print_tbreak, 189
                                          pp_set_ellipsis_text, 189
                                          pp_set_formatter_out_channel, 189
                                          pp_set_formatter_output_functions, 189
                                          pp_set_margin, 189
                                          pp_set_max_boxes, 189
                                          pp_set_max_indent, 189
                                          pp_set_tab, 189
                                          pred, 163
                                          pred_num, 242
                                          prerr_char, 168
                                          prerr_endline, 168
                                          prerr_float, 168
                                          prerr_int, 168
                                          prerr_newline, 168
                                          prerr_string, 168
                                          print, 204
                                          print_as, 184
                                          print_bool, 184
                                          print_break, 184
                                          print_char, 167, 184
                                          print_cut, 184
                                          print_endline, 167
                                          print_float, 167, 184
                                          print_flush, 184
                                          print_if_newline, 185
                                          print_int, 167, 184
                                          print_newline, 167, 185
                                          print_space, 184
                                          print_stat, 191
                                          print_string, 167, 184

print_tab, 186
print_tbreak, 186

Printexc (module), 203 printf, 189, 205 Printf (module), 204

push, 209

Queue (module), 205

quo_num, 240
raise, 160

Random (module), 206

ratio_of_num, 242

read, 222, 254

read_float, 168
read_int, 168
read_key, 263
read_line, 168
readdir, 226
readlink, 227
really_input, 171
receive, 253

recv, 233, 255

recvfrom, 233, 255

red, 259 ref, 172

regexp, 246
regexp_case_fold, 246
register, 180
register_exception, 180

remove, 194, 212, 266 rename, 213, 224

replace, 266
replace_first, 247

rev, 197

rewinddir, 226

rgb, 259

rhs_end, 203
rhs_start, 203
rindex, 212
rindex_from, 212
rmdir, 225
round_num, 241
search_backward, 246
search_forward, 246

seek_in, 172
seek_out, 170

select, 227, 251, 254, 255

self, 250

send, 233, 253, 255 sendto, 233, 255 set, 178, 191, 211, 215 Set (module), 206

set_approx_printing, 243
set_close_on_exec, 225
set_color, 259
set_ellipsis_text, 187
set_error_when_null_denominator, 243
set_floating_precision, 243
set_font, 261
set_formatter_out_channel, 187
set_formatter_output_functions, 187
set_line_width, 260
set_margin, 185
set_max_boxes, 185
set_max_indent, 185
set_nonblock, 225
set_normalize_ratio, 243
set_normalize_ratio_when_printing, 243
set_tab, 186
set_text_size, 261
setgid, 230
setitimer, 230
setsid, 237
setsockopt, 233
setuid, 230
shutdown, 233
shutdown_connection, 234
sigabrt, 214
sigalrm, 214
sigchld, 214
sigcont, 214
sigfpe, 214
sighup, 214
sigill, 214
sigint, 214
sigkill, 214
sign_num, 241

signal, 214, 253

sigpipe, 214

sigprof, 214
sigquit, 214
sigsegv, 214
sigstop, 214
sigterm, 214
sigtstp, 214
sigttin, 214
sigttou, 214
sigusr1, 214
sigusr2, 214
sigvtalrm, 214

sin, 165

sinh, 165
size_x, 259
size_y, 259

sleep, 229, 255 snd, 166 socket, 232, 255

socketpair, 232, 255

Sort (module), 208

sound, 264

split, 199, 248

sprintf, 205
sqrt, 165
square_num, 240

Stack (module), 208 Stack_overflow (exception), 161 stat, 191, 224

std_formatter, 187

stderr, 167, 221 stdin, 167, 221 stdout, 167, 221 Str (module), 245 Stream (module), 209

string, 181

String (module), 210

string_after, 248
string_before, 248
string_match, 246
string_of_bool, 166
string_of_float, 166
string_of_inet_addr, 231
string_of_int, 166
string_of_num, 242

sub, 179, 211

                                          sub_num, 240
                                          substitute_first, 247
                                          substring, 181
                                          succ, 163
                                          succ_num, 241
                                          symbol_end, 203
                                          symbol_start, 203
                                          symlink, 227
                                          sync, 254

Sys (module), 212

Sys_error (exception), 161 system, 221, 254

                                          take, 205

tan, 165

                                          tanh, 165
                                          tcdrain, 237
                                          tcflow, 237
                                          tcflush, 237
                                          tcgetattr, 236
                                          tcsendbreak, 237
                                          tcsetattr, 236
                                          temp_file, 182
                                          text_size, 261

Thread (module), 250 ThreadUnix (module), 254

                                          time, 229
                                          timed_read, 255
                                          timed_write, 255
                                          times, 229

tl, 197

                                          to_buffer, 202
                                          to_channel, 201
                                          to_list, 179
                                          to_string, 202, 204
                                          total_size, 202
                                          transp, 262
                                          truncate, 166, 223
                                          try_lock, 252
                                          umask, 224
                                          uncapitalize, 212

Unix (module), 218

Unix_error (exception), 219

                                          unlink, 224
                                          unlock, 252

uppercase, 181, 212
usage, 178
utimes, 229

wait, 220, 253, 254

wait_next_event, 263
wait_pid, 251
wait_read, 251
wait_timed_read, 251
wait_timed_write, 251

                                          wait_write, 251

waitpid, 221, 254 Weak (module), 215

                                          white, 259
                                          word_size, 213
                                          wrap, 254

write, 222, 254

                                          yellow, 259

278

Index of keywords

and, see let, type, 75

as, 59–61, 73–75

assert, 85

begin, 63, 64

class, 73, 75, 78, 81

closed, 73, 75
constraint, 71, 73–75
do, see while, for
done, see while, for
downto, see for
else, see if

end, 63, 64, 73, 75–77, 79, 80 exception, 73, 76, 78, 79, 81 external, 76, 77, 79, 80

false, 56

for, 63, 67 fun, 63–65

function, 63–65

functor, 76, 79, 82

if, 63, 64, 67

in, see let
inherit, 73–75
lazy, 85

let, 63, 64, 66, 79, 80

match, 63, 64, 67

                                          method, 73, 75

module, 76, 78, 79, 81

                                          mutable, 71–75

new, 63, 70

                                          not, 64
                                          of, see type, exception

open, 76, 79, 81 or, 63, 64, 67

                                          parser, 83
                                          private, 73–75
                                          protected, 73, 75
                                          rec, see let

sig, 76, 77

                                          struct, 79, 80

then, see if, 75

                                          to, see for
                                          true, 56

try, 63, 64, 68 type, 71, 76–79, 81

val, 73–77

                                          virtual, 73, 75

when, 63, 65

                                          while, 67
                                          with, see match, try, 76, 79